### v2

# Information Physics of Intelligence: Unifying Logical Depth and Entropy under Thermodynamic Constraints

Abstract
The rapid scaling of artificial intelligence models has revealed a fundamental tension between model capacity (storage) and inference efficiency (computation). While classical information theory focuses on transmission and storage limits, it lacks a unified physical framework to quantify the thermodynamic costs of generating information from compressed laws versus retrieving it from memory. In this paper, we propose a theoretical framework that treats information processing as an enabling mapping from ontological states to carrier states. We introduce a novel metric, Derivation Entropy, which quantifies the effective work required to compute a target state from a given logical depth. By analyzing the interplay between Shannon entropy (storage) and computational complexity (time/energy), we demonstrate the existence of a critical phase transition point. Below this threshold, memory retrieval is thermodynamically favorable; above it, generative computation becomes the optimal strategy. This ”Energy-Time-Space” conservation law provides a physical explanation for the efficiency of generative models and offers a rigorous mathematical bound for designing next-generation, energy-efficient AI architectures. Our findings suggest that the minimization of Derivation Entropy is a governing principle for the evolution of both biological and artificial intelligence.
I Introduction
The fundamental nature of information has long been a subject of debate across physics, computer science, and philosophy. In our previous works, we established a formalized meta-framework positing that information is fundamentally the enabling mapping from an ontological state to a carrier state [xu2025information, xu2024research, qiu2025research]. Under this view, intelligence is not merely the accumulation of data, but an active physical process of mapping the infinite complexity of the external environment (ontology) onto the finite constraints of an internal representation (carrier). This mapping inherently necessitates a primal trade-off: the system must decide whether to store the mapping results statically (memory) or to regenerate them dynamically via algorithms (computation).
Classically, these two aspects have been treated separately. Shannon’s information theory [shannon1948mathematical] rigorously defines the limits of static storage and transmission, quantifying the irreducible randomness of the carrier state. However, Shannon entropy is agnostic to the causal history or the computational effort required to generate that state. Conversely, Algorithmic Information Theory (AIT), particularly through the lens of Kolmogorov complexity [kolmogorov1965three] and Bennett’s concept of Logical Depth [bennett1988logical], focuses on the computational time required to reconstruct a structure from its compressed description. While Logical Depth captures the dynamic value of the ”mapping process,” it lacks a direct unification with the statistical mechanics governing the carrier’s physical existence.
This separation becomes problematic when we consider that information processing is strictly bound by physical laws. As Brillouin [brillouin1956science] and Landauer [landauer1961irreversibility] demonstrated, the manipulation of information—specifically the erasure of logical states—is inseparably linked to thermodynamic entropy and heat dissipation. Therefore, the ”enabling mapping” from ontology to carrier is not just a logical operation but a thermodynamic cycle. Current artificial intelligence systems, such as Large Language Models (LLMs) based on the Transformer architecture [vaswani2017attention, brown2020language], are rapidly approaching the physical limits of this cycle. The empirical scaling laws observed in these systems suggest a deep, underlying fundamental trade-off connecting parameter size (storage), training FLOPs (computation), and energy consumption, yet a unified theoretical explanation remains elusive.
In this paper, we propose the ”Information Physics of Intelligence,” a theoretical framework that unifies Logical Depth and Shannon Entropy under rigorous thermodynamic constraints. We define Derivation Entropy ($H_{derive}$), which quantifies the total physical cost of the ontological-to-carrier mapping. By integrating the thermodynamic cost of storage with the algorithmic complexity of retrieval, we derive the Energy-Time-Space Triality Bound. This bound reveals that minimizing storage space (lowering Shannon entropy) inevitably necessitates a compensatory increase in computational time (increasing Logical Depth) or energy flux, governed by Landauer’s limit.
Our contributions are threefold:
  1. 1.
    We formalize the ”ontological-to-carrier mapping” within a thermodynamic context, bridging the gap between abstract information measures [xu2025general] and physical complexity [bennett1982thermodynamics].
  2. 2.
    We mathematically derive the phase transition point where intelligent systems must switch from a retrieval-based strategy to a derivation-based strategy to maintain thermodynamic optimality.
  3. 3.
    We provide a first-principles explanation for the efficiency bottlenecks in modern AI models, offering a governing equation for designing next-generation, energy-efficient cognitive architectures.
II System Model and Preliminaries
We establish the formal framework for knowledge systems, introducing the concepts of enabling mappings, logical depth, and their information-theoretic characterization.
II-A Knowledge Systems
Definition 1 (Knowledge System).
A knowledge system is a tuple $\mathcal{K}=(K,Q,\mathcal{L},\vdash,P_{Q})$ where:
  • $K$ is a knowledge base consisting of facts and rules
  • $Q$ is a query space
  • $\mathcal{L}$ is a logical system with atomic propositions $\mathcal{A}$
  • $\vdash$ is an inference system
  • $P_{Q}$ is a probability distribution over queries
Definition 2 (Query Answering).
Given knowledge base $K$ and query $q\in Q$, the answer $a=\mathcal{A}(q,K)$ is derived through inference: $K\vdash q\Rightarrow a$.
II-B Enabling Mappings and Information
Definition 3 (Enabling Mapping).
Let $X$, $Y$ be sets with state spaces $\mathcal{S}(X)$ and $\mathcal{S}(Y)$. A mapping $f:\mathcal{S}(X)\to\mathcal{S}(Y)$ is an enabling mapping (denoted $f:\mathcal{S}(X)\Rightarrow\mathcal{S}(Y)$) if it satisfies:
  1. 1.
    Surjectivity: $\forall y\in\mathcal{S}(Y),\exists x\in\mathcal{S}(X):y=f(x)$;
  2. 2.
    Physical causation: The realization of state $y$ is physically dependent on state $x$;
  3. 3.
    Energy requirement: The transformation requires minimum energy
    $E(f)\geq k_{B}T\ln\left(\frac{|\mathcal{S}(X)|}{|\mathrm{im}(f)|}\right),$ (1)
    where $\mathrm{im}(f)=\{f(x):x\in\mathcal{S}(X)\}$ is the image of $f$.
    For surjective mappings, $|\mathrm{im}(f)|=|\mathcal{S}(Y)|$, so
    $E(f)\geq k_{B}T\ln\left(\frac{|\mathcal{S}(X)|}{|\mathcal{S}(Y)|}\right).$ (2)
Physical interpretation: This reflects Landauer’s principle—compressing from $|\mathcal{S}(X)|$ states to $|\mathrm{im}(f)|$ states erases $\log_{2}\left(|\mathcal{S}(X)|/|\mathrm{im}(f)|\right)$ bits of information, requiring at least $k_{B}T\ln\left(|\mathcal{S}(X)|/|\mathrm{im}(f)|\right)$ joules of energy to be dissipated as heat.
Remark 1 (Energy Budget Interpretation).
The energy $E$ in the following theorem represents the total thermodynamic cost of the mapping process. It accounts for:
  • Erasure cost (irreversible): Information compression from $\mathcal{S}(O,T_{0})$ to $\mathcal{S}(C,T_{r})$ requires energy dissipation as specified by Landauer’s principle (Definition 3).
  • Maintenance cost (barrier): Sustaining the carrier state $C$ against thermal fluctuations at temperature $T$ requires an energy barrier proportional to the information stored. Following the Brillouin-Szilard principle [brillouin1956science][bennett1982thermodynamics], maintaining $n$ bits of information requires an energy barrier of at least $n\cdot k_{B}T\ln 2$.
For the lower bounds derived in Theorem 1, we use the total energy budget $E$ as a conservative estimate. A more refined analysis separating erasure and maintenance costs would strengthen these bounds but is not necessary for the inequalities presented here.
Definition 4 (Information).
Information is a six-tuple $I=\langle O,T_{o},S(O,T_{o}),C,T_{r},S(C,T_{r})\rangle$ where:
  • $O\subseteq W\cup M$: ontology (objective or subjective reality)
  • $T_{o}\subseteq\mathbb{R}^{+}$: occurrence time
  • $S(O,T_{o})$: state set of ontology at time $T_{o}$
  • $C\subseteq W$: carrier (objective reality)
  • $T_{r}\subseteq\mathbb{R}^{+}$: reflection time (with $T_{o}\subseteq T_{r}$)
  • $S(C,T_{r})$: state set of carrier at time $T_{r}$
satisfying $I:S(O,T_{o})\rightrightarrows S(C,T_{r})$ is an enabling mapping[xu2024research].
Theorem 1 (Energy-Constrained Representation Capacity).
For any enabling mapping $\mathcal{I}:S(O,T_{o})\rightrightarrows S(C,T_{r})$ requiring energy $E$, the following hold:
  1. 1.
    (Carrier capacity lower bound)
    $\log_{2}|\mathcal{S}(C,T_{r})|\geq\log_{2}|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T\ln 2}.$ (3)
  2. 2.
    (Information preservation bound) Under any joint distribution $P_{O,C}$ consistent with the deterministic mapping $C=f(O)$,
    $I(O;C)\leq\frac{E}{k_{B}T\ln 2},$ (4)
    and consequently,
    $H(O\mid C)\geq H(O)-\frac{E}{k_{B}T\ln 2}.$ (5)
  3. 3.
    (Uniform case—optimal compression) If $\mathcal{S}(O,T_{0})$ is uniformly distributed and the mapping partitions states uniformly (i.e., all fibers have equal size), then equality holds in (3):
    $\log_{2}|\mathcal{S}(C,T_{r})|=\log_{2}|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T\ln 2}.$ (6)
    For non-uniform partitions, inequality (3) is strict.
Proof.
We prove each part in turn.
(1) Carrier capacity lower bound.
By Definition 3, for the surjective mapping $\mathcal{I}:S(O,T_{o})\rightrightarrows S(C,T_{r})$, the energy satisfies
$E\geq k_{B}T\ln\left(\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\right).$ (7)
Rearranging gives
$\ln|\mathcal{S}(C,T_{r})|\geq\ln|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T}.$ (8)
Converting to base-2 logarithms (dividing both sides by $\ln 2$) yields
$\log_{2}|\mathcal{S}(C,T_{r})|\geq\log_{2}|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T\ln 2},$ (9)
which establishes (3).
Physical intuition: Given energy budget $E$, the mapping can erase at most $E/(k_{B}T\ln 2)$ bits of information. Therefore, the carrier must retain at least $\log_{2}|\mathcal{S}(O,T_{0})|-E/(k_{B}T\ln 2)$ bits, which determines the minimum carrier capacity.
(2) Information preservation bound.
Since $C=\mathcal{I}(O)$ is a deterministic function, $H(C\mid O)=0$, and thus
$I(O;C)=H(C)-H(C\mid O)=H(C)\leq\log_{2}|\mathcal{S}(C,T_{r})|.$ (10)
From thermodynamics, the maximum amount of information that can be maintained against thermal noise with energy $E$ is $E/(k_{B}T\ln 2)$ bits. [^1] This follows from the Brillouin-Szilard principle: maintaining information against thermal fluctuations requires an energy barrier proportional to $k_{B}T$ per bit. See Bennett [bennett1982thermodynamics][bennett2003notes] for a detailed thermodynamic analysis of information storage and the fundamental limits imposed by thermal noise. Therefore,
$I(O;C)\leq\frac{E}{k_{B}T\ln 2},$ (11)
proving (4).
Using the identity $H(O\mid C)=H(O)-I(O;C)$, we immediately obtain
$H(O\mid C)\geq H(O)-\frac{E}{k_{B}T\ln 2},$ (12)
which is (5).
(3) Uniform case.
Assume $O$ is uniformly distributed over $\mathcal{S}(O,T_{0})$, so $H(O)=\log_{2}|\mathcal{S}(O,T_{0})|$.
If the mapping $\mathcal{I}$ partitions $\mathcal{S}(O,T_{0})$ into fibers of equal size, then
$|\mathcal{I}^{-1}(c)|=\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\quad\text{for all }c\in\mathcal{S}(C,T_{r}).$ (13)
Under uniform partitioning, $C$ is also uniformly distributed, so $H(C)=\log_{2}|\mathcal{S}(C,T_{r})|$. The information erased is exactly
$\Delta I=H(O)-H(C)=\log_{2}\left(\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\right)\text{ bits}.$ (14)
By Landauer’s principle, the energy consumed is
$E=k_{B}T\ln\left(\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\right).$ (15)
Solving for $\log_{2}|\mathcal{S}(C,T_{r})|$ gives
$\log_{2}|\mathcal{S}(C,T_{r})|=\log_{2}|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T\ln 2},$ (16)
which is the equality (6).
For non-uniform partitions, the distribution of $C$ becomes non-uniform with $H(C)<\log_{2}|\mathcal{S}(C,T_{r})|$. The actual information erased is
$\Delta I=H(O)-H(C)>\log_{2}\left(\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\right),$ (17)
requiring energy
$E>k_{B}T\ln\left(\frac{|\mathcal{S}(O,T_{0})|}{|\mathcal{S}(C,T_{r})|}\right).$ (18)
Consequently, the inequality (3) becomes strict:
$\log_{2}|\mathcal{S}(C,T_{r})|>\log_{2}|\mathcal{S}(O,T_{0})|-\frac{E}{k_{B}T\ln 2}.$ (19)
Physical interpretation: Uniform partitions maximize entropy production efficiency—they erase the minimum amount of information for a given compression ratio $|\mathcal{S}(O)|/|\mathcal{S}(C)|$, making them thermodynamically optimal. Non-uniform partitions waste energy by erasing more information than necessary.
II-C Logical Structure and Atomic Decomposition
Definition 5 (Logical System).
A logical system $\mathcal{L}=(\mathcal{A},\text{WFF},\vdash,\models)$ consists of[qiu2025research][xu2025information]:
  • $\mathcal{A}$: finite set of atomic propositions
  • WFF: well-formed formulas built from $\mathcal{A}$ and connectives $\{\neg,\land,\lor,\to\}$
  • $\vdash$: proof system (e.g., natural deduction, sequent calculus)
  • $\models$: model-theoretic semantics
Definition 6 (Logical Closure).
The logical closure $\text{Cn}(S)$ of a formula set $S\subseteq\text{WFF}$ is:
$\text{Cn}(S)=\{\varphi\in\text{WFF}:S\vdash\varphi\}$ (20)
Definition 7 (Atomic Decomposition).
For state $s$ with interpretation $I(s)\subseteq\text{WFF}$, the atomic decomposition $\text{Atom}(s)$ is the unique minimal set satisfying:
  1. 1.
    $\text{Atom}(s)\subseteq\mathcal{A}$ (atoms only)
  2. 2.
    $\text{Cn}(\text{Atom}(s))=\text{Cn}(I(s))$ (closure equivalence)
  3. 3.
    $\forall A^{\prime}\subsetneq\text{Atom}(s):\text{Cn}(A^{\prime})\subsetneq\text{Cn}(I(s))$ (minimality)
  4. 4.
    $\forall\alpha\in\text{Atom}(s):\text{Atom}(s)\setminus\{\alpha\}\nvdash\alpha$ (independence)
Theorem 2 (Atomic Basis Theorem).
Every finitely expressible state $s$ has a unique atomic decomposition $\text{Atom}(s)$.
Proof.
Existence. Let $\text{Atoms}_{0}=\{\text{all atomic propositions in }I(s)\}$. Define the reduction operator $R$ on subsets of $A$ by
$R(A^{\prime})=\{\alpha\in A^{\prime}:A^{\prime}\setminus\{\alpha\}\not\vdash\alpha\}.$ (21)
We claim that $R$ is idempotent, i.e., $R(R(A^{\prime}))=R(A^{\prime})$.
To see this, let $B=R(A^{\prime})$. Clearly $R(B)\subseteq B$. For the reverse inclusion, let $\alpha\in B$. Then $A^{\prime}\setminus\{\alpha\}\not\vdash\alpha$. Since $B\subseteq A^{\prime}$, we have $B\setminus\{\alpha\}\subseteq A^{\prime}\setminus\{\alpha\}$. By monotonicity of the proof system ($\Gamma\subseteq\Delta$ and $\Gamma\vdash\phi$ implies $\Delta\vdash\phi$), it follows that $B\setminus\{\alpha\}\not\vdash\alpha$. Hence $\alpha\in R(B)$, so $R(B)=B$.
Now set $\text{Atom}(s)=R(\text{Atoms}_{0})$. By construction:
  1. 1.
    $\text{Atom}(s)\subseteq\text{Atoms}_{0}\subseteq A$;
  2. 2.
    For all $\alpha\in\text{Atom}(s)$, $\text{Atom}(s)\setminus\{\alpha\}\not\vdash\alpha$ (independence);
  3. 3.
    $\text{Cn}(\text{Atom}(s))=\text{Cn}(\text{Atoms}_{0})=\text{Cn}(I(s))$, because removing derivable atoms does not change the closure;
  4. 4.
    Minimality follows from independence: any proper subset would miss some non-derivable atom.
Uniqueness. Suppose $\text{Atom}_{1}(s)$ and $\text{Atom}_{2}(s)$ are two minimal atomic decompositions of $s$. Assume for contradiction that $\text{Atom}_{1}(s)\neq\text{Atom}_{2}(s)$. Without loss of generality, let $\alpha\in\text{Atom}_{1}(s)\setminus\text{Atom}_{2}(s)$. Since $\text{Cn}(\text{Atom}_{2}(s))=\text{Cn}(I(s))=\text{Cn}(\text{Atom}_{1}(s))$, we have $\text{Atom}_{2}(s)\vdash\alpha$.
But $\text{Atom}_{1}(s)$ is independent, so $\text{Atom}_{1}(s)\setminus\{\alpha\}\not\vdash\alpha$. However, because $\text{Atom}_{2}(s)\subseteq A$ and $\alpha\in A$ is an atomic proposition, the only way for $\text{Atom}_{2}(s)\vdash\alpha$ to hold in a standard propositional logic is if $\alpha\in\text{Atom}_{2}(s)$ (since atomic propositions are logically independent unless explicitly equated, which is excluded by minimality). This contradicts $\alpha\notin\text{Atom}_{2}(s)$.
Therefore, $\text{Atom}_{1}(s)=\text{Atom}_{2}(s)$. ∎
Definition 8 (Logical Depth).
Fix a sound and complete proof system $\mathcal{P}$ (e.g., natural deduction for first-order logic). The logical depth of formula $\varphi$ conditioned on knowledge state $\psi$ is:
$\text{Ld}(\varphi|\psi)=\min\{n:\exists\pi\in\mathcal{P},\\ \pi\text{ derives }\varphi\text{ from }\text{Atom}(\psi),|\pi|=n\}$ (22)
where $\text{Atom}(\psi)$ is the atomic decomposition of $\psi$ (Definition 7), and $|\pi|$ counts inference rule applications in $\pi$.
Remark 2 (Proof System Dependence).
Logical depth depends on the choice of $\mathcal{P}$, but for complete proof systems in the same logic, depths differ by at most a constant factor. Throughout this paper, we fix $\mathcal{P}$ to be natural deduction for first-order logic with equality. For practical knowledge systems (databases, knowledge graphs), we use simpler fragments (Horn clauses, description logic) where $\mathcal{P}$ reduces to forward chaining or tableau calculus. The key insight is that $\text{Ld}(\varphi|\psi)$ captures the minimal derivation complexity rather than worst-case runtime.
Proposition 1 (Relation to Bennett’s Logical Depth).
Our definition relates to Bennett’s original logical depth [bennett1988logical] as follows: Bennett defines $\text{depth}_{t}(x)$ as the runtime of the fastest program generating string $x$ in time $\leq t$. For knowledge systems where queries $\varphi$ are derived from base facts $\psi$ via logical inference, our $\text{Ld}(\varphi|\psi)$ corresponds to Bennett’s depth with $t$ proportional to proof length. Specifically, if each inference step requires $O(1)$ time, then $\text{depth}_{O(n)}(\varphi)\approx\text{Ld}(\varphi|\psi)$ for minimal proof length $n$.
Proposition 2 (Logical Depth Properties).
For any state $s$ and formulas $\varphi,\psi$:
  1. 1.
    Zero depth: $\text{Ld}_{s}(\varphi)=0\Leftrightarrow\varphi\in\text{Atom}(s)$
  2. 2.
    Monotonicity: If $\text{Atom}(s_{1})\subseteq\text{Atom}(s_{2})$, then $\text{Ld}_{s_{2}}(\varphi)\leq\text{Ld}_{s_{1}}(\varphi)$
  3. 3.
    Subadditivity: $\text{Ld}_{s}(\varphi\land\psi)\leq\text{Ld}_{s}(\varphi)+\text{Ld}_{s}(\psi)+1$
  4. 4.
    Transitivity: $\text{Ld}_{s}(\chi)\leq\text{Ld}_{s}(\varphi)+\text{Ld}_{\{\varphi\}}(\chi)$
Proof.
(1) Direct from definition. (2) More atoms provide more starting points, reducing derivation steps. (3) Derive $\varphi$ and $\psi$ separately, then apply conjunction. (4) Chain derivations: first derive $\varphi$ from $\text{Atom}(s)$, then derive $\chi$ from $\varphi$. ∎
II-D Hybrid Storage-Computation Strategies
Definition 9 (Hybrid Strategy).
A hybrid strategy is a triple $\Pi=(S,C,R)$ where:
  • $S\subseteq\text{Cn}(K)$: precomputed facts stored in memory
  • $C:Q\times 2^{\text{Cn}(K)}\to Q$: computation function deriving answers
  • $R:Q\to\{\text{storage},\text{compute},\text{hybrid}\}$: routing policy
Definition 10 (Strategy Cost).
The cost of strategy $\Pi$ for query distribution $P_{Q}$ is:
$\text{Cost}(\Pi)=$ $w_{E}\cdot E(\Pi)+w_{T}\cdot T(\Pi)+w_{S}\cdot|S|$ (23)
where $E(\Pi)=\mathbb{E}_{q\sim P_{Q}}[E_{\text{query}}(q|\Pi)]$ (24)
$T(\Pi)=\mathbb{E}_{q\sim P_{Q}}[T_{\text{query}}(q|\Pi)]$ (25)
$|S|=\text{storage size in bits}$ (26)
III Logical Depth as an Information Metric
We now establish that logical depth is an information-theoretic quantity dual to Shannon entropy, bridging the gap between logical reasoning and information processing.
III-A Derivation Entropy and Information Content
Definition 11 (Derivation Entropy).
The derivation entropy from $\psi$ to $\varphi$ is defined as:
$H_{\text{derive}}(\varphi|\psi)=\text{Ld}(\varphi|\psi)\cdot\ln(2)$ (27)
measured in nats, where $\ln(2)\approx 0.693$ is the conversion factor from inference steps to natural units of information.
Remark 3.
Derivation entropy quantifies the “information processing work” required to extract $\varphi$ from $\psi$ through logical inference. Each inference step corresponds to processing $\ln(2)$ nats of information, analogous to how each bit erasure in Landauer’s principle requires energy $k_{B}T\ln(2)$.
Assumption 1 (Semantic Information per Inference Step).
In the information-rich regime, each inference step in a typical derivation reduces the conditional entropy $H(q\mid\cdot,K)$ by at least $\Omega(\ln 2)$ nats in expectation. That is, if $\mathcal{D}_{i}$ is the derived knowledge after $i$ steps, then
$\mathbb{E}\big[H(q\mid\mathcal{D}_{i},K)-H(q\mid\mathcal{D}_{i+1},K)\big]=\Omega(\ln 2),$ (28)
where the expectation is taken over queries and goal-directed derivation strategies commonly used in knowledge base reasoning.
Justification. The operational information reflects implementation costs (e.g., encoding derivation traces), while semantic information captures the thermodynamic work of generating new knowledge. In physical terms, Landauer’s principle applies to the erasure of intermediate semantic states, not syntactic indices. Empirically:
  • Modus ponens generates one new fact $\psi$ from $\varphi$ and $\varphi\to\psi$, contributing $\approx\ln 2$ nats of semantic information.
  • Forward chaining over a knowledge graph produces a new triple, whose information content is independent of the graph size.
Thus, while operational costs grow with knowledge base size, semantic information per step remains bounded below by $\ln 2$ nats.
Remark 4 (Conversion Factor Interpretation).
The factor $\ln(2)\approx 0.693$ converts from inference steps (logical complexity) to nats (semantic information). This parallels Landauer’s principle, where $k_{B}T\ln(2)$ converts from bits (information) to joules (thermodynamic cost). Both conversions reflect the fundamental connection between discrete computational operations and continuous physical quantities.
Definition 12 (Information Content).
Let $\varphi$ and $\psi$ be propositional formulas. Define the information content of $\varphi$ given $\psi$ as:
$H(\varphi\mid\psi):=K(\varphi\mid\text{Atom}(\psi))\cdot\ln 2$ (29)
where $K(\varphi\mid\text{Atom}(\psi))$ is the Kolmogorov complexity of $\varphi$ relative to the atom set $\text{Atom}(\psi)$.
Definition 13 (Information-Rich Regime).
A query $q$ given knowledge base $K$ is said to be in the information-rich regime if its conditional information content satisfies:
$H(q\mid K)\gg\log_{2}\big(|\text{Atom}(K)|\big).$
This condition ensures that the semantic information of the query dominates the structural overhead of indexing atoms, making the logarithmic gap in Theorem 3 negligible (Corollary 1).
Theorem 3 (Derivation Entropy Bounds).
For any formulas $\varphi,\psi$ with $\psi\vdash\varphi$, let: $m=|\mathrm{Atom}(\psi)|$, $n=L_{d}(\varphi\mid\psi)$, $c=\max\{\lceil\log_{2}|R|\rceil,\,k\}$, where $|R|$ is the number of inference rules and $k$ bounds the arity of rules. Then the derivation entropy satisfies:
$H_{\mathrm{derive}}(\varphi\mid\psi)\geq\frac{H(\varphi\mid\psi)}{c\cdot\log_{2}(m+n)}-O(\log m),$ (30)
$H_{\mathrm{derive}}(\varphi\mid\psi)\leq\frac{H(\varphi\mid\psi)\cdot c\cdot\log_{2}(m+n)}{\ln 2}+O(\log m).$ (31)
Equivalently,
$H(\varphi\mid\psi)=\Theta\big(H_{\mathrm{derive}}(\varphi\mid\psi)\cdot\log(m+n)\big).$ (32)
In the information-rich regime where $H(\varphi\mid\psi)\gg\log(m+n)$, the logarithmic overhead dominates structural constants, yielding the asymptotic equivalence stated in Corollary 1.
Proof.
Lower bound (30). Any derivation $\pi$ of length $n$ can be encoded by specifying, for each step $i$: (i) the inference rule ($\lceil\log_{2}|R|\rceil$ bits), (ii) the premises ($k\cdot\lceil\log_{2}(m+i)\rceil$ bits). The total length is at most $n\cdot c\cdot\log_{2}(m+n)+O(n)$. By Kolmogorov complexity, $K(\varphi\mid\text{Atom}(\psi))\leq nc\log_{2}(m+n)+O(1)$, where the O(1) term accounts for the universal Turing machine overhead in decoding the derivation. . Using $H(\varphi\mid\psi)\leq K(\varphi\mid\text{Atom}(\psi))\cdot\ln 2+O(\log m)$, we obtain
$H(\varphi\mid\psi)\leq nc\log_{2}(m+n)\cdot\ln 2+O(\log m).$
Rearranging and noting $H_{\mathrm{derive}}=n\ln 2$ gives (30).
Upper bound (31). Conversely, the minimal program that outputs $\varphi$ given $\text{Atom}(\psi)$ can be simulated by a derivation that explicitly constructs each bit of the program. This yields a derivation of length $O(K(\varphi\mid\text{Atom}(\psi)))$, and hence $n=O(K(\varphi\mid\text{Atom}(\psi)))$. Thus $H_{\mathrm{derive}}=n\ln 2=O(H(\varphi\mid\psi))$. To recover the logarithmic factor, note that simulating a Turing machine step may require $O(\log(m+n))$ logical steps to manage state encoding. A tighter analysis (see Li & Vitányi, 1997) gives $n\leq c^{\prime}\cdot K(\varphi\mid\text{Atom}(\psi))\cdot\log(m+n)$, which implies (31). ∎
Remark 5 (Tightness and Physical Interpretation).
The logarithmic gap $\log_{2}(m+n)$ between $H(\varphi|\psi)$ and $H_{\mathrm{derive}}(\varphi|\psi)$ represents:
  1. 1.
    Structural overhead of logical reasoning: Explicit step-by-step derivations must specify rule choices and premise selections, while direct encoding can be more compact.
  2. 2.
    Motivation for storage: Since $H(\varphi|\psi)=O(H_{\mathrm{derive}}(\varphi|\psi)/\log(m+n))$, storing formulas (cost $\sim H(\varphi|\psi)$) is cheaper than deriving them (cost $\sim H_{\mathrm{derive}}(\varphi|\psi)$) by a logarithmic factor.
  3. 3.
    Tightness: For typical formulas with $H(\varphi|\psi)=\Omega(m)$ and $n=O(\mathrm{poly}(m))$, both bounds are tight up to constant factors, confirming that $H_{\mathrm{derive}}$ and $H(\varphi|\psi)$ are equivalent up to logarithmic terms.
Corollary 1 (Asymptotic Equivalence).
For formulas in the information-rich regime where $H(\varphi|\psi)\gg\log(m+n)$, the derivation entropy is asymptotically equivalent to the information content divided by the logarithm of the knowledge size:
$H_{\mathrm{derive}}(\varphi|\psi)=\Theta\left(\frac{H(\varphi|\psi)}{\log(m+n)}\right).$ (33)
Remark 6.
The original condition $H(q\mid K)\gg\log_{2}\big(\max\{|\text{Atom}(K)|,\,L_{d}(q\mid K)\}\big)$ is asymptotically equivalent to Definition 13 for all practical query classes where the logical depth grows at most polynomially with the knowledge base size, i.e., $L_{d}(q\mid K)\leq\mathrm{poly}(|\text{Atom}(K)|)$. This includes Datalog, SPARQL, and template-based natural language queries commonly used in knowledge-intensive AI systems.
III-B Storage-Computation Duality
We now present our central result: the duality between storage (information entropy) and computation (logical depth).
Theorem 4 (Frequency-Weighted Storage-Computation Duality).
Let $K$ be a knowledge base and $q\in Q$ a query in the information-rich regime (Definition 13). For any hybrid strategy $\Pi=(S,C,R)$ with precomputed storage $S\subseteq\mathrm{Cn}(K)$, define:
  • $H_{q}=H(q\mid K)$: semantic information content of $q$,
  • $f_{q}=P_{Q}(q)$: access frequency of query $q$,
  • $m_{S}=|\mathrm{Atom}(S)|$: number of atomic facts in storage,
  • $n_{S}=L_{d}(q\mid S)$: logical depth of deriving $q$ from $S$,
  • $c>0$: system-dependent constant from Theorem 3.
Then the amortized semantic cost per access satisfies:
$\mathrm{Cost}_{\mathrm{amort}}(\Pi;q):=\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S)\\ \geq H_{q}\left(1+\frac{1}{c\log(m_{S}+n_{S})}-\frac{1}{f_{q}}\right)-O(\log m_{S}).$ (34)
Moreover, under the mild assumption that $n_{S}\leq\mathrm{poly}(|\mathrm{Atom}(K)|)$, we have the asymptotic lower bound:
$\mathrm{Cost}_{\mathrm{amort}}(\Pi;q)\geq H_{q}\left(1+\frac{1}{c\log|\mathrm{Atom}(K)|}-\frac{1}{f_{q}}\right)-o(H_{q}).$ (35)
Critical frequency threshold: There exists a critical frequency
$f_{c}=1+\frac{1}{c\log|\mathrm{Atom}(K)|}\approx 1+\frac{1}{\log|\mathrm{Atom}(K)|}$ (36)
such that:
  • For $f_{q}<f_{c}$: Storage is inefficient—the amortized cost (35) exceeds $H_{q}$ (pure storage cost), and computation is preferred.
  • For $f_{q}>f_{c}$: Storage is efficient—the amortized cost falls below $H_{q}$, justifying precomputation.
  • At $f_{q}=f_{c}$: The system is at the break-even point, where storage and computation have equal amortized cost.
This bound is tight up to logarithmic factors: there exist strategies achieving
$\mathrm{Cost}_{\mathrm{amort}}(\Pi;q)\leq H_{q}\left(1+\frac{1}{c\log|\mathrm{Atom}(K)|}+\frac{1}{f_{q}}\right)+o(H_{q}).$ (37)
Proof.
We prove each part in turn.
Part 1: Amortized lower bound (34).
By the chain rule for conditional information (Definition 12),
$H_{q}=H(q\mid K)=H(q\mid S)+I(q;S\mid K),$ (38)
where $I(q;S\mid K)$ is the mutual information between query $q$ and storage $S$ given the knowledge base $K$. Since mutual information is bounded by entropy,
$I(q;S\mid K)\leq|S|.$ (39)
Thus, from (38) and (39),
$|S|\geq H_{q}-H(q\mid S).$ (40)
From Theorem 3 (lower bound),
$H_{\mathrm{derive}}(q\mid S)\geq\frac{H(q\mid S)}{c\log(m_{S}+n_{S})}-O(\log m_{S}).$ (41)
Now consider the amortized cost per access. Since the storage cost $|S|$ is a one-time investment amortized over $f_{q}$ accesses (where $f_{q}=P_{Q}(q)$ is the probability/frequency of query $q$), while the computation cost $H_{\mathrm{derive}}(q\mid S)$ is incurred on every access, the amortized cost per access is:
$\mathrm{Cost}_{\mathrm{amort}}(\Pi;q)=\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S).$ (42)
Substituting (40) and (41) into (42):
$\mathrm{Cost}_{\mathrm{amort}}(\Pi;q)\geq\frac{H_{q}-H(q\mid S)}{f_{q}}+\frac{H(q\mid S)}{c\log(m_{S}+n_{S})}$ (43)
$-O(\log m_{S})$
$=\frac{H_{q}}{f_{q}}+H(q\mid S)\left(\frac{1}{c\log(m_{S}+n_{S})}-\frac{1}{f_{q}}\right)$
$-O(\log m_{S}).$
Since $0\leq H(q\mid S)\leq H_{q}$, to obtain a lower bound, we consider the worst case for the coefficient of $H(q\mid S)$ in (43):
  • If $\frac{1}{c\log(m_{S}+n_{S})}-\frac{1}{f_{q}}\geq 0$, the minimum occurs at $H(q\mid S)=0$, giving $\mathrm{Cost}_{\mathrm{amort}}\geq H_{q}/f_{q}$.
  • If $\frac{1}{c\log(m_{S}+n_{S})}-\frac{1}{f_{q}}<0$, the minimum occurs at $H(q\mid S)=H_{q}$, giving:
$\mathrm{Cost}_{\mathrm{amort}}(\Pi;q)$ $\geq\frac{H_{q}}{f_{q}}+H_{q}\left(\frac{1}{c\log(m_{S}+n_{S})}-\frac{1}{f_{q}}\right)$ (44)
$\quad-O(\log m_{S})$
$=H_{q}\left(\frac{1}{c\log(m_{S}+n_{S})}\right)-O(\log m_{S}).$
Combining both cases and rearranging:
$\mathrm{Cost}_{\mathrm{amort}}(\Pi;q)\geq H_{q}\left(1+\frac{1}{c\log(m_{S}+n_{S})}-\frac{1}{f_{q}}\right)-O(\log m_{S}),$ (45)
which establishes (34).
Part 2: Asymptotic bound and critical frequency.
In the information-rich regime (Definition 13), $H_{q}\gg\log|\mathrm{Atom}(K)|$. Assuming $n_{S}\leq\mathrm{poly}(|\mathrm{Atom}(K)|)$, we have $\log(m_{S}+n_{S})=O(\log|\mathrm{Atom}(K)|)$. Thus, the bound (34) becomes:
$\mathrm{Cost}_{\mathrm{amort}}(\Pi;q)\geq H_{q}\left(1+\frac{1}{c\log|\mathrm{Atom}(K)|}-\frac{1}{f_{q}}\right)-o(H_{q}),$ (46)
which is (35).
The critical frequency $f_{c}$ is the point where the amortized cost equals the pure storage cost $H_{q}$:
$H_{q}\left(1+\frac{1}{c\log|\mathrm{Atom}(K)|}-\frac{1}{f_{c}}\right)=H_{q}\\ \Longrightarrow\quad f_{c}=1+\frac{1}{c\log|\mathrm{Atom}(K)|}.$ (47)
For typical proof systems, $c\approx 1$, giving $f_{c}\approx 1+1/\log|\mathrm{Atom}(K)|$.
Part 3: Tightness.
We construct strategies achieving the upper bound (37):
  1. 1.
    Pure storage ($S\ni q$): Then $H(q\mid S)=0$, $H_{\mathrm{derive}}=0$, and $\mathrm{Cost}_{\mathrm{amort}}=|S|/f_{q}\approx H_{q}/f_{q}$.
  2. 2.
    Hybrid strategy: Store intermediate facts reducing $H(q\mid S)$ to $O(H_{q}/\log|\mathrm{Atom}(K)|)$. Then:
    $\mathrm{Cost}_{\mathrm{amort}}\approx\frac{H_{q}}{f_{q}}+\frac{H_{q}}{c(\log|\mathrm{Atom}(K)|)^{2}}\\ \leq H_{q}\left(\frac{1}{f_{q}}+\frac{1}{c\log|\mathrm{Atom}(K)|}\right)+o(H_{q}),$ (48)
    which matches (37) up to lower-order terms.
Thus, the bound is tight up to logarithmic factors.
Physical interpretation.
The frequency-weighted bound (35) reveals a fundamental trade-off:
  • Low-frequency queries ($f_{q}\ll f_{c}$): The term $H_{q}/f_{q}$ dominates, making amortized cost large. Computation is preferred.
  • High-frequency queries ($f_{q}\gg f_{c}$): The term $H_{q}/c\log|\mathrm{Atom}(K)|$ dominates, making amortized cost small. Storage is preferred.
  • Critical frequency ($f_{q}\approx f_{c}$): The system is at the break-even point, where storage and computation have equal amortized efficiency.
This explains the empirical success of caching policies (LRU, LFU) that prioritize high-frequency items, and provides a rigorous information-theoretic foundation for threshold-based storage selection (Policy 1). ∎
Remark 7 (Frequency Weighting and Amortization).
The critical frequency $f_{c}$ defined in Theorem 4 has several important interpretations:
  1. 1.
    Economic perspective: Storage cost $|S|$ is a “capital investment” amortized over $f_{q}$ accesses, while computation cost $H_{\mathrm{derive}}(q\mid S)$ is an “operational expense” incurred per access. The critical frequency $f_{c}$ is the break-even point where total cost equals pure storage cost.
  2. 2.
    Thermodynamic perspective: By Landauer’s principle (Theorem 1), the amortized energy cost per access is:
    $E_{\mathrm{amort}}(q)=\frac{|S|\cdot k_{B}T\ln 2}{f_{q}}+L_{d}(q\mid S)\cdot k_{B}T\ln 2.$ (49)
    For $f_{q}>f_{c}$, this is less than the energy of direct storage, justifying precomputation from a physical energy perspective.
  3. 3.
    Scaling behavior: As knowledge base size grows ($|\mathrm{Atom}(K)|\to\infty$), the critical frequency approaches 1:
    $f_{c}=1+\frac{1}{c\log|\mathrm{Atom}(K)|}\to 1^{+}.$ (50)
    This means that in very large knowledge systems (e.g., trillion-parameter LLMs, billion-triple knowledge graphs), even moderately frequent queries ($f_{q}>1$) benefit from storage due to the logarithmic compression advantage of logical inference.
  4. 4.
    Connection to classical caching: Classical caching policies (LRU, LFU) implicitly approximate frequency-weighted optimization. However, they ignore logical depth $L_{d}(q\mid K)$, leading to suboptimal decisions. Our threshold policy (Policy 1) corrects this by weighting frequency by computational depth:
    $\text{Store }q\text{ if }f_{q}\cdot L_{d}(q\mid K)>\tau,$ (51)
    where $\tau=\Theta(\log|\mathrm{Atom}(K)|)$ is the scale-aware threshold.
    Remark 8 (Single vs. Multi-Query Cost Accounting).
    Theorem 4 (frequency-weighted duality) addresses the cost of answering a single query $q$ repeatedly over $N_{q}=N\cdot f_{q}$ accesses, where $f_{q}=P_{Q}(q)$ is the access probability. The amortized cost per access is:
    $\text{Cost}_{\text{amort}}(q)=\frac{|S|}{N\cdot f_{q}}+H_{\text{derive}}(q\mid S).$ (52)
    This is the correct metric when optimizing for a specific query $q$ that will be executed multiple times (e.g., a dashboard refresh query in a database).
    For a general-purpose system serving $N$ queries drawn i.i.d. from distribution $P_{Q}$ over query set $\mathcal{Q}$, the expected cost per access is:
    $\mathbb{E}_{q\sim P_{Q}}[\text{Cost}]=\frac{|S|}{N}+\mathbb{E}_{q\sim P_{Q}}[H_{\text{derive}}(q\mid S)].$ (53)
    In the asymptotic regime ($N\to\infty$), the storage term vanishes:
    $\lim_{N\to\infty}\mathbb{E}[\text{Cost}]=\mathbb{E}[H_{\text{derive}}(q\mid S)]\\ =H(Q\mid S,K)=H(Q\mid K)-I(S;Q\mid K),$ (54)
    where the second equality uses $H(Q\mid S,K)=\mathbb{E}_{q\sim P_{Q}}[H(q\mid S,K)]$ and the third uses the standard identity $H(Q\mid S,K)=H(Q\mid K)-I(S;Q\mid K)$ (see Cover & Thomas [cover2012elements], Theorem 2.4.1). Thus, maximizing mutual information $I(S;Q\mid K)$ is equivalent to minimizing expected computation cost.
    Why naive averaging fails. One might incorrectly attempt to “average” the single-query amortized cost (52) over $P_{Q}$:
    $\mathbb{E}_{q\sim P_{Q}}\left[\frac{|S|}{P_{Q}(q)}+H_{\text{derive}}(q\mid S)\right]\stackrel{{\scriptstyle?}}{{=}}\sum_{i=1}^{|\mathcal{Q}|}P_{Q}(q_{i})\\ \cdot\frac{|S|}{P_{Q}(q_{i})}+\mathbb{E}[H_{\text{derive}}]=|S|\cdot|\mathcal{Q}|+\mathbb{E}[H_{\text{derive}}].$ (55)
    This is invalid because it conflates two distinct operational scenarios:
    • In (52), we amortize the one-time storage investment $|S|$ over the $N\cdot f_{q}$ accesses to that specific query.
    • In (53), we amortize $|S|$ over all $N$ accesses across all queries.
    The term $|S|\cdot|\mathcal{Q}|$ in (55) has no physical meaning—it grows linearly with the number of possible queries, whereas storage cost should depend only on actual usage.
    Example III.1 (Numerical verification).
    Consider a query set $\mathcal{Q}=\{q_{1},q_{2},q_{3}\}$ with distribution $P_{Q}=(0.5,0.3,0.2)$, storage $|S|=100$ bits, computation costs $H_{\text{derive}}(q_{i}\mid S)=(10,20,30)$ bits, and $N=1000$ total accesses. Expected computation cost is:
    $\mathbb{E}[H_{\text{derive}}]=0.5\cdot 10+0.3\cdot 20+0.2\cdot 30=17~\text{bits}.$
    Incorrect formula (55):
    $|S|\cdot|\mathcal{Q}|+\mathbb{E}[H_{\text{derive}}]=100\cdot 3+17=317~\text{bits}.$
    Correct formula (53):
    $\frac{|S|}{N}+\mathbb{E}[H_{\text{derive}}]=\frac{100}{1000}+17=17.1~\text{bits}.$
    The $18\times$ discrepancy confirms the invalidity of naive averaging. The correct formula amortizes the storage investment over all accesses, yielding $0.1$ bits/access, whereas the incorrect formula attempts to charge $|S|$ separately for each query type, effectively triple-counting the storage cost.
Practical implication: The critical frequency $f_{c}$ provides a quantitative decision rule for capacity allocation in knowledge systems:
  • LLMs: Parametric storage should be allocated to facts with $f_{q}\cdot L_{d}(q\mid K)>\log N$, where $N$ is the number of parameters. For GPT-4 scale ($N\approx 1.7\times 10^{12}$), this gives $\log N\approx 40$ bits, meaning only queries requiring >40>40 bits of computation at moderate frequency should be stored parametrically.
  • Knowledge graphs: Materialized views should be created for join patterns with $f_{q}>1+1/\log|E|$, where $|E|$ is the number of edges. For Freebase ($|E|\approx 10^{9}$), this gives $f_{c}\approx 1.05$, meaning even slightly frequent multi-hop queries benefit from precomputation.
  • Edge devices: On-device caching should prioritize queries with $f_{q}>1/\eta$, where $\eta$ is the device-to-cloud speed ratio. For 4G networks ($\eta\approx 0.1$), $f_{c}\approx 10$, meaning only queries accessed >10>10 times per session should be cached locally.
Corollary 2 (Frequency-Weighted Extreme Strategy Costs).
Under the conditions of Theorem 3:
  1. 1.
    Pure storage: Store the answer to query $q$. Then
    $\mathrm{Cost}_{\mathrm{amort}}=\frac{|S|}{f_{q}}\approx\frac{H_{q}}{f_{q}},\quad H_{\mathrm{derive}}=0.$ (56)
    This is optimal if and only if $f_{q}\geq f_{c}=1+\frac{1}{c\log|\mathrm{Atom}(K)|}$.
  2. 2.
    Pure computation: Use $S=\emptyset$. Then by Theorem 3,
    $\mathrm{Cost}_{\mathrm{amort}}=H_{\mathrm{derive}}(q\mid K)=\Theta\left(\frac{H_{q}}{\log|\mathrm{Atom}(K)|}\right).$ (57)
    This is optimal if and only if $f_{q}\leq f_{c}$.
  3. 3.
    Optimal hybrid: For queries near the critical frequency ($f_{q}\approx f_{c}$), an optimal hybrid strategy achieves
    $\mathrm{Cost}_{\mathrm{amort}}=H_{q}\left(1-\Theta\left(\frac{1}{\log|\mathrm{Atom}(K)|}\right)\right)+o(H_{q}).$ (58)
    This strategy stores intermediate facts reducing $H(q\mid S)$ to $O(H_{q}/\log|\mathrm{Atom}(K)|)$, balancing storage and computation costs.
  4. 4.
    Frequency stratification: For a query distribution $P_{Q}$, the optimal strategy partitions queries into three classes:
    • High-frequency class ($f_{q}>f_{c}+\epsilon$): Pure storage. Expected cost: $\mathbb{E}[\mathrm{Cost}]\approx\mathbb{E}[H_{q}/f_{q}]$.
    • Medium-frequency class ($|f_{q}-f_{c}|\leq\epsilon$): Hybrid storage-computation. Expected cost: $\mathbb{E}[\mathrm{Cost}]\approx H_{q}(1+o(1))$.
    • Low-frequency class ($f_{q}<f_{c}-\epsilon$): Pure computation. Expected cost: $\mathbb{E}[\mathrm{Cost}]\approx\mathbb{E}[H_{q}/\log|\mathrm{Atom}(K)|]$.
    where $\epsilon=\Theta(1/\log^{2}|\mathrm{Atom}(K)|)$ is the transition width.
Thus, in large knowledge systems ($|\mathrm{Atom}(K)|\to\infty$), the amortized cost of computation for low-frequency queries becomes asymptotically cheaper than storage by a logarithmic factor, while high-frequency queries always benefit from precomputation.
Proof.
Parts (1) and (2) follow directly from Theorem 4 by substituting the extreme cases ($S=\{q\}$ for pure storage, $S=\emptyset$ for pure computation).
For part (3), we construct an explicit hybrid strategy achieving (61). Let $S^{*}\subseteq\text{Atom}(K)$ be the storage set obtained by greedily selecting atoms maximizing mutual information $I(S;q\mid K)$ subject to $|S^{*}|\leq\beta\cdot H_{q}$ for some constant $\beta\in(0,1)$. By submodularity of mutual information (Theorem 6), this achieves
$I(S^{*};q\mid K)\geq\left(1-\frac{1}{e}\right)\cdot H_{q}.$ (59)
By the chain rule, $H(q\mid S^{*},K)=H_{q}-I(S^{*};q\mid K)\leq H_{q}/e$. From Theorem 3 (lower bound),
$H_{\text{derive}}(q\mid S^{*})\geq\frac{H(q\mid S^{*})}{c\log(|S^{*}|+L_{d}(q\mid S^{*}))}\geq\frac{H_{q}}{e\cdot c\log|{\rm Atom}(K)|}.$ (60)
The amortized cost is
$\text{Cost}_{\text{amort}}(q)$ $=\frac{|S^{*}|}{f_{q}}+H_{\text{derive}}(q\mid S^{*})$
$\leq\frac{\beta H_{q}}{f_{q}}+\frac{H_{q}}{e\cdot c\log|{\rm Atom}(K)|}$
$=H_{q}\left(\frac{\beta}{f_{q}}+\frac{1}{e\cdot c\log|{\rm Atom}(K)|}\right).$ (61)
For queries at the critical frequency $f_{q}\approx f_{c}=1+\frac{1}{c\log|{\rm Atom}(K)|}$, choosing $\beta=1/(ef_{c})=\Theta(1/\log|{\rm Atom}(K)|)$ yields
$\text{Cost}_{\text{amort}}(q)=H_{q}\left(1+O\left(\frac{1}{\log|{\rm Atom}(K)|}\right)\right)\\ =H_{q}\left(1-\Theta\left(\frac{1}{\log|{\rm Atom}(K)|}\right)\right)+o(H_{q}),$ (62)
establishing (61).
For part (4), the frequency stratification follows by comparing the costs in parts (1)-(3): pure storage is optimal when $H_{q}/f_{q}<H_{q}/\log|{\rm Atom}(K)|$ (i.e., $f_{q}>\log|{\rm Atom}(K)|$), pure computation when $f_{q}\ll 1$, and hybrid strategies dominate in the intermediate regime $f_{q}\approx f_{c}$ where both costs are comparable. ∎
Remark 9 (Relationship to Original Formulation).
The frequency-weighted Theorem 4 generalizes the original storage-computation duality bound as follows:
Original bound (without frequency):
$\mathrm{Cost}_{\mathrm{sem}}(\Pi;q):=|S|+H_{\mathrm{derive}}(q\mid S)\\ \geq H_{q}\left(1+\frac{1}{c\log|\mathrm{Atom}(K)|}\right)-o(H_{q}).$ (63)
Frequency-weighted bound (Theorem 4):
$\mathrm{Cost}_{\mathrm{amort}}(\Pi;q):=\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S)\\ \geq H_{q}\left(1+\frac{1}{c\log|\mathrm{Atom}(K)|}-\frac{1}{f_{q}}\right)-o(H_{q}).$ (64)
Key differences:
  1. 1.
    Cost amortization: The frequency-weighted bound amortizes the one-time storage cost $|S|$ over $f_{q}$ accesses, reflecting the economic reality that precomputation is a capital investment with diminishing marginal returns.
  2. 2.
    Critical frequency: The original bound is universal (holds for all queries), while the frequency-weighted bound reveals a phase transition at $f_{c}=1+1/(c\log|\mathrm{Atom}(K)|)$ separating storage-efficient from computation-efficient queries.
  3. 3.
    Single-query vs. multi-query: The original bound optimizes for a single query execution, while the frequency-weighted bound optimizes for repeated query executions under a distribution $P_{Q}$. This is critical for practical systems where the same query may be issued multiple times (e.g., “What is the capital of France?” in a chatbot).
  4. 4.
    Tightness: Both bounds are tight up to logarithmic factors, but achieve tightness under different conditions:
    • Original: Tight for $f_{q}=1$ (single execution).
    • Frequency-weighted: Tight for $f_{q}\approx f_{c}$ (repeated executions near break-even).
Unified formulation: Both bounds can be unified by considering the expected total cost over $N$ accesses:
$\mathbb{E}[\mathrm{Total\ Cost}]=|S|+N\cdot f_{q}\cdot H_{\mathrm{derive}}(q\mid S),$ (65)
where $N\cdot f_{q}$ is the expected number of times query $q$ is issued. Dividing by $N\cdot f_{q}$ yields the amortized cost per access:
$\mathbb{E}[\mathrm{Cost\ per\ Access}]=\frac{|S|}{N\cdot f_{q}}+H_{\mathrm{derive}}(q\mid S)\\ \approx\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S)\quad\text{(for large $N$)}.$ (66)
Thus, the frequency-weighted bound is the asymptotic limit of the original bound under repeated accesses, making it the correct metric for optimizing long-running knowledge systems (databases, LLMs, knowledge graphs) rather than one-shot query answering.
III-C Mutual Information and Submodularity
Theorem 5 (Storage Efficiency via Mutual Information).
The efficiency of storage set $S$ for query distribution $P_{Q}$ is:
$\eta(S)=\frac{I(S;Q)}{|S|}\in[0,1]$ (67)
with equality at 1 if and only if $S$ is a sufficient statistic for $Q$.
Theorem 6 (Submodularity of Mutual Information).
The mutual information $I(S;Q)$ is submodular in $S$:
$I(S_{1}\cup S_{2};Q)+I(S_{1}\cap S_{2};Q)\leq I(S_{1};Q)+I(S_{2};Q)$ (68)
Proof.
This follows from the submodularity of entropy. For any sets $A,B$:
$H(A\cup B)+H(A\cap B)$ $\leq H(A)+H(B)$ (69)
$\Rightarrow I(A\cup B;Q)+I(A\cap B;Q)$ $\leq I(A;Q)+I(B;Q)$ (70)
by expanding mutual information as I(X;Q)=H(Q)H(Q|X)I(X;Q)=H(Q)-H(Q|X) and using the fact that conditional entropy is supermodular. ∎
Corollary 3 (Greedy Approximation).
A greedy algorithm for selecting storage items achieves a (11/e)(1-1/e) approximation ratio for maximizing $I(S;Q)$ subject to $|S|\leq M_{\max}$.
IV Thermodynamic Constraints and Physical Limits
This section establishes the physical constraints on information processing in knowledge systems, connecting our information-theoretic framework to thermodynamics via Landauer’s principle.
IV-A Landauer’s Principle and Energy Bounds
Theorem 7 (Landauer Energy Bound for Computation).
Deriving formula $\varphi$ from knowledge base $K$ requires minimum energy:
$E_{\text{compute}}(\varphi|K)\geq\text{Ld}(\varphi|K)\cdot k_{B}T\ln(2)$ (71)
where $k_{B}=1.38\times 10^{-23}$ J/K is Boltzmann’s constant and $T$ is the operating temperature.
Proof.
Each inference step in deriving $\varphi$ from $K$ performs an irreversible logical operation. By Landauer’s principle [landauer1961irreversibility], each irreversible operation erasing or transforming 1 bit of information requires minimum energy $k_{B}T\ln(2)$.
Consider a derivation with $\text{Ld}(\varphi|K)$ steps. Each step:
  1. 1.
    Applies an inference rule $r_{i}:\{\psi_{1},\ldots,\psi_{k}\}\vdash\psi^{\prime}$
  2. 2.
    Writes conclusion $\psi^{\prime}$ to working memory
  3. 3.
    Erases intermediate results (at least 1 bit per step)
The total bit erasures across all steps is at least $\text{Ld}(\varphi|K)$. By Landauer’s principle, total energy:
$E_{\text{compute}}\geq\text{Ld}(\varphi|K)\cdot k_{B}T\ln(2)$ (72)
Theorem 8 (Storage Energy Bound).
Maintaining storage $S$ containing $|S|$ bits for time $t$ requires energy:
$E_{\text{storage}}(S,t)\geq|S|\cdot k_{B}T\ln(2)\cdot\frac{t}{t_{\text{refresh}}}$ (73)
where $t_{\text{refresh}}$ is the thermal decoherence time.
Proof.
Physical storage media suffer thermal noise causing bit flips at rate $\lambda_{\text{flip}}\sim\exp(-E_{\text{barrier}}/k_{B}T)$. The refresh period is $t_{\text{refresh}}=1/\lambda_{\text{flip}}$. Each refresh cycle requires re-writing $|S|$ bits, each costing $k_{B}T\ln(2)$ energy. Over time $t$, number of refreshes is $\lceil t/t_{\text{refresh}}\rceil$, yielding the stated bound. ∎
IV-B Energy-Time-Space Triality Bound
We now prove our main physical constraint theorem.
Theorem 9 (Fundamental Triality Bound).
For any hybrid strategy $\Pi=(S,C,R)$ answering query distribution $P_{Q}$ over knowledge base $K$ in the information-rich regime:
$E(\Pi)\cdot T(\Pi)\cdot M^{-1}(\Pi)\geq\Omega(H(Q|K))\cdot k_{B}T\ln(2)$ (74)
where $\Omega(H(Q|K))$ represents the computational cost induced by the conditional entropy $H(Q|K)$, with dimension of time. Via Landauer’s principle, both sides can be dimensionally reduced to time, ensuring consistency.
This bound is asymptotically tight and achieved by the optimal hybrid strategy in the information-rich regime (see Section VI).
Proof.
[Preamble] Since queries are in the information-rich regime:
$H(Q|K)\gg\log_{2}(|\text{Atom}(K)|),\log_{2}(Ld(Q|K))$ (75)
By Corollary 1, this implies:
$Ld(Q|K)\cdot\ln(2)\approx H(Q|K)\quad[\text{asymptotically}]$ (76)
We analyze three cases using this approximation.
Case 1: Pure Computation Strategy ($\Pi_{c}$)
Setup: - No precomputed storage: $S=\emptyset$, hence $M(\Pi_{c})\to 0$.
- Time cost: Average number of inference steps:
$T(\Pi_{c})=\mathbb{E}_{q\sim P_{Q}}[Ld(q|K)]=Ld(Q|K)$ (77)
- Energy cost: By Theorem 7,
$E(\Pi_{c})=\mathbb{E}_{q\sim P_{Q}}[E_{\text{compute}}(q|K)]\geq Ld(Q|K)\cdot k_{B}T\ln(2)$ (78)
Energy-Time Product: Since we are in the information-rich regime, we use the asymptotic approximation:
$Ld(Q|K)\cdot\ln(2)\approx H(Q|K)$ (79)
Thus,
$\begin{split}E(\Pi_{c})\cdot T(\Pi_{c})&\geq(Ld(Q|K))^{2}\cdot k_{B}T\ln(2)\\ &\geq\Omega(H(Q|K)^{2}/\ln(2))\cdot k_{B}T\ln(2)\\ &=\Omega(H(Q|K)^{2}\cdot k_{B}T)\end{split}$ (80)
Case 2: Pure Storage Strategy ($\Pi_{s}$)
Setup: - Precompute all answers: $S=\{(q,A(q,K)):q\in\text{supp}(P_{Q})\}$. - Storage requirement:
$M(\Pi_{s})=H(Q|K)$ (81)
- Maintenance energy: By Theorem 9, over average query inter-arrival time $t_{\text{avg}}=1/\lambda_{Q}$:
$E(\Pi_{s})=H(Q|K)\cdot k_{B}T\ln(2)\cdot t_{\text{avg}}\cdot t_{\text{refresh}}$ (82)
Energy-Storage Product:
$E(\Pi_{s})\cdot M^{-1}(\Pi_{s})=k_{B}T\ln(2)\cdot t_{\text{avg}}\cdot t_{\text{refresh}}$ (83)
For reasonable systems where $t_{\text{avg}}\sim t_{\text{refresh}}$:
$E(\Pi_{s})\cdot M^{-1}(\Pi_{s})\geq\Omega(k_{B}T\ln(2))$ (84)
In the information-rich regime, this bound holds as stated.
Case 3: Hybrid Strategy ($\Pi_{h}$)
Setup: - We consider a hybrid strategy that balances both precomputation and on-demand computation.
Analysis: By the Storage-Computation Duality Theorem (Theorem 4):
$\min[\text{Storage}]+\min[\text{Computation}]\geq H(Q|K)\cdot(1+\ln(2))$ (85)
In the information-rich regime, the total cost can be approximated as:
$E(\Pi_{h})\cdot T(\Pi_{h})\cdot M^{-1}(\Pi_{h})\geq\Omega(H(Q|K)\cdot k_{B}T\ln(2))$ (86)
This completes the analysis for all three cases, showing consistency with the triality bound.
Corollary 4 (Critical Storage Capacity).
The critical storage capacity is:
$M_{\text{critical}}=\frac{H(Q)}{\log\left(\frac{E_{\text{budget}}}{H(Q|K)\cdot k_{B}T\ln(2)}\right)}$ (87)
Below this capacity, storage is the bottleneck; above it, computation dominates.
IV-C Entropy Production and Irreversibility
Theorem 10 (Minimum Entropy Production).
Answering query $q$ produces minimum environmental entropy:
$\Delta S_{\text{env}}\geq\frac{I(S;q)}{T}$ (88)
Proof.
By the second law of thermodynamics, total entropy change:
$\Delta S_{\text{total}}=\Delta S_{\text{system}}+\Delta S_{\text{env}}\geq 0$ (89)
The system extracts information $I(S;q)$ from storage $S$. This reduces system entropy by $I(S;q)/\ln(2)$ nats. To maintain non-negative total entropy:
$\Delta S_{\text{env}}\geq\frac{I(S;q)}{T}$ (90)
V Optimal Hybrid Strategies
V-A Threshold Policies for Hybrid Strategies
Theorem 4 implies that the marginal benefit of storing additional information diminishes as the knowledge base grows. For a query $q$ with access frequency $f_{q}$, the expected cost of computing on-demand is
$\mathbb{E}[\text{Compute Cost}]=f_{q}\cdot H_{\mathrm{derive}}(q\mid K)\\ =\Theta\!\left(f_{q}\cdot\frac{H_{q}}{\log|\text{Atom}(K)|}\right),$ (91)
while the amortized cost of storing its answer is
$\mathbb{E}[\text{Store Cost}]=H_{q}+\lambda\cdot\text{(storage overhead)},$ (92)
where $\lambda$ penalizes storage capacity usage.
A query should be stored if and only if
$f_{q}\cdot\frac{H_{q}}{\log|\text{Atom}(K)|}>H_{q}\quad\Longleftrightarrow\quad f_{q}>\Theta(\log|\text{Atom}(K)|).$ (93)
Thus, we obtain a scale-aware threshold policy:
Policy 1 (Scale-Aware Threshold).
Store query $q$ if its normalized frequency satisfies
$f_{q}\cdot\frac{L_{d}(q\mid K)}{H_{q}}>\tau,$ (94)
where the threshold $\tau=\Theta(\log|\text{Atom}(K)|)$ grows logarithmically with knowledge base size.
This policy generalizes classical caching (e.g., LFU) by incorporating computational depth $L_{d}(q|K)$ and semantic information $H_{q}$, and explains the phase transition in storage utility observed empirically.
V-B Phase Transition
Theorem 11 (Storage Utility Phase Transition).
Define $\alpha=E_{\text{storage}}/E_{\text{compute}}$. There exists critical point
$\alpha_{c}=\frac{H(Q)}{\mathbb{E}[\text{Ld}(Q|K)]\cdot\ln(2)}$ (95)
such that: (i) for $\alpha<\alpha_{c}$, $\partial L/\partial M\sim-\exp(-\alpha/\alpha_{c})$; (ii) for $\alpha>\alpha_{c}$, $\partial L/\partial M\sim-1/M$.
VI Numerical Validation and Simulations
All datasets satisfy $H(q\mid K)\gg\log|\mathrm{Atom}(K)|$ (verified empirically), placing them firmly in the information-rich regime. We validate our theoretical predictions through extensive numerical experiments on synthetic knowledge systems. Knowledge bases range from 1K to 100K atoms. Query distributions include uniform and Zipfian ($\alpha\in\{1.0,1.2,1.5,1.8,2.0\}$). Physical parameters are: temperature T=300T=300K.
In Experiment 1, we validate Theorem 4 (Frequency-Weighted Duality) by testing individual queries sampled from realistic distributions. For each query $q$, we compute its access frequency $f_{q}$ from a Zipfian distribution, then verify that the optimal amortized cost satisfies the theoretical bound $\text{Cost}_{\text{amort}}(q)\geq H_{q}(1+\frac{1}{c\log|K|}-\frac{1}{f_{q}})$. We test 50 randomly sampled queries from each of three knowledge bases: Small (1K atoms), Medium (10K atoms), Large (100K atoms), with average logical depths 3–5.
TABLE I: Single-Query Duality Theorem Validation (Theorem 4)
KB Size Entities Queries Tested Satisfaction Rate Avg Margin
Small 1.00K 50 100.0% 6785.725
Medium 10.00K 50 100.0% 280828.276
Large 100.00K 50 100.0% 10329841.734
Note: Satisfaction Rate = fraction of queries satisfying $\text{Cost}_{\text{amort}}(q)\geq H_{q}(1+\frac{1}{c\log|K|}-\frac{1}{f_{q}})$. Avg Margin = average difference between observed cost and theoretical bound.
The results demonstrate perfect validation of Theorem 4: all 150 tested queries (50 per knowledge base) satisfy the frequency-weighted duality bound. The large positive margins (6.8K–10.3M nats) indicate that the bound is not tight in practice, which is expected since the theorem provides a fundamental lower limit rather than an achievable optimum. The increasing margin with knowledge base size reflects the $1/f_{q}$ term in the bound—larger knowledge bases have more queries with lower individual frequencies, leading to more conservative bounds.
[…]
Physical Interpretation: At room temperature with $H(Q|K)=10^{6}$ bits, the fundamental limit is:
$E\cdot T\cdot M^{-1}\geq 10^{6}\times 1.38\times 10^{-23}\times 0.693\approx 10^{-17}\text{ J}\cdot\text{s/bit}$ (96)
In Experiment 2, we vary storage capacity from 0
Figure 1: Phase transition in query latency. The figure shows: (top-left) latency vs. storage capacity, (top-right) gradient $\partial L/\partial M$ marking the phase transition signature, (bottom-left) cache hit rate, and (bottom-right) latency vs. storage in bits on log scale. The red dashed vertical lines mark the transition point at approximately 10% storage capacity, where latency improvement shifts from rapid to gradual, demonstrating the phase transition phenomenon predicted by theory.
Results (Figure 1): The phase transition occurs at approximately 10% of the query set, identified by the inflection point in the latency gradient. Below this point, increasing storage yields rapid latency improvement (steep gradient); above it, further storage provides diminishing returns (gradual gradient). The transition validates the theoretical prediction that storage utility undergoes a qualitative shift at a critical capacity determined by information-theoretic and workload parameters.
In Experiment 3, we compare our FreqDepth strategy ($\text{score}=\text{freq}\times\text{depth}$) against three baselines across cache sizes 10-500 on the medium KB (1000 queries). We focus on the constrained regime (cache=50, 5% of queries) where strategy differences are most pronounced.
TABLE II: Baseline Comparison at Cache Size = 50 (5% of queries)
Strategy Hit Rate Latency Improv.
LRU 87.2% 1.54 baseline
LFU 87.9% 1.50 2.6%
TrueMI 51.6% 2.96 -92%
FreqDepth (Ours) 91.5% 1.36 11.6%
Figure 2: Baseline strategy comparison across cache sizes (10-500). The three panels show: (left) cache hit rate, (center) average query latency, and (right) total computation cost. Our FreqDepth strategy (freq $\times$ depth heuristic) consistently outperforms LRU and LFU across all metrics and cache sizes. TrueMI (mutual information maximization) performs poorly due to ignoring access frequency.
At the constrained cache size of 50 (Table 2), our FreqDepth strategy achieves 91.5% hit rate vs. 87.2% for LRU, with 11.6% latency improvement. Figure 2 shows performance across the full range (cache sizes 10-500): FreqDepth consistently outperforms LRU/LFU at all sizes. The theoretically optimal TrueMI performs poorly (51.6% hit rate at cache=50) because it greedily selects queries that maximize global mutual information without considering access frequency—it caches informationally rich but rarely accessed queries. In contrast, FreqDepth’s $\text{freq}\times\text{depth}$ heuristic balances information gain with access probability, achieving practical superiority. At larger cache sizes (e.g., cache $\geq$ 300), all classical strategies converge to $95\%$ hit rates as capacity constraints relax.
In Experiment 4, we test robustness via sensitivity analysis across: (i) five Zipf exponents $\alpha\in\{1.0,1.2,1.5,1.8,2.0\}$ modeling query frequency distributions; (ii) five average logical depths $\{2,3,5,7,10\}$; (iii) four knowledge base scales from 100 to 5000 entities:
Figure 3: Sensitivity analysis. Left pair: cost vs. storage fraction for different Zipf parameters $\alpha$ and optimal $\beta^{*}$ vs. $\alpha$. Right pair: cost vs. storage fraction for different logical depths and optimal $\beta^{*}$ vs. depth. Theoretical bounds hold across all configurations.
[…]
Figure 4 shows that $H(Q)$ scales logarithmically with knowledge base size as predicted by information theory. The framework’s bounds hold consistently across all tested configurations, validating theoretical predictions across diverse parameter regimes.
VII Applications
We demonstrate practical applications of our framework across three domains.
VII-A Large Language Models
Consider a GPT-4 scale model with 1.7T parameters. The knowledge base contains $H(K)=10^{15}$ bits, and queries follow a power-law distribution with $H(Q)=10^{12}$ bits.
Parameter Settings. Average logical depth: $\mathbb{E}[\text{Ld}]=5$ inference steps. Energy budget: $E_{\text{budget}}=1000$ kWh. Temperature: T=300T=300K.
Critical Storage Calculation. Using Corollary 4:
$M_{\text{critical}}=\frac{10^{12}}{\log(1000\times 3.6\times 10^{9}/(10^{6}\times k_{B}T\ln 2))}\approx 2.5\times 10^{11}\text{ bits}$ (97)
Optimal Strategy. Following our threshold policy framework, we compute $\text{Ld}^{*}=2.5$ and $f^{*}=0.85$. The optimal strategy:
  • Parametric storage (70%): Facts with $\text{Ld}\leq 2$ and frequency >0.85>0.85
  • RAG retrieval (25%): Facts with $3\leq\text{Ld}\leq 8$
  • Dynamic reasoning (5%): Complex queries with $\text{Ld}>8$
Performance Gains. Compared to pure parametric storage: 45% energy reduction, with latency distribution P50=50ms, P90=200ms, P99=2000ms.
VII-B Knowledge Graphs
A medical knowledge graph with $10^{7}$ entities and $10^{8}$ relationships faces multi-hop query workloads.
Query Analysis.
Query Type Ld Frequency Example
Single-hop 1 40% Drug $\to$ Disease
Double-hop 2 35% Symptom $\to$ Treatment
Triple-hop 3 20% Patient $\to$ Protocol
Complex $\geq 4$ 5% Multi-condition
[…]
VII-C Edge Intelligence
A smartphone assistant with 4GB RAM and 15Wh battery must operate for 24 hours.
Resource Allocation. Energy budget per 24h: $E_{\text{budget}}=4.5$ Wh (30% to AI). Storage limit: $M_{\max}=2$ GB (50% to knowledge).
Hybrid Strategy. Device storage (200 MB): Common sense, personal habits, quick facts. Cloud coordination: Low-frequency knowledge and complex reasoning ($\text{Ld}>10$). Adaptive caching: Weekly updates based on usage patterns.
Energy Distribution. Device inference (40%), network communication (30%), storage maintenance (20%), overhead (10%). Achieves 95% on-device query completion within 24h battery constraint.
VIII Related Work
VIII-A Information Theory and Kolmogorov Complexity
Shannon’s seminal 1948 paper [shannon1948mathematical] introduced entropy $H(X)=-\sum_{x}p(x)\log p(x)$ as the fundamental measure of information content, establishing the mathematical foundations for data compression and communication. This framework has been extensively developed, with Cover and Thomas [cover2012elements] providing the standard modern treatment encompassing source coding, channel capacity, and rate-distortion theory.
Parallel to Shannon’s probabilistic approach, algorithmic information theory developed the notion of descriptional complexity. Kolmogorov [kolmogorov1965three], Solomonoff [solomonoff1964formal], and Chaitin [chaitin1969simplicity] independently introduced Kolmogorov complexity $K(x)$ as the length of the shortest program generating string $x$. This captures a different facet of information: while Shannon entropy measures average information content over a distribution, Kolmogorov complexity measures the individual information content of a single object.
The minimum description length (MDL) principle [rissanen1978modeling, grunwald2007minimum] applies these ideas to model selection and inference, balancing model complexity against data fit. However, MDL focuses on static description length, not the dynamic computational cost of deriving consequences.
Bennett’s logical depth [bennett1988logical] addresses this gap by measuring computational sophistication: $\text{Ld}(x)$ is the runtime of the fastest program generating $x$ from a short input. This distinguishes between ”shallow” information (random bits, easy to describe but not to compute from scratch) and ”deep” information (organized structures requiring significant computation to generate).
Our work extends Bennett’s framework in three critical ways:
  1. 1.
    Conditional depth: We formalize $\text{Ld}(\varphi|\psi)$ for conditional derivation, enabling analysis of incremental inference.
  2. 2.
    Information-theoretic duality: We prove that derivation entropy $H_{\text{derive}}=\text{Ld}\cdot\ln(2)$ is dual to Shannon entropy, establishing a precise quantitative relationship.
  3. 3.
    Physical grounding: We connect logical depth to thermodynamic energy costs via Landauer’s principle, providing physical constraints on computation.
Related notions include computational depth [antunes2006sophistication], which measures the difficulty of generating an object from random bits, and logical depth sophistication [koppel1987structure], which captures ”useful” information. However, these works focus on individual objects rather than knowledge systems, and lack our connection to storage-computation trade-offs and thermodynamic limits.
VIII-B Thermodynamics of Computation
Landauer’s landmark 1961 paper [landauer1961irreversibility] established the first deep connection between information theory and thermodynamics, proving that erasing one bit of information requires dissipating minimum energy $E_{\text{erase}}\geq k_{B}T\ln(2)$ where $k_{B}$ is Boltzmann’s constant and $T$ is temperature. This result has profound implications: computation is fundamentally irreversible and energy-consuming at the physical level.
Bennett extended this foundation in several directions. His work on reversible computation [bennett1973logical] showed that logically reversible operations can be performed with arbitrarily small energy dissipation, though at the cost of increased time and space. His comprehensive 1982 review [bennett1982thermodynamics] synthesized the thermodynamics of computation, connecting Landauer’s principle to Maxwell’s demon, the second law, and computational complexity.
Recent years have seen experimental confirmation of Landauer’s principle. Bérut et al. [berut2012experimental] demonstrated bit erasure approaching the Landauer limit using a colloidal particle in optical tweezers. Theoretical refinements [sagawa2010thermodynamic] have connected information thermodynamics to feedback control and Maxwell’s demon.
Lloyd [lloyd2000ultimate] derived ultimate physical limits to computation from quantum mechanics, showing that a system of mass $m$ and energy $E$ confined to volume $V$ can perform at most $E/(h\pi)$ operations per second and store at most $S=k_{B}\ln d$ bits where $d$ is the dimensionality of the Hilbert space.
Our work makes three novel contributions beyond existing thermodynamics:
  1. 1.
    Logical depth-energy connection: We prove that each inference step requires energy $k_{B}T\ln(2)$ by connecting logical depth to bit erasure.
  2. 2.
    Triality bound: We derive the first joint bound on energy, time, and storage: $E\cdot T\cdot M^{-1}\geq\Omega(H(Q|K))\cdot k_{B}T\ln(2)$, where $\Omega(H(Q|K))$ represents computational cost with time dimension. Previous work bounded resources independently; our triality captures fundamental coupling.
  3. 3.
    System-level analysis: Rather than isolated gates or operations, we analyze end-to-end knowledge system workflows, accounting for storage maintenance, query routing, and dynamic workloads.
This shift from gate-level to system-level thermodynamics enables practical application to modern computing systems (LLMs, knowledge graphs, edge devices) where architectural decisions dominate energy consumption.
VIII-C Time-Space Trade-offs in Computation
Classical computational complexity theory has extensively studied time-space trade-offs. Yao [yao1985should] pioneered this area with results on sorting and searching. Beame et al. [beame1991general] provided general sequential time-space trade-offs showing that for many problems, $T(n)\cdot S(n)=\Omega(n^{2})$ where $T$ is time and $S$ is space. Savage’s work on data structures [arora2009computational] established fundamental limits on pointer-based representations.
Competitive analysis of caching. Sleator and Tarjan’s seminal work on competitive paging algorithms established that LRU achieves $O(k)$-competitive ratio where $k$ is cache size. However, competitive analysis focuses on worst-case performance against adversarial inputs, whereas our information-theoretic framework characterizes average-case performance under probabilistic query distributions. Recent work on learning-augmented algorithms [kraska2018case] attempts to bridge this gap using machine learning, but lacks theoretical guarantees.
Our work differs by: (i) adopting an information-theoretic perspective quantifying information content rather than worst-case complexity, (ii) incorporating thermodynamic energy costs via Landauer’s principle, (iii) optimizing average-case performance over query distributions using logical depth, and (iv) providing asymptotic bounds rather than competitive ratios. The constant $(1+\ln(2))$ emerges from information-theoretic first principles rather than algorithm-specific analysis.
VIII-D Knowledge Base and Caching Systems
Materialized Views. Database systems extensively use materialized views to trade storage for query latency [gupta1997materialized, chaudhuri1995optimizing]. These systems employ cost-based optimizers but lack theoretical guarantees. Our framework provides the first information-theoretic lower bounds on this trade-off.
Caching Policies. Classical caching algorithms like LRU, LFU, and ARC [megiddo2003arc] rely on heuristics based on recency and frequency. While effective, they provide no optimality guarantees. Our threshold strategy achieves provable (11/e)(1-1/e) approximation via submodular optimization [nemhauser1978analysis, krause2014submodular].
Learned Systems. Recent work applies machine learning to database optimization [kraska2018case], exploiting data patterns. These approaches are complementary to our theoretical framework, which provides fundamental limits that learned systems cannot surpass.
VIII-E Knowledge Graphs and Semantic Systems
Large-scale knowledge graphs [bollacker2008freebase, carlson2010toward] and semantic networks [miller1995wordnet] face significant storage-computation challenges. Embedding methods [bordes2013translating] compress knowledge at the cost of inexact retrieval. Our framework quantifies this trade-off: compression reduces storage entropy but increases logical depth for reconstruction.
Modern question-answering systems combine parametric knowledge (stored in model weights) with non-parametric retrieval [karpukhin2020dense, lewis2020retrieval]. This hybrid approach implicitly navigates the storage-computation trade-off. Our theoretical analysis provides principled guidelines for allocating capacity between these components.
VIII-F Large Language Models and Retrieval
Transformer-based language models [vaswani2017attention, brown2020language] store knowledge parametrically, requiring massive storage and computation. Retrieval-augmented generation (RAG) [lewis2020retrieval] reduces this burden by offloading factual knowledge to external databases.
Our analysis reveals that optimal LLM architectures should store high-frequency, high-logical-depth knowledge parametrically, while retrieving low-frequency or simple facts. This explains the empirical success of RAG: it exploits the storage-computation duality by allocating expensive parametric storage to knowledge with high $\text{Ld}(q|K)\times\text{freq}(q)$ product.
Recent systems like ZeRO [rajbhandari2020zero] and distributed training frameworks [zhao2020pytorch] address memory limitations in large models. Our triality bound provides theoretical justification for their design choices.
VIII-G Energy-Efficient Computing
The energy crisis in computing [horowitz20141, barroso2007case] has motivated extensive work on energy-proportional systems. Prior work optimizes specific components (processors, memory, networks) independently. Our triality bound reveals fundamental coupling between energy, time, and storage that cannot be overcome through component-level optimization alone.
Sub-threshold computing [lei2017optimal] reduces energy per operation but increases latency. Our framework quantifies this trade-off through the triality product, showing that gains in energy efficiency must be balanced against increased time or storage costs.
VIII-H Phase Transitions in Computer Science
Phase transitions in computational complexity [monasson1999determining] and constraint satisfaction [krzakala2007gibbs] reveal sudden changes in problem difficulty. Our phase transition at critical storage capacity $M_{c}$ is analogous but fundamentally different: it arises from information-theoretic constraints rather than problem structure.
The transition we observe is sharp and universal across problem domains, determined solely by $H(Q)$, $\mathbb{E}[\text{Ld}]$, and physical constants. This universality distinguishes it from problem-specific phase transitions in SAT solvers or random graphs.
VIII-I Information-Theoretic Learning and Prediction
Information-theoretic approaches to learning have a rich history. Merhav and Feder [merhav1998universal] developed universal prediction schemes achieving min-max regret bounds under entropy constraints. Their work focuses on passive learning where the learner observes data and predicts. In contrast, our framework addresses active computation where the system decides whether to retrieve precomputed results or compute on-demand.
The connection arises through regret minimization: our storage selection problem is equivalent to minimizing regret in an online learning game where ”experts” are candidate storage sets. However, standard online learning assumes unit costs per expert consultation, whereas in our setting, computation cost scales with logical depth $\text{Ld}(q|S)$, requiring depth-aware algorithms.
VIII-J Submodularity and Greedy Algorithms
Submodular optimization [nemhauser1978analysis, krause2014submodular, buchbinder2018submodular] provides approximation guarantees for many combinatorial problems. We prove that mutual information $I(S;Q)$ is submodular in storage set $S$, enabling (11/e)(1-1/e) approximation via greedy selection [badanidiyuru2014fast].
This connection is non-trivial: while entropy is submodular, the optimal storage set must balance multiple objectives (coverage, diversity, computation reduction), requiring careful analysis. Recent advances in submodular maximization [buchbinder2018submodular] could potentially improve our greedy algorithms, though the (11/e)(1-1/e) bound is optimal for general submodular functions.
VIII-K Online Learning and Adaptive Strategies
For dynamic query distributions, we develop online algorithms achieving $O(\sqrt{T\log n})$ regret. This connects to online convex optimization [hazan2016introduction] and multi-armed bandits [auer2002nonstochastic, cesa2006prediction].
Unlike classical online learning where actions affect only immediate reward, storage decisions have long-term effects through reduced computation. We address this via a novel reduction to online submodular maximization.
IX Conclusion
In this paper, we have established a unified physical framework for intelligence, grounded in the fundamental axiom that information constitutes the enabling mapping from ontological states to carrier states [xu2025information, xu2024research]. By rigorously analyzing the thermodynamic costs associated with this mapping, we have demonstrated that the separation between ”static storage” (Shannon Entropy) and ”dynamic computation” (Logical Depth) is artificial; they are, in fact, conjugate variables governed by a fundamental constraint: the Energy-Time-Space Triality Bound.
Our proposed metric, Derivation Entropy ($H_{derive}$), serves as the quantitative bridge between these domains. It reveals that the efficiency of any intelligent system—biological or artificial—is strictly bounded by the trade-off between the entropic cost of maintaining memory and the dissipative heat of performing logical derivation [landauer1961irreversibility, bennett1982thermodynamics]. Specifically, we identified a Critical Derivation Threshold, a phase transition point beyond which it becomes thermodynamically cheaper to re-compute information from compressed laws rather than to retrieve it from vast databases.
This theoretical finding has profound implications for the design of next-generation AI architectures. As current models like Transformers approach the limits of memory bandwidth and energy consumption [vaswani2017attention, horowitz20141], our framework suggests that future systems may benefit from optimizing the Derivation-Retrieval Ratio rather than simply scaling parameters. Such systems could dynamically navigate the ”Logical Depth vs. Shannon Entropy” curve, acting as adaptive engines that minimize the total action of the ontological mapping process.
Ultimately, this work suggests that the ”simplicity” of a scientific theory or an AI model is not just an aesthetic preference but may reflect a thermodynamic imperative. By aligning the formalization of general states [qiu2025research] with physical constraints, we provide a roadmap for constructing intelligent systems that are not only logically powerful but physically sustainable. Future research will extend this framework to non-equilibrium open systems and quantum computing regimes, further exploring the ultimate physical limits of the enabling mapping.
Acknowledgment
During the writing and revision of this paper, we received many insightful comments from Associate Professor Wang Rui of the School of Computer Science at Shanghai Jiao Tong University and also gained much inspiration and assistance from regular academic discussions with doctoral students Yiming Wang, Chun Li, Hu Xu, Siyuan Qiu, Junxuan He, and Xiao Wang. We hereby express our sincere gratitude to them.