### v3

# Information Physics of Intelligence: Unifying Logical Depth and Entropy under Thermodynamic Constraints

Abstract
The rapid scaling of artificial intelligence models has revealed a fundamental tension between model capacity (storage) and inference efficiency (computation). While classical information theory focuses on transmission and storage limits, it lacks a unified physical framework to quantify the thermodynamic costs of generating information from compressed laws versus retrieving it from memory. In this paper, we propose a theoretical framework that treats information processing as an enabling mapping from ontological states to carrier states. We introduce a novel metric, Derivation Entropy, which quantifies the effective work required to compute a target state from a given logical depth. By analyzing the interplay between Shannon entropy (storage) and computational complexity (time/energy), we demonstrate the existence of a critical phase transition point. Below this threshold, memory retrieval is thermodynamically favorable; above it, generative computation becomes the optimal strategy. This ”Energy-Time-Space” conservation law provides a physical explanation for the efficiency of generative models and offers a rigorous mathematical bound for designing next-generation, energy-efficient AI architectures. Our findings suggest that the minimization of Derivation Entropy is a governing principle for the evolution of both biological and artificial intelligence.
I Introduction
The fundamental nature of information has long been a subject of debate across physics, computer science, and philosophy. In our previous works, we established a formalized meta-framework positing that information is fundamentally the enabling mapping from an ontological state to a carrier state [xu2025information, xu2024research, qiu2025research]. Under this view, intelligence is not merely the accumulation of data, but an active physical process of mapping the infinite complexity of the external environment (ontology) onto the finite constraints of an internal representation (carrier). This mapping inherently necessitates a primal trade-off: the system must decide whether to store the mapping results statically (memory) or to regenerate them dynamically via algorithms (computation).
Classically, these two aspects have been treated separately. Shannon’s information theory [shannon1948mathematical] rigorously defines the limits of static storage and transmission, quantifying the irreducible randomness of the carrier state. However, Shannon entropy is agnostic to the causal history or the computational effort required to generate that state. Conversely, Algorithmic Information Theory (AIT), particularly through the lens of Kolmogorov complexity [kolmogorov1965three] and Bennett’s concept of Logical Depth [bennett1988logical], focuses on the computational time required to reconstruct a structure from its compressed description. While Logical Depth captures the dynamic value of the ”mapping process,” it lacks a direct unification with the statistical mechanics governing the carrier’s physical existence.
This separation becomes problematic when we consider that information processing is strictly bound by physical laws. As Brillouin [brillouin1956science] and Landauer [landauer1961irreversibility] demonstrated, the manipulation of information—specifically the erasure of logical states—is inseparably linked to thermodynamic entropy and heat dissipation. Therefore, the ”enabling mapping” from ontology to carrier is not just a logical operation but a thermodynamic cycle. Current artificial intelligence systems, such as Large Language Models (LLMs) based on the Transformer architecture [vaswani2017attention, brown2020language], are rapidly approaching the physical limits of this cycle. The empirical scaling laws observed in these systems suggest a deep, underlying fundamental trade-off connecting parameter size (storage), training FLOPs (computation), and energy consumption, yet a unified theoretical explanation remains elusive.
In this paper, we propose the ”Information Physics of Intelligence,” a theoretical framework that unifies Logical Depth and Shannon Entropy under rigorous thermodynamic constraints. We define Derivation Entropy ($H_{derive}$), which quantifies the total physical cost of the ontological-to-carrier mapping. By integrating the thermodynamic cost of storage with the algorithmic complexity of retrieval, we derive the Energy-Time-Space Triality Bound. This bound reveals that minimizing storage space (lowering Shannon entropy) inevitably necessitates a compensatory increase in computational time (increasing Logical Depth) or energy flux, governed by Landauer’s limit.
Our contributions are threefold:
  1. 1.
    We formalize the ”ontological-to-carrier mapping” within a thermodynamic context, bridging the gap between abstract information measures [xu2025general] and physical complexity [bennett1982thermodynamics].
  2. 2.
    We mathematically derive the phase transition point where intelligent systems must switch from a retrieval-based strategy to a derivation-based strategy to maintain thermodynamic optimality.
  3. 3.
    We provide a first-principles explanation for the efficiency bottlenecks in modern AI models, offering a governing equation for designing next-generation, energy-efficient cognitive architectures.
II System Model and Preliminaries
We establish the formal framework for knowledge systems, introducing the concepts of enabling mappings, derivation depth, and their information-theoretic characterization.
II-A The Underlying Logical System $\mathcal{L}$
Throughout this work, we fix a formal logical system $\mathcal{L}$ that serves as the foundation for expressing semantic states and reasoning about information. Specifically, $\mathcal{L}$ is taken to be first-order logic with least fixed-point operators (FO(LFP)) [immerman1999descriptive], extended with multiple sorts, including at least $\mathsf{Obj}$ (for ontological entities), $\mathsf{Time}$ (for discrete time points), and $\mathsf{Carrier}$ (for physical carriers); function and relation symbols are typed accordingly (e.g., a state predicate may have signature $\mathsf{State}:\mathsf{Obj}\times\mathsf{Time}\to\mathsf{Bool}$). The language of $\mathcal{L}$ includes:
  • A countable set of sorted variables, constants, function symbols, and relation symbols;
  • Standard Boolean connectives and first-order quantifiers over each sort;
  • A least fixed-point operator $\mu$ enabling inductive definitions of relations (e.g., reachability, state transitions).
Crucially, we restrict attention to finite structures over a discrete, bounded time domain $T=\{t_{0},t_{1},\dots,t_{n}\}$ for some $n\in\mathbb{N}$. Under this restriction, $L$ enjoys the following essential properties:
  • Consistency: The proof system of $\mathcal{L}$ is sound and free of contradiction.
  • Effective model checking: For any finite structure $\mathcal{M}$ and formula $\varphi\in\mathcal{L}$, the satisfaction relation $\mathcal{M}\models\varphi$ is decidable.
  • Expressive adequacy: By the Immerman–Vardi theorem, FO(LFP) captures exactly the class of polynomial-time computable properties over ordered finite structures. This makes $\mathcal{L}$ sufficiently expressive to model the state evolution of a wide range of real-world information systems under finite-horizon autoregressive decoding.
II-B Enabling Mappings and Information
Definition 1 (State Expressibility).
Let $\mathcal{L}$ be the logical system defined above. The state set $S(X,T)$ of an object set $X$ over a time set $T$ is said to be expressible in $\mathcal{L}$ if it can be represented by assigning $X$, $T$, and their subsets and elements as interpretations of a set of well-formed formulas (WFFs) in $\mathcal{L}$.
In all subsequent discussions of this paper, we consider only states expressible in the logical system $\mathcal{L}$[qiu2025research, xu2025information].
Definition 2 (Enabling Mapping).
Let $O$ be an ontological space, $C$ a carrier space, and $T_{o},T_{r}\subseteq\mathbb{R}^{+}$ be time domains. A correspondence
$\mathcal{E}:S_{O}(O,T_{o})\rightrightarrows S_{C}(C,T_{r})$ (1)
is called an enabling mapping if the following conditions hold:
(E1) Realizability:
For every ontological state $s_{o}\in S_{O}(O,T_{o})$, the image $\mathcal{E}(s_{o})$ is a non-empty subset of $S_{C}(C,T_{r})$. That is, there exists at least one valid carrier state $s_{c}\in\mathcal{E}(s_{o})$ that can physically represent $s_{o}$.
(E2) Causality:
For every $s_{o}$ and any realization $s_{c}\in\mathcal{E}(s_{o})$, we have $s_{o}\prec s_{c}$, where $\prec$ denotes a causal dependency relation, implying: (i) The occurrence time of $s_{o}$ precedes or is concurrent with the realization of $s_{c}$; (ii) The ontological state $s_{o}$ is a necessary condition for the intentional generation of $s_{c}$.
(E3) Physicality:
The realization of any mapping instance $s_{o}\mapsto s_{c}$ (where $s_{c}\in\mathcal{E}(s_{o})$) consumes non-zero physical resources (e.g., energy, time).
Remark 1.
[…]
Definition 3 (Information).
Information is an enabling mapping [xu2024research]:
$\mathcal{I}:S_{O}(O,T_{o})\rightrightarrows S_{C}(C,T_{r})$ (2)
where
  • $O$: ontology (objective or subjective reality)
  • $T_{o}\subseteq\mathbb{R}^{+}$: occurrence time
  • $S_{O}=S_{O}(O,T_{o})$: state set of ontology
  • $C$: carrier (objective reality)
  • $T_{r}\subseteq\mathbb{R}^{+}$: reflection time
  • $S_{C}=S_{C}(C,T_{r})$: state set of carrier
Equivalently, information can be represented as a sextuple $\mathcal{I}=\langle O,T_{o},S_{O},C,T_{r},S_{C}\rangle$. It is also assumed to be effectively computable on semantic inputs
Definition 4 (Ideal Information).
Let $S_{O}$ and $S_{C}$ be defined as above. Ideal Information $\mathcal{I}$ is an enabling mapping satisfying two additional strict conditions:
Completeness (Surjectivity on Domain):
Every semantic state $s_{o}\in S_{O}$ has a physical representation. Formally, $\forall s_{o}\in S_{O},\mathcal{I}(s_{o})\neq\emptyset$. (Inherited from E1).
Reliability (Unambiguous Decoding):
While a semantic state may map to multiple physical carriers (synonyms), any specific physical carrier maps back to at most one semantic state. Formally, if $s_{c}\in\mathcal{I}(s_{o_{1}})$ and $s_{c}\in\mathcal{I}(s_{o_{2}})$, then $s_{o_{1}}=s_{o_{2}}$.
Remark 2.
The set $\mathcal{I}(s_{o})\subseteq S_{C}$ constitutes the synonymous carrier set of the semantic state $s_{o}\in S_{O}$. Therefore, the semantic characteristics of an Ideal Information instance $\mathcal{I}=\langle O,T_{o},S_{O},C,T_{r},S_{C}\rangle$ can be fully characterized by directly studying the semantic state space $S_{O}$, with each $s_{o}\in S_{O}$ being equipped with its associated synonymous carrier set $\mathcal{I}(s_{o})$.
II-C Information Cost of Energy and Time
According to Definitions 23, state transitions of an information carrier are constrained by physical conditions. Building on fundamental results such as the energy requirements of high-frequency harmonic oscillators in classical physics [Nyquist1928Thermal] and the Margolus–Levitin bound in quantum mechanics [Margolus1998maximum], we adopt the following abstract physical principle.
Assumption 1 (Physical Cost Principle).
For a physical system evolving from a state $s_{1}$ to a distinguishable state $s_{2}$, the required energy $E$ and time $\tau$ satisfy
$E\cdot\tau\;\geq\;\kappa,$ (3)
where $\kappa$ is a universal constant with the dimension of action.
Intuitively, Assumption 1 states that every physically distinguishable transition of the carrier system incurs a non-zero resource cost that cannot be made arbitrarily small: the product of the available energy scale and the time scale must exceed a fixed action threshold $\kappa$. In particular, a physical process that implements an information mapping cannot realize arbitrarily many mutually distinguishable semantic alternatives without paying at least a proportional cost in the joint resource $E\cdot\tau$.
In the context of Ideal Information (Definition 4), the semantic state space $S_{O}(O,T_{o})$ has finite cardinality, and the corresponding amount of information is measured in bits as
$I\;=\;\log_{2}|S_{O}(O,T_{o})|.$ (4)
Because Ideal Information is complete and reliably decodable, any physical implementation must, in effect, be able to distinguish among all $|S_{O}(O,T_{o})|=2^{I}$ possible semantic inputs via mutually distinguishable carrier states. This leads to the following resource lower bound.
Theorem 1 (Resource Bound for Ideal Enabling Mapping).
Let $\mathcal{I}:S_{O}(O,T_{o})\rightrightarrows S_{C}(C,T_{r})$ be an Ideal Information instance, and let $I=\log_{2}|S_{O}(O,T_{o})|$ denote the amount of information transmitted (in bits) as in (4). Then any physical implementation of $\mathcal{I}$ satisfies
$E\cdot\tau\;\geq\;\kappa\cdot I,$ (5)
where $E$ is the energy resource, $\tau$ is the implementation time, and $\kappa$ is the universal constant from Assumption 1.
Since $\mathcal{I}$ is Ideal Information in the sense of Definition 4, the mapping $\mathcal{I}:S_{O}(O,T_{o})\rightrightarrows S_{C}(C,T_{r})$ is complete and reliable: every semantic state $s_{o}\in S_{O}(O,T_{o})$ has at least one physical realization $s_{c}\in S_{C}(C,T_{r})$, and distinct semantic states are not confounded at the carrier level. Hence any physical implementation of $\mathcal{I}$ must realize at least $|S_{O}(O,T_{o})|$ pairwise distinguishable carrier states, one (or a synonymous set) for each semantic state.
Equivalently, the implementation process must be capable of distinguishing among $|S_{O}(O,T_{o})|=2^{I}$ alternative semantic inputs. This corresponds to an information content of $I=\log_{2}|S_{O}(O,T_{o})|$ bits.
Consider now the physical evolution of the carrier system during the implementation of $\mathcal{I}$. Each time the system evolves from one distinguishable physical state to another in such a way that it increases or commits a new semantic distinction (for instance, by encoding one more bit of choice among alternative semantic states), Assumption 1 applies: the associated energy and time resources $E_{i}$ and $\tau_{i}$ must satisfy
$E_{i}\cdot\tau_{i}\;\geq\;\kappa.$
Let $N$ denote the number of such information-relevant, distinguishable state transitions required by the implementation. Since each such transition can encode at most a constant number of bits of new distinguishable semantic information, we have $N\geq I$ up to a fixed multiplicative constant. Absorbing this constant into $\kappa$ if necessary, we may assume without loss of generality that $N\geq I$.
Summing the resource inequality over all $N$ transitions yields
$E\cdot\tau\;\geq\;\sum_{i=1}^{N}E_{i}\cdot\tau_{i}\;\geq\;N\kappa\;\geq\;\kappa I,$
where $E$ and $\tau$ denote, respectively, the total energy resource and the total implementation time available to the physical process. This establishes the claimed lower bound (5). ∎
In the following, we interpret the constant $\kappa$ in Assumption 1 in the usual thermodynamic setting of information processing. Specifically, for systems operating in a thermal equilibrium environment at temperature $T$ and undergoing quasistatic transformations, the minimal resource cost per bit can be identified with the Landauer bound $k_{B}T\ln 2$ measured in terms of dissipated heat. Under this interpretation, Theorem 1 specializes to a Landauer-type inequality for Ideal Information.
Corollary 1 (Landauer Principle for Ideal Information).
In a thermal equilibrium environment at temperature $T$, an Ideal Information instance implemented via a quasistatic process satisfies the following lower bound on energy dissipation:
$E_{\mathrm{diss}}\;\geq\;k_{B}T\ln 2\cdot I,$ (6)
where $I=\log_{2}|S_{O}(O,T_{O})|$ is the amount of information transmitted losslessly (in bits).
In a thermal environment at temperature $T$, Landauer’s principle[landauer1961irreversibility] states that any logically irreversible operation that erases or thermodynamically commits one bit of information leads, in the quasistatic limit, to a minimal heat dissipation of at least $k_{B}T\ln 2$. In the setting of Theorem 1, implementing an ideal information instance that reliably encodes $I=\log_{2}|S_{O}(O,T_{o})|$ bits requires at least $I$ effective information-bearing operations, each of which is thermodynamically equivalent to committing one bit of semantic information to a physical carrier state.
Identifying the per-bit resource constant in Assumption 1 with $\kappa_{T}=k_{B}T\ln 2$ in this thermal, quasistatic regime, Theorem 1 implies that the total dissipated energy satisfies
$E_{\mathrm{diss}}\;\geq\;\kappa_{T}\cdot I\;=\;k_{B}T\ln 2\cdot I,$
which is exactly the bound in (6). This yields the claimed Landauer-type lower bound for any lossless implementation of Ideal Information. ∎
II-D Atomic Information
In an Ideal Information instance $\mathcal{I}=\langle O,T_{o},S_{O},C,T_{r},S_{C}\rangle$ (Definition 4), the ontological state space $S_{O}$ is equipped with a predecessor function
$P_{O}:S_{O}\to 2^{S_{O}},$ (7)
which captures the semantic causal structure among states. Intuitively, $P_{O}(s)$ is the set of semantic states that can directly lead to $s$ within the ontological dynamics of $O$. We impose the following axioms on $P_{O}(\cdot)$.
Axiom 1 (Strict Temporal Precedence).
For any $s_{o}(o,t),s^{\prime}_{o^{\prime}}(o^{\prime},t^{\prime})\in S_{O}$, if $s^{\prime}_{o^{\prime}}\in P_{O}(s_{o})$, then
$t^{\prime}<t.$
Axiom 2 (Finite and Computable Predecessors).
For any $s\in S_{O}$, the predecessor set $P_{O}(s)$ is finite, and there exists an effective procedure that, given $s$, computes $P_{O}(s)$.
Axiom 3 (Lower-Bounded and Discrete Time).
The time domain $T_{o}$ is bounded from below and discrete, i.e., there exists $t_{\min}\in T_{o}$ and $\Delta t>0$ such that for any strictly decreasing sequence $\{t_{n}\}$ in $T_{o}$ we have $t_{n}\geq t_{\min}$ and $t_{n}-t_{n+1}\geq\Delta t$.
Axiom 1 states that predecessors must occur strictly earlier in ontological time; Axiom 2 excludes infinitely branching causal histories and ensures that local causal structure is effectively inspectable; and Axiom 3 guarantees that there is no infinite regress of ever earlier states in finite time. Together, these axioms imply that the backward causal unfolding of any semantic state is a finite, effectively traversable structure. This leads to the following result.
Theorem 2 (Existence and Computability of Semantic Atomic States).
For any target semantic state $s_{\text{target}}\in S_{O}$, there exists a finite, non-empty set of semantic atomic states $A(s_{\text{target}})\subseteq S_{O}$ such that:
  1. (i)
    Each $s_{\text{atom}}\in A(s_{\text{target}})$ is predecessor-free:
    $P_{O}(s_{\text{atom}})=\varnothing.$
  2. (ii)
    Every $s_{\text{atom}}\in A(s_{\text{target}})$ can reach $s_{\text{target}}$ along the causal structure induced by $P_{O}(\cdot)$, i.e., there exists a finite sequence
    $s_{\text{atom}}=s_{0},s_{1},\dots,s_{n}=s_{\text{target}}$
    such that $s_{i}\in P_{O}(s_{i+1})$ for all $i=0,\dots,n-1$.
  3. (iii)
    The mapping $s_{\text{target}}\mapsto A(s_{\text{target}})$ is computable.
Fix any $s_{\text{target}}\in S_{O}$. Consider the backward expansion of $s_{\text{target}}$ along the predecessor function $P_{O}(\cdot)$: start from $s_{\text{target}}$, enumerate $P_{O}(s_{\text{target}})$, then for each $s^{(1)}\in P_{O}(s_{\text{target}})$ enumerate $P_{O}(s^{(1)})$, and so on.
By Axiom 1, every time we move from a state to one of its predecessors, the time-stamp strictly decreases: if $s^{\prime}_{o^{\prime}}(o^{\prime},t^{\prime})\in P_{O}(s_{o}(o,t))$, then $t^{\prime}<t$. Since $T_{o}$ is lower-bounded and discrete (Axiom 3), there is no infinite strictly decreasing temporal sequence. Therefore, any backward traversal starting from $s_{\text{target}}$ must terminate in finitely many steps. Equivalently, the backward expansion tree is of finite depth.
By Axiom 2, each node has a finite predecessor set, so the branching factor of the backward expansion tree is finite. Thus the entire backward tree rooted at $s_{\text{target}}$ consists of finitely many nodes. The leaves of this finite tree are exactly those states with no predecessors, i.e., all $s\in S_{O}$ such that $P_{O}(s)=\varnothing$ and that can reach $s_{\text{target}}$ along the tree.
Define $A(s_{\text{target}})$ to be the set of all such leaves. Since the tree is finite and non-empty (it contains at least $s_{\text{target}}$), $A(s_{\text{target}})$ is finite and non-empty, and every element has no predecessors, establishing (i) and (ii).
For (iii), consider the following recursive procedure
TRACEATOMS(s)(s):
  • If $P_{O}(s)=\varnothing$, output $\{s\}$.
[…]
Theorem 2 shows that, for any semantic state of interest, there exists a finite and computable collection of causally irreducible semantic states that underpin it. We now upgrade this semantic view to the level of physically realizable information.
Definition 5 (Atomic Information).
Let $\mathcal{I}=\langle O,T_{o},S_{O},C,T_{r},S_{C}\rangle$ be an Ideal Information instance and let $\alpha\in S_{O}$ be any semantic state. The Atomic Information associated with $\alpha$, denoted by $\mathcal{I}_{\alpha}$, is defined as the restriction of $\mathcal{I}$ to the singleton $\{\alpha\}$:
$\mathcal{I}_{\alpha}\;=\;\langle O,\,T_{o},\,\{\alpha\},\,C,\,T_{r},\,\mathcal{I}(\alpha)\rangle,$ (8)
where
$\mathcal{I}(\alpha)=\left\{\,s_{c}\in S_{C}\mid\begin{array}[]{@{}l@{}}\text{the realization of }s_{c}\text{ yields}\\ \text{the semantic state }\alpha\end{array}\right\}$
is the set of carrier states synonymous with $\alpha$ under the enabling mapping $\mathcal{I}:S_{O}\rightrightarrows S_{C}$.
Because $\mathcal{I}$ is Ideal Information (Definition 4), it satisfies completeness and unambiguous decoding. In particular, for every semantic state $\alpha\in S_{O}$ we have $\mathcal{I}(\alpha)\neq\varnothing$, so there always exists at least one carrier state $s_{c}\in S_{C}$ that physically realizes $\alpha$. Thus each $\mathcal{I}_{\alpha}$ defined in (8) is a physically realizable information entity that tightly couples an abstract semantic state with the set of its synonymous physical representations.
Combining Theorem 2 with the above definition yields the following corollary.
Corollary 2 (Existence and Computability of Atomic Information).
Let $\mathcal{I}=\langle O,T_{o},S_{O},C,T_{r},S_{C}\rangle$ be an Ideal Information instance, and let $s_{\text{target}}\in S_{O}$ be any semantic state. Then:
  1. 1.
    There exists a finite, non-empty set of semantic atomic states $A(s_{\text{target}})\subseteq S_{O}$ such that each $\alpha\in A(s_{\text{target}})$ is predecessor-free and causally contributes to $s_{\text{target}}$ as in Theorem 2.
  2. 2.
    For every $\alpha\in A(s_{\text{target}})$, the corresponding Atomic Information
    $\mathcal{I}_{\alpha}\;=\;\langle O,T_{o},\{\alpha\},C,T_{r},\mathcal{I}(\alpha)\rangle$
    is well-defined and physically realizable, and the finite set $\{\mathcal{I}_{\alpha}\mid\alpha\in A(s_{\text{target}})\}$ is computable from $s_{\text{target}}$.
Item (1) is exactly Theorem 2.
For item (2), since $\mathcal{I}$ is Ideal Information, its completeness property ensures $\mathcal{I}(\alpha)\neq\varnothing$ for every $\alpha\in S_{O}$; in particular, for each $\alpha\in A(s_{\text{target}})$, $\mathcal{I}(\alpha)$ is a non-empty set of synonymous carrier states, so $\mathcal{I}_{\alpha}$ in (8) is well-defined and physically realizable.
Moreover, Theorem 2 provides a total, effective procedure TRACEATOMS that, given $s_{\text{target}}$, computes $A(s_{\text{target}})$. Under the system model adopted in this paper, the enabling mapping $\mathcal{I}(\cdot)$ is also assumed to be computable for each input semantic state. Therefore, for every $\alpha\in A(s_{\text{target}})$ we can effectively enumerate $\mathcal{I}(\alpha)$ and construct the corresponding Atomic Information $\mathcal{I}_{\alpha}$. Since $A(s_{\text{target}})$ is finite, the set $\{\mathcal{I}_{\alpha}\mid\alpha\in A(s_{\text{target}})\}$ is finite and computable, as claimed. ∎
II-E Semantic Inference Structure
The predecessor function $P_{O}(\cdot)$ introduced above captures the semantic causal structure among ontological states. In order to connect this causal view with the proof-theoretic notion of derivation in the underlying logical system $\mathcal{L}$, we make the following alignment assumption.
Assumption 2 (Alignment of Causality and Inference).
For every semantic state $s\in S_{O}$, the predecessor set $P_{O}(s)$ coincides with the set of its immediate inference premises in $\mathcal{L}$. More precisely:
  • Soundness of predecessors: For every $s^{\prime}\in P_{O}(s)$, there exists a single inference step in $\mathcal{L}$ that derives $s$ from $s^{\prime}$ (possibly together with elements of $P_{O}(s)\setminus\{s^{\prime}\}$ as side premises).
  • Completeness of predecessors: Conversely, whenever there is a single-step inference in $\mathcal{L}$ whose conclusion is $s$, all semantic premises that this inference depends on are contained in $P_{O}(s)$.
In other words, the directed graph induced by $P_{O}(\cdot)$ is (up to inessential bureaucratic details of the proof system) identical to the derivation graph generated by single-step inference rules of $\mathcal{L}$ over $S_{O}$.
On this basis, we define the global set of semantic atomic states as
$\text{Atom}(S_{O})\coloneqq\{\,s\in S_{O}\mid P_{O}(s)=\emptyset\,\}.$ (9)
Intuitively, $\text{Atom}(S_{O})$ collects all ontological states that have no semantic predecessors and therefore serve as the “primitive” or “axiomatic” basis for subsequent derivations in $\mathcal{L}$. By Axioms 1–3, $\text{Atom}(S_{O})$ is well-defined; it may be finite or infinite, but any particular target state $s_{\text{target}}$ depends only on a finite subset of $\text{Atom}(S_{O})$ identified in Theorem 2.
II-F Derivation Depth
With the atomic basis and the semantic inference structure in place, we can now define the complexity of any derived state in terms of its distance from the atomic layer along the causal/inference graph.
Definition 6 (Derivation Depth).
Under Assumption 2, the predecessor function $P_{O}(\cdot)$ induces the immediate semantic inference structure of $\mathcal{L}$ over $S_{O}$. For any state $s\in S_{O}$, the Derivation Depth of $s$, denoted $\text{Dd}(s)$, is defined inductively as follows:
$\text{Dd}(s)\coloneqq\begin{cases}0,&\text{if }P_{O}(s)=\emptyset,\\[2.15277pt] 1+\displaystyle\min_{s^{\prime}\in P_{O}(s)}\text{Dd}(s^{\prime}),&\text{otherwise.}\end{cases}$ (10)
Equivalently, $\text{Dd}(s)$ is the length of the shortest finite derivation chain
$s_{0},s_{1},\dots,s_{n}=s$ (11)
such that
  1. 1.
    $s_{0}\in\text{Atom}(S_{O})$ (i.e., $P_{O}(s_{0})=\emptyset$);
  2. 2.
    $s_{i}\in P_{O}(s_{i+1})$ for all $i=0,\dots,n-1$.
In this case we write
[…]
In particular, for any atomic state $\alpha\in\text{Atom}(S_{O})$, we have $\text{Dd}(\alpha)=0$.
Theorem 3 (Computability of Derivation Depth).
For any state $s\in S_{O}$, the derivation depth $\text{Dd}(s)$ defined in (10) is a unique, finite, and computable integer.
We view the directed graph over $S_{O}$ induced by the predecessor function $P_{O}(\cdot)$. By Axiom 1 (Strict Temporal Precedence), every edge $s^{\prime}\in P_{O}(s)$ leads from $s$ to a strictly earlier time-stamp $t^{\prime}<t$. By Axiom 3 (Lower-Bounded and Discrete Time), there is no infinite strictly decreasing sequence of time-stamps. Hence, every backward path
$s=s_{n},s_{n-1},\dots,s_{0}$
following predecessors $s_{i-1}\in P_{O}(s_{i})$ has finite length. Equivalently, the backward expansion tree rooted at any $s\in S_{O}$ has finite depth.
By Axiom 2 (Finite and Computable Predecessors), each state $s$ has a finite, effectively computable predecessor set $P_{O}(s)$. Thus, the backward expansion from $s$ yields a finite, effectively traversable directed acyclic graph (DAG) whose edges are given by the predecessor relation.
Now consider the recursive definition (10):
$\text{Dd}(s)=\begin{cases}0,&\text{if }P_{O}(s)=\emptyset,\\[2.15277pt] 1+\displaystyle\min_{s^{\prime}\in P_{O}(s)}\text{Dd}(s^{\prime}),&\text{otherwise.}\end{cases}$
Well-foundedness. Every recursive call from $s$ to $s^{\prime}\in P_{O}(s)$ strictly decreases the associated time-stamp (by Axiom 1). Since time is bounded below and discrete (Axiom 3), there is no infinite descending chain of such calls. Therefore, the recursion is well-founded.
Termination. The recursion depth is bounded by the maximal length of a backward predecessor chain starting from $s$, which is finite as argued above. Hence, the evaluation of $\text{Dd}(s)$ terminates after finitely many steps.
Uniqueness and computability. For any non-atomic state $s$ with $P_{O}(s)\neq\emptyset$, the set $\{\text{Dd}(s^{\prime})\mid s^{\prime}\in P_{O}(s)\}$ consists of finitely many well-defined integers (by the well-founded recursion). The minimum over this finite set is unique. Moreover, because $P_{O}(s)$ is effectively computable for every $s$ (Axiom 2), we can compute $\text{Dd}(s)$ by a bottom-up dynamic-programming traversal of the finite backward DAG: first assign depth 0 to all atomic states ($P_{O}(s)=\emptyset$), then iteratively compute $\text{Dd}(s)$ for states whose predecessors’ depths are already known. Thus, $\text{Dd}(s)$ is a unique, finite, and computable integer for every $s\in S_{O}$. ∎
Proposition 1 (Derivation depth as Bennett-style logical depth).
Let $\mathcal{L}$ be the underlying logical system and let $P$ be a fixed sound and complete proof system for $\mathcal{L}$. Let $\Gamma$ be a finite base theory, and let $\varphi$ be any formula such that $\Gamma\vdash_{P}\varphi$. We write $Dd(\varphi\mid\Gamma)$ for the minimal derivation depth of $\varphi$ with respect to the semantic/inference structure induced by $\Gamma$ in the sense of Definition 6.
Fix a universal Turing machine $U$, and fix a computable encoding scheme $\langle\cdot\rangle$ that encodes formulas, proofs, and finite sets of formulas as finite strings. Assume that the following two conditions hold.
  1. (EPS)
    Efficient proof simulation. There exists a constant $c_{1}>0$ such that, for any proof $\pi$ in $P$ of length [^1] Here the length of $\pi$ is measured as the number of inference steps in $P$. $n$ whose conclusion is $\varphi$, the universal machine $U$, given as input $\langle\mathrm{Atom}(\Gamma)\rangle$ and $\langle\pi\rangle$, can verify $\pi$ and output $\langle\varphi\rangle$ within time at most $c_{1}\cdot n$.
  2. (NPC)
    No super-linear proof compression into time. There exists a constant $c_{2}>0$ such that, for any program $p$ for $U$ with oracle access to $\mathrm{Atom}(\Gamma)$ that outputs $\langle\varphi\rangle$ in time $T$, there is a proof $\pi$ in $P$ of $\varphi$ from $\Gamma$ of length at most $c_{2}\cdot T$.
Then there exist constants $c,C>0$ (depending only on $U$, $P$ and the encoding scheme, but not on $\Gamma$ or $\varphi$) such that for all $\varphi$ with $\Gamma\vdash_{P}\varphi$ we have
$\begin{split}c\cdot\mathrm{Dd}(\varphi\mid\Gamma)\;\leq\;&\mathrm{depth}_{C\cdot\mathrm{Dd}(\varphi\mid\Gamma)}\bigl(\langle\varphi\rangle\mid\langle\mathrm{Atom}(\Gamma)\rangle\bigr)\\ \;\leq\;&C\cdot\mathrm{Dd}(\varphi\mid\Gamma),\end{split}$ (12)
where $\mathrm{depth}_{t}(x\mid y)$ denotes Bennett’s logical depth of string $x$ at significance level $t$, conditional on $y$, with respect to the universal machine $U$.
In particular, under (EPS) and (NPC) we have
$\mathrm{depth}_{O(Dd(\varphi\mid\Gamma))}\bigl(\langle\varphi\rangle\mid\langle\mathrm{Atom}(\Gamma)\rangle\bigr)\;=\;\Theta\bigl(Dd(\varphi\mid\Gamma)\bigr),$ (13)
where the implicit constants in the $O(\cdot)$ and $\Theta(\cdot)$ notation depend only on $U$, $P$, and the chosen encodings.
III Derivation Depth as an Information Metric
We now establish that derivation depth is an information-theoretic quantity dual to Shannon entropy, bridging the gap between logical reasoning and information processing.
III-A Derivation Entropy and Information Content
Definition 7 (Derivation Entropy – General Form).
The derivation entropy from $\psi$ to $\phi$ is defined as:
$H_{\mathrm{derive}}(\phi|\psi)=\mathrm{Dd}(\phi|\psi)\cdot\lambda,$ (14)
where $\mathrm{Dd}(\phi|\psi)$ is the derivation depth (Definition 6) and $\lambda>0$ is the information cost per inference step.
Remark 3 (Physical Interpretation of $\lambda$).
The parameter $\lambda$ quantifies the information-theoretic cost per inference step:
  1. 1.
    Classical thermodynamics: Each logically irreversible step erases $\geq 1$ bit. By Landauer’s principle, $\lambda=\ln 2$ nats.
  2. 2.
    Quantum mechanics: By the Margolus–Levitin bound, $\lambda=\pi/2$ in units of $\hbar$.
  3. 3.
    Proof system structure: Selecting from $|\mathcal{R}|$ rules per step suggests $\lambda\gtrsim\ln|\mathcal{R}|$ nats, as a structural lower bound on the information cost of inference.
TABLE I: Analogy Between Thermodynamic and Information-Theoretic Costs
Domain Discrete Unit Continuous Unit Conversion Factor
Thermodynamics 1 bit erasure Energy (J) $k_{B}T\ln 2$
Information Physics 1 inference step Information (nats) $\ln 2$
Assumption 3 (Classical Thermodynamic Regime).
The knowledge system operates in the classical thermodynamic regime: (1) Inference steps are logically irreversible at thermal equilibrium; (2) Each step erases $O(1)$ bits; (3) Landauer erasure dominates other costs. Under this assumption, $\lambda=\ln 2$ nats.
Corollary 3 (Derivation Entropy in Classical Regime).
Under Assumption 3: $H_{\mathrm{derive}}(\phi|\psi)=\mathrm{Dd}(\phi|\psi)\cdot\ln 2$.
Lemma 1 (Derivation Encoding).
Let $\mathcal{P}$ be a proof system with rule set $R$ and maximum rule arity $k$. Any derivation $\pi$ of length $n$ from $m$ premises satisfies:
$|\mathrm{Enc}(\pi)|\leq n\cdot c_{L}\cdot\log(m+n)+O(\log(m+n)),$ (15)
where $c_{L}=\max\{\lceil\log_{2}|R|\rceil,k\}$.
A derivation $\pi=((r_{1},\vec{p}_{1}),\ldots,(r_{n},\vec{p}_{n}))$ consists of $n$ steps. Each step $(r_{i},\vec{p}_{i})$ encodes:
  • Rule: $r_{i}\in R$ in $\lceil\log_{2}|R|\rceil$ bits.
  • Premises: $\vec{p}_{i}=(p_{i,1},\ldots,p_{i,k_{i}})$ with $k_{i}\leq k$ pointers, each in $\{1,\ldots,m+i-1\}$, requiring $\leq k\cdot\lceil\log_{2}(m+n)\rceil$ bits total.
Per-step encoding: $\lceil\log_{2}|R|\rceil+k\cdot\lceil\log_{2}(m+n)\rceil\leq c_{L}\cdot(1+\log_{2}(m+n))$.
Over $n$ steps: $n\cdot c_{L}\cdot(1+\log_{2}(m+n))$.
Adding $O(\log(m+n))$ header bits for $m,n$, and noting that $n\cdot c_{L}\leq n\cdot c_{L}\cdot\log(m+n)$ when $\log(m+n)\geq 1$:
$|\mathrm{Enc}(\pi)|\leq n\cdot c_{L}\cdot\log(m+n)+O(\log(m+n)).\qed$
Definition 8 (Information-Rich Regime).
Let $S_{O}$ be a knowledge base and let $q$ be a query (logical formula) over $S_{O}$. We define the conditional algorithmic entropy of $q$ given $S_{O}$ as
$H(q\mid S_{O})\;:=\;K\bigl(q\mid\langle\text{Atom}(S_{O})\rangle\bigr),$ (16)
where $K(\cdot\mid\cdot)$ denotes conditional Kolmogorov complexity with respect to a fixed universal Turing machine and a computable encoding scheme $\langle\cdot\rangle$.
We say that $q$ lies in the information-rich regime (with respect to $S_{O}$) if
$H(q\mid S_{O})\;\geq\;\omega\bigl(\log|\text{Atom}(S_{O})|\bigr),$ (17)
that is, equivalently,
$\frac{H(q\mid S_{O})}{\log|\text{Atom}(S_{O})|}\;\longrightarrow\;\infty\qquad\text{as }|\text{Atom}(S_{O})|\to\infty.$ (18)
For query classes satisfying the polynomial-depth condition
$Dd(q\mid S_{O})\;\leq\;\mathrm{poly}\bigl(|\text{Atom}(S_{O})|\bigr),$ (19)
(e.g., Datalog, SPARQL, and template-based natural-language queries), conditions (17)–(18) are asymptotically equivalent to
$H(q\mid S_{O})\;\gg\;\log\max\bigl\{|\text{Atom}(S_{O})|,\,Dd(q\mid S_{O})\bigr\},$ (20)
as $|\text{Atom}(S_{O})|\to\infty$.
Lemma 2 (Derivation Incompressibility in the Information-Rich Regime).
Let $S_{O}$ be a knowledge base and $q$ a logical formula such that:
  1. (i)
    $n=Dd(q\mid S_{O})\geq 1$,
  2. (ii)
    $m=|\mathrm{Atom}(S_{O})|$,
  3. (iii)
    $q$ lies in the information-rich regime (Definition 8),
  4. (iv)
    $n=\mathrm{poly}(m)$ (e.g., for bounded-depth reasoning systems).
Then there exists a constant $c_{2}>0$, depending only on the proof system and encoding scheme, such that the conditional Kolmogorov complexity of $q$ given $S_{O}$ satisfies
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)\geq c_{2}\cdot n\log(m+n)-O(\log m).$
By Lemma 1, any derivation of $q$ from $\mathrm{Atom}(S_{O})$ of length $n$ can be encoded in at most
$L_{\text{enc}}=c_{1}\cdot n\log(m+n)+O(\log m)$
bits for some constant $c_{1}>0$. This establishes an upper bound on $K(q\mid\langle\mathrm{Atom}(S_{O})\rangle)$.
Now suppose, for contradiction, that no such lower bound holds. Then for every $\varepsilon>0$, there exist arbitrarily large $m$ and corresponding $q$ such that
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)<\varepsilon\cdot n\log(m+n).$
Since $n=\mathrm{poly}(m)$, we have $\log(m+n)=\Theta(\log m)$, so the above implies
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)=o\bigl(n\log m\bigr).$
Consequently, the conditional algorithmic entropy satisfies
$H(q\mid S_{O})=K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)\cdot\ln 2=o\bigl(n\log m\bigr).$
But because $n\geq 1$ and $n=\mathrm{poly}(m)$, we have $n\log m\geq\log m$, and in fact $n\log m\gg\log m$ as $m\to\infty$. Hence
$H(q\mid S_{O})=o\bigl(n\log m\bigr)\not\gg\log m,$
which contradicts condition (iii)—that $q$ lies in the information-rich regime. Therefore, there must exist a constant $c_{2}>0$ such that
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)\geq c_{2}\cdot n\log(m+n)-O(\log m),$
as claimed. ∎
Throughout this subsection, all asymptotic notation ($O(\cdot)$, $\Theta(\cdot)$, $\omega(\cdot)$, $\gg$, etc.) is taken with respect to
$m\;:=\;|\text{Atom}(S_{O})|\;\longrightarrow\;\infty,$
under the standing assumption that
$n\;:=\;Dd(q\mid S_{O})\;=\;\mathrm{poly}(m).$
Theorem 4 (Derivation Depth as an Information Metric).
Let $S_{O}$ be a knowledge base and let $q$ be a query satisfying conditions (i)(iv) of Lemma 2 (in particular, $q$ lies in the information-rich regime of Definition 8 and $n:=Dd(q\mid S_{O})=\mathrm{poly}(m)$ with $m:=|\text{Atom}(S_{O})|$). Define
$H(q\mid S_{O})\;:=\;K\bigl(q\mid\langle\text{Atom}(S_{O})\rangle\bigr)\cdot\ln 2,$ (21)
and, under the classical thermodynamic regime (Assumption 3), define the derivation entropy as
$H_{\mathrm{derive}}(q\mid S_{O})\;:=\;Dd(q\mid S_{O})\cdot\ln 2.$ (22)
Then
$H_{\mathrm{derive}}(q\!\mid\!S_{O})\!=\!\Theta\!\left(\frac{H(q\mid S_{O})}{\!\log\!\big(|\text{Atom}(S_{O})|\!+\!\mathrm{Dd}(q\mid S_{O})\big)}\right)\!,$ (23)
where the implicit constants in the $\Theta(\cdot)$ notation depend only on the choice of universal Turing machine, proof system, and encoding scheme, but not on $S_{O}$ or $q$.
Let $m=|\text{Atom}(S_{O})|$ and $n=Dd(q\mid S_{O})$. By Lemma 1 and Lemma 2, there exist constants $c_{1},c_{2}>0$ such that
$c_{2}\,n\log(m+n)-O(\log m)\;\leq\;K\bigl(q\mid\langle\text{Atom}(S_{O})\rangle\bigr)\\ \;\leq\;c_{1}\,n\log(m+n)+O(\log m).$
Multiplying both sides by $\ln 2$ yields
$c_{2}^{\prime}\,n\log(m+n)-O(\log m)\;\leq\;H(q\mid S_{O})\\ \;\leq\;c_{1}^{\prime}\,n\log(m+n)+O(\log m),$
for suitable constants $c_{1}^{\prime},c_{2}^{\prime}>0$ with $c_{i}^{\prime}=c_{i}\ln 2$. Because $n=\mathrm{poly}(m)$, we have $\log(m+n)=\Theta(\log m)$, and the $O(\log m)$ term is asymptotically dominated by $n\log(m+n)$. Hence
$H_{\mathrm{derive}}(q\!\mid\!S_{O})\!=\!\Theta\!\left(\!\frac{H(q\mid S_{O})}{\log\!\big(|\text{Atom}(S_{O})|\!+\!\mathrm{Dd}(q\mid S_{O})\big)}\!\right)\!,$
Rearranging gives
$n\;=\;\Theta\!\left(\frac{H(q\mid S_{O})}{\log(m+n)}\right).$
Finally, substituting $H_{\mathrm{derive}}(q\mid S_{O})=n\cdot\ln 2$ and recalling that $m=|\text{Atom}(S_{O})|$ and $n=Dd(q\mid S_{O})$, we obtain
$H_{\mathrm{derive}}(q\!\mid\!S_{O})\!=\!\Theta\!\left(\frac{H(q\mid S_{O})}{\!\log\!\big(|\text{Atom}(S_{O})|\!+\!\mathrm{Dd}(q\mid S_{O})\big)}\right)\!,$
which is exactly (23). This completes the proof. ∎
III-B Frequency-weighted storage–computation tradeoff
In this subsection we combine the semantic information measures developed in Section II with the derivation-depth analysis in Section III.A to obtain a quantitative tradeoff between storage and computation under a given query distribution. Intuitively, if a query $q$ will be accessed frequently, it may be advantageous to “precompute and store” part of the information needed to answer $q$; if $q$ is rare, it is better to compute the answer on demand from the base knowledge $S_{O}$. Our main result in this subsection, Theorem 5, formalizes this intuition in terms of semantic information content, derivation entropy, and frequency-weighted amortized cost.
Definition 9 (Access Frequency).
For a query distribution $P_{Q}$ over query set $Q$ and a total of $N$ queries, the access frequency (expected access count) of query $q$ is:
$f_{q}:=N\cdot P_{Q}(q).$ (24)
Thus $f_{q}\in[0,N]$ with $\sum_{q\in Q}f_{q}=N$.
Throughout this subsection we assume Definition 9 has been fixed, and we keep its notation. In particular, we let $P_{Q}$ denote the query distribution over a set $Q$ of admissible queries, and
$f_{q}:=N\cdot P_{Q}(q)$
denote the expected access count of query $q$ over $N$ total queries. [^2] As in Definition 9, $f_{q}$ is a dimensionless quantity: it is a pure count (expected number of accesses) without an intrinsic time unit. When we speak of amortized cost per access, we will divide a total bit-length (storage) or total semantic derivation cost (in nats) by this dimensionless $f_{q}$ to obtain a per-access cost in the corresponding information unit.
We now state the central theorem of this section.
Theorem 5 (Frequency-weighted storage–computation tradeoff).
Let $S_{O}$ be a knowledge base with atomic semantic state set $\text{Atom}(S_{O})$, and let $q$ be a query in the information-rich regime with respect to $S_{O}$ in the sense of Definition 8. Let $P_{Q}$ be a query distribution over a query set $Q$, and define $f_{q}:=N\cdot P_{Q}(q)$ as in Definition 9.
Consider a hybrid strategy $\Pi=(S,C,R)$ where:
  1. 1.
    $S$ is an auxiliary storage structure (e.g., materialized facts, indices, intermediate derivations) that is effectively encodable as a binary string $\text{enc}{S}$ of length $\text{abs}{S}$ bits.
  2. 2.
    $C$ and $R$ describe the residual computation (inference) carried out over the combined knowledge base $S_{O}\cup S$.
We impose the following assumptions on $S$ and on inference over $S_{O}\cup S$:
  1. 1.
    (SA1) Information-theoretic measurement of storage. The bit length $|S|$ of the encoding $\text{enc}{S}$ satisfies
    $H(S)\;\leq\;|S|\;\leq\;H(S)+O(1),$ (25)
    where $H(S)$ is the Shannon entropy of $S$ for a suitable prior over admissible storage states. In particular,
    $I(q;S\mid S_{O})\;\leq\;H(S)\;\leq\;|S|.$ (26)
  2. 2.
    (SA2) Structural regularity of the extended knowledge base. The combined knowledge base $S_{O}\cup S$ admits a well-defined atomic semantic state set $\text{Atom}(S_{O}\cup S)$, a predecessor function satisfying Axioms 1-3, and hence a derivation depth function $Dd(\cdot\mid S_{O}\cup S)$ as in Definition 6. We write
    $m_{S}:=|\text{Atom}(S_{O}\cup S)|,n_{S}:=Dd(q\mid S_{O}\cup S).$ (27)
    We assume $n_{S}=\mathrm{poly}(m_{S})$ (polynomial-depth regime) and that $q$ remains in the information-rich regime with respect to $S_{O}\cup S$ (in the sense of Definition 8, replacing $S_{O}$ by $S_{O}\cup S$). Moreover, the size/depth parameters of $S_{O}\cup S$ are of the same polynomial order as those of $S_{O}$.
  3. 3.
    (SA3) Applicability of the derivation-depth bound. The proof-system/encoding conditions (EPS) and (NPC) of Proposition 1 hold not only for the base theory induced by $S_{O}$, but also for the theory induced by $S_{O}\cup S$. Consequently, the asymptotic relation of Theorem 4 applies to inference from $S_{O}\cup S$: there exist constants $c_{1},c_{2}>0$ such that
    $c_{1}\,\frac{H(q\mid S_{O}\cup S)}{\log(m_{S}+n_{S})}-O(1)\!\;\leq\;\!H_{\mathrm{derive}}(q\mid S_{O}\cup S)\\ \;\leq\;c_{2}\,\frac{H(q\mid S_{O}\cup S)}{\log(m_{S}+n_{S})}+O(1),$ (28)
    where
    $H_{\mathrm{derive}}(q\mid S_{O}\cup S):=Dd(q\mid S_{O}\cup S)\,\lambda$
    is the derivation entropy (Definition 7) with per-step cost $\lambda>0$, and all entropies are measured in nats.
Under these assumptions, define:
  • The semantic information content of $q$ (Definition 8):
    $H_{q}:=H(q\mid S_{O}):=K\big(\text{enc}{q}\mid\text{enc}{\text{Atom}(S_{O})}\big)\,\ln 2.$ (29)
  • The effective derivation cost factor associated with $S$:
    $L_{S}:=\log(m_{S}+n_{S}),$ (30)
    which captures the logarithmic compression advantage available through derivations over $S_{O}\cup S$.
  • The amortized cost per access of answering $q$ using strategy $\Pi$:
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi;q):=\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S).$ (31)
Then there exist positive constants $c,C>0$ (depending only on the universal Turing machine, proof system, and encoding conventions, but not on $S_{O}$, $q$, $S$, or $P_{Q}$) such that
$\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;\min\left\{\frac{H_{q}}{f_{q}},\;C\,\frac{H_{q}}{L_{S}}\right\}+O(1).$ (32)
In particular, up to additive $O(1)$ and multiplicative constant factors, there exists a strategy $\Pi$:
  1. (1)
    Low-frequency regime ($f_{q}\leq L_{S}$, on-demand computation dominated):
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;C\,\frac{H_{q}}{L_{S}}+O(1).$ (33)
  2. (2)
    High-frequency regime ($f_{q}>L_{S}$, precomputation/storage dominated):
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;\frac{H_{q}}{f_{q}}+O(1).$ (34)
The break-even frequency scale is given (up to constant factors) by
$f_{c}\;\asymp\;L_{S}\;=\;\log(m_{S}+n_{S}).$ (35)
Queries with $f_{q}\gg f_{c}$ are asymptotically more cost-effective to support via storage (larger $S$); queries with $f_{q}\ll f_{c}$ are asymptotically more cost-effective to answer via on-demand derivation.
We sketch the main steps.
Step 0: Information-theoretic preliminaries
By (SA1) and the data-processing inequality,
$I(q;S\mid S_{O})\;\leq\;H(S)\;\leq\;|S|.$ (36)
Using the mutual-information identity
$H(q\mid S_{O})=H(q\mid S_{O},S)+I(q;S\mid S_{O}),$ (37)
we obtain
$H(q\mid S_{O},S)=H_{q}-I(q;S\mid S_{O})\geq H_{q}-|S|.$ (38)
Thus $S$ can reduce the conditional entropy of $q$ (given $S_{O}$) by at most $|S|$ nats.
Step 1: Applying Theorem 4
By (SA2) and (SA3), Theorem 4 applies to $S_{O}\cup S$ and yields
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)\;\leq\;c_{2}\,\frac{H(q\mid S_{O}\cup S)}{L_{S}}+O(1).$ (39)
Using the bounds
$H_{q}-|S|\;\leq\;H(q\mid S_{O}\cup S)\;\leq\;H_{q},$ (40)
we obtain
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)\;\leq\;c_{2}\,\frac{H_{q}}{L_{S}}+O\!\left(\frac{|S|}{L_{S}}\right)+O(1).$ (41)
Step 2: Amortized cost
By definition,
$\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)=\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S).$ (42)
Substituting the previous inequality,
$\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;\frac{|S|}{f_{q}}+c_{2}\,\frac{H_{q}}{L_{S}}+O\!\left(\frac{|S|}{L_{S}}\right)+O(1).$ (43)
This decomposition makes explicit the storage-amortization term $|S|/f_{q}$ and the derivation-entropy term $\sim H_{q}/L_{S}$.
Step 3: Extremal strategies
To obtain a clean upper bound in terms of $(H_{q},f_{q},L_{S})$ alone, we compare against two constructive extremes.
(a) Pure computation strategy $\Pi_{\mathrm{comp}}$: take $S=\varnothing$. Then, by Theorem 4 applied to $S_{O}$,
$\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{comp}};q)=H_{\mathrm{derive}}(q\mid S_{O})\\ \;\asymp\;\frac{H_{q}}{\log\big(|\text{Atom}(S_{O})|+Dd(q\mid S_{O})\big)}\\ \;\leq\;C_{0}\,\frac{H_{q}}{L_{S}}+O(1)$ (44)
for some constant $C_{0}>0$, where the last inequality uses that $L_{S}=\log(m_{S}+n_{S})$ is of the same asymptotic order as $\log(|\text{Atom}(S_{O})|+Dd(q\mid S_{O}))$ under the polynomial-depth assumptions.
(b) Pure storage strategy $\Pi_{\mathrm{store}}$: by the coding theorem and the definition of $H_{q}$, there exists a conditional code for the answers to $q$ (given $S_{O}$) of length at most $H_{q}+O(1)$ nats. We can store such a near-optimal representation in $S$, so that $|S|\leq H_{q}+O(1)$, and design $C$ and $R$ so that answering $q$ from $S_{O}\cup S$ requires only constant derivation cost:
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)=O(1).$ (45)
Hence
$\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{store}};q)=\frac{|S|}{f_{q}}+O(1)\;\leq\;\frac{H_{q}}{f_{q}}+O(1).$ (46)
Since we are interested in achievable costs, the optimal amortized cost
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q):=\inf_{\Pi}\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)$ (47)
is bounded above by the minimum of these two explicit strategies:
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\\ \;\leq\;\min\big\{\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{comp}};q),\,\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{store}};q)\big\}\\ \;\leq\;\min\left\{C_{0}\,\frac{H_{q}}{L_{S}},\;\frac{H_{q}}{f_{q}}\right\}+O(1).$ (48)
Equivalently, there exists $C>0$ such that
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\leq\;\max\left\{\frac{H_{q}}{f_{q}},\;C\,\frac{H_{q}}{L_{S}}\right\}+O(1),$ (49)
establishing the claimed bound.
Step 4: Frequency regimes
From the last inequality we obtain two asymptotic regimes:
  • If $f_{q}\leq L_{S}$, then $\frac{H_{q}}{f_{q}}\geq\frac{H_{q}}{L_{S}}$ and
    $\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\leq\;\frac{H_{q}}{f_{q}}+O(1).$ (50)
  • If $f_{q}>L_{S}$, then $\frac{H_{q}}{f_{q}}<\frac{H_{q}}{L_{S}}$ and
    $\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\leq\;C\,\frac{H_{q}}{L_{S}}+O(1).$ (51)
The transition occurs when $f_{q}$ is on the order of $L_{S}$, i.e., $f_{c}\asymp L_{S}=\log(m_{S}+n_{S})$. ∎
Remark 4 (Physical and engineering interpretation).
Theorem 5 admits a natural physical and engineering interpretation.
(1) Units and dimensions. By construction, $H_{q}$ and $H_{\mathrm{derive}}$ are measured in nats (or bits up to a constant factor), while $|S|$ is measured in bits. The frequency $f_{q}$ is a dimensionless expected count of accesses. Therefore:
  • $|S|/f_{q}$ has the same information dimension as $|S|$ (bits), but represents an average storage cost per access;
  • $H_{\mathrm{derive}}$ is a semantic “work” or “action” measured in nats;
  • the bound in Theorem 5 compares two quantities of the same information-theoretic dimension: per-access storage and per-access semantic computation.
In a physical implementation, both can be converted into expected energy dissipation using Landauer’s principle or related models of irreversible computation.
(2) Two limiting protocols. The proof explicitly constructs two limiting protocols:
  1. 1.
    A pure computation protocol, where no auxiliary storage is used ($S=\varnothing$) and the cost is determined entirely by derivation from $S_{O}$; this yields an amortized cost of order $H_{q}/\log(\cdot)$.
  2. 2.
    A pure storage protocol, where we store a near-optimally compressed representation of the answer to $q$, making the derivation cost $O(1)$ and the amortized cost dominated by $H_{q}/f_{q}$.
Theorem 5 states that, up to constant factors and lower-order terms, no hybrid protocol can asymptotically outperform the better of these two extremes, once $H_{q}$, $f_{q}$, and $L_{S}$ are fixed.
(3) Break-even frequency and caching. The critical scale $f_{c}\asymp L_{S}=\log(m_{S}+n_{S})$ plays the role of a break-even frequency between caching and on-demand computation:
  • If $f_{q}\ll f_{c}$, the term $H_{q}/f_{q}$ dominates; it is more efficient to answer $q$ by recomputation each time than to allocate storage to cache its answers.
  • If $f_{q}\gg f_{c}$, the term $CH_{q}/L_{S}$ dominates; investing storage to reduce $H(q\mid S_{O}\cup S)$ (and hence $H_{\mathrm{derive}}$) becomes asymptotically advantageous.
This mirrors classical engineering heuristics for cache and index design: high-frequency queries justify aggressive precomputation and indexing, while low-frequency queries are better handled by direct inference.
(4) Role of $L_{S}$. The factor $L_{S}=\log(m_{S}+n_{S})$ reflects the same logarithmic compression phenomenon as in Theorem 4: in the information-rich regime, the derivation cost per nat of semantic information scales like $1/\log(\text{size}+\text{depth})$. A larger, deeper, but still well-structured theory $S_{O}\cup S$ allows the system to answer semantically complex queries with fewer effective derivation steps per nat.
(5) Robustness. The qualitative form of the tradeoff—and the scale $f_{c}\asymp L_{S}$ at which caching becomes beneficial—is robust to:
  • changes in the underlying universal Turing machine (which affect $H_{q}$ only by an additive constant);
  • variations in the proof system, as long as (EPS) and (NPC) hold and Theorem 4 applies;
  • different encodings of $S$, as long as (SA1) holds, i.e., the code is within $O(1)$ of the Shannon-optimal length.
Thus the storage–computation tradeoff is controlled mainly by the semantic and structural properties of $S_{O}$ and $S$, rather than by low-level representational details.
Remark 5 (Amortized storage cost and query frequencies).
In Theorem 5 we analyzed the amortized cost of serving a single fixed query $q$ that is expected to be accessed repeatedly. Recall that for a query distribution $P_{Q}$ over a finite query set $Q$ and a total horizon of $N$ queries, we define in Definition 9 the expected number of accesses of $q$ by
$f_{q}\;\coloneqq\;N\cdot P_{Q}(q).$
Here $f_{q}$ is a frequency (an expected count), not a probability. Under this convention, the one-time storage investment $|S|$ is amortized over the $f_{q}$ accesses of $q$, leading to the single-query amortized cost
$\operatorname{Cost}_{\mathrm{amort}}(q)\;=\;\frac{|S|}{f_{q}}\;+\;H_{\mathrm{derive}}(q\mid S_{O}\cup S)\\ \;=\;\frac{|S|}{N\cdot P_{Q}(q)}\;+\;H_{\mathrm{derive}}(q\mid S_{O}\cup S).$ (52)
Here $|S|$ is the one-time storage cost (in bits) needed to maintain $S$, and $H_{\mathrm{derive}}(q\mid S_{O}\cup S)$ is the per-access inference cost (in bits) to derive $q$ from $(S_{O},S)$.
A tempting but incorrect way to “average” this cost across query types is to write
$\mathbb{E}_{q\sim P_{Q}}\Bigl[\frac{|S|}{P_{Q}(q)}+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\Bigr]\\ \;=\;\sum_{q_{i}\in Q}P_{Q}(q_{i})\Bigl(\frac{|S|}{P_{Q}(q_{i})}+H_{\mathrm{derive}}(q_{i}\mid S_{O}\cup S)\Bigr).$ (53)
By direct calculation, this yields
$\sum_{q_{i}}P_{Q}(q_{i})\,\frac{|S|}{P_{Q}(q_{i})}\;=\;|S|\cdot|Q|,$
and hence
$\mathbb{E}_{q\sim P_{Q}}\Bigl[\frac{|S|}{P_{Q}(q)}+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\Bigr]\\ \;=\;|S|\cdot|Q|\;+\;\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr].$
This expression is conceptually wrong: it treats $|S|$ as if it were incurred separately for each query type, effectively charging $|S|$ a total of $|Q|$ times. This contradicts the amortization model underlying Theorem 5 and Definition 9, where $|S|$ is a single one-time investment that is shared across all queries.
To see the problem numerically, consider an example where $Q=\{q_{1},q_{2},q_{3}\}$,
$P_{Q}(q_{1})=0.5,\quad P_{Q}(q_{2})=0.3,\quad P_{Q}(q_{3})=0.2,$
the storage cost is $|S|=100$ bits, and
$H_{\mathrm{derive}}(q_{1}\mid S_{O}\cup S)=10,$
$H_{\mathrm{derive}}(q_{2}\mid S_{O}\cup S)=20,$
$H_{\mathrm{derive}}(q_{3}\mid S_{O}\cup S)=30.$
(bits per access). Let $N=1000$ be the total number of queries. Then the expected total inference cost is
$N\cdot\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]$
= $1000\cdot(0.5\cdot 10+0.3\cdot 20+0.2\cdot 30)$
= $17{,}000\text{ bits},$
and the total cost over all queries is
$|S|+17{,}000\;=\;17{,}100\text{ bits}.$
The correct average cost per query is therefore
$\frac{|S|}{N}\;+\;\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]\\ \;=\;\frac{100}{1000}+17\;=\;17.1\text{ bits per query}.$ (54)
In contrast, the naive averaging in (53) yields
$|S|\cdot|Q|+17\;=\;100\cdot 3+17\;=\;317\text{ bits per query},$
which is clearly nonsensical: the one-time cost $|S|$ has been multiplied by the number of query types $|Q|$ instead of being spread over the actual number of query occurrences $N$.
In summary, the correct way to account for storage in a multi-query setting is:
  • treat $|S|$ as a single, one-time storage investment;
  • amortize it over the total number $N$ of query occurrences, giving a per-query storage contribution of $|S|/N$;
  • add the expected per-query inference cost $\mathbb{E}_{q\sim P_{Q}}[H_{\mathrm{derive}}(q\mid S_{O}\cup S)]$.
This yields the expected cost per query as in (54):
$\mathbb{E}[\operatorname{Cost}]\;=\;\frac{|S|}{N}\;+\;\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr].$ (55)
If we now view $Q$ as a random variable with distribution $P_{Q}$, and interpret $H_{\mathrm{derive}}(q\mid S_{O}\cup S)$ as a per-instance algorithmic cost compatible with a Shannon conditional entropy $H(Q\mid S_{O},S)$, then
$\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]\;\approx\;H(Q\mid S_{O},S).$
Consequently, the expected Shannon cost term can be decomposed as
$H(Q\mid S_{O},S)\;=\;H(Q\mid S_{O})-I(S;Q\mid S_{O}),$ (56)
where $I(S;Q\mid S_{O})$ is the conditional mutual information between $S$ and $Q$ given $S_{O}$. Equation (56) shows that the storage $S$ is beneficial to the extent that it captures information about the future queries $Q$ beyond what is already in $S_{O}$: maximizing $I(S;Q\mid S_{O})$ minimizes the residual cost $H(Q\mid S_{O},S)$, and thus reduces the expected inference burden in (55).
Remark 6 (From single-query amortization to system-wide averages).
Remark 5 clarifies that Theorem 5 naturally applies first to a single fixed query $q$ that is accessed $f_{q}=N\cdot P_{Q}(q)$ times. In that setting, the amortized cost per access of $q$ is
$\operatorname{Cost}_{\mathrm{amort}}(q)\;=\;\frac{|S|}{f_{q}}\;+\;H_{\mathrm{derive}}(q\mid S_{O}\cup S),$ (57)
where $|S|/f_{q}$ is the per-access share of the one-time storage cost $|S|$. This formula is meaningful only when $f_{q}$ is understood as a number of accesses (frequency), in accordance with Definition 9.
When we move from this single-query view to a multi-query workload with distribution $P_{Q}$ and horizon $N$, the correct aggregation is:
$\text{total storage cost}\;=\;|S|,$
$\text{total inference cost}\;=\;\sum_{q\in Q}f_{q}\cdot H_{\mathrm{derive}}(q\mid S_{O}\cup S)$
$\;=\;N\cdot\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr],$
$\text{total number of queries}\;=\;\sum_{q\in Q}f_{q}\;=\;N.$
Hence the system-wide average cost per query is exactly
$\mathbb{E}[\operatorname{Cost}]\;=\;\frac{|S|}{N}\;+\;\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr],$ (58)
which is consistent with (55). In particular, the one-time storage cost $|S|$ appears only once in the numerator and is amortized over the $N$ actual query occurrences, rather than over the number of query types $|Q|$.
If, incorrectly, one were to start from the single-query formula (57), replace $f_{q}$ by $N\cdot P_{Q}(q)$, and then naively average over $q$ as in
$\sum_{q_{i}\in Q}P_{Q}(q_{i})\Bigl(\frac{|S|}{P_{Q}(q_{i})}+H_{\mathrm{derive}}(q_{i}\mid S_{O}\cup S)\Bigr),$ (59)
one would again obtain the spurious term $|S|\cdot|Q|$ discussed in Remark 5. This mistake arises from double-counting the storage cost $|S|$ as if it were separately incurred for each query type. The correct interpretation is that $|S|$ is a fixed capital cost that supports the entire query workload, and therefore must be amortized over the full horizon $N$ of query instances.
From an information-theoretic perspective, if we interpret
$\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]\;\approx\;H(Q\mid S_{O},S),$
with $H(\cdot\mid\cdot)$ now denoting the Shannon conditional entropy of the random variable $Q$, then the second term in (58) can be written as
$H(Q\mid S_{O},S)\;=\;H(Q\mid S_{O})-I(S;Q\mid S_{O}),$
where $I(S;Q\mid S_{O})$ is the conditional mutual information between $S$ and $Q$ given $S_{O}$. This identity shows that, at the system level, the storage $S$ is beneficial to the extent that it reduces the Shannon uncertainty of future queries $Q$ beyond what $S_{O}$ already provides. In other words, $S$ is “useful” precisely when it encodes structure about the query distribution $P_{Q}$ so as to increase $I(S;Q\mid S_{O})$ and thereby decrease the expected inference cost term in (58).
We remark that the per-query algorithmic complexity $H_{\mathrm{derive}}(q\mid S_{O}\cup S)$ used in Theorems 15 and the Shannon entropy $H(Q\mid S_{O},S)$ appearing here live at different levels: the former describes instance-wise description length or inference difficulty, while the latter describes average-case uncertainty over the random variable $Q$. They are, however, compatible in the sense that suitable averaging of algorithmic costs over $q\sim P_{Q}$ recovers the Shannon quantities, which is why both viewpoints lead to the same amortized-cost decomposition.
Practical implication: The critical frequency $f_{c}$ provides a quantitative decision rule for capacity allocation in knowledge systems. For a query type $q$ with expected frequency $f_{q}$ over a horizon of $N$ queries, precomputing and storing its answer (or relevant structure) is worthwhile when the one-time storage cost per query $\frac{|S_{q}|}{f_{q}}$ is smaller than the expected per-query inference cost $H_{\mathrm{derive}}(q\mid S_{O})$ (or, in a Shannon view, its expected uncertainty $H(Q\mid S_{O})$). Equivalently, the storage investment is justified when $f_{q}>f_{c}$, where the threshold $f_{c}$ depends on the available capacity and the inference difficulty of $q$.
Concretely, this yields the following design rules in different settings:
  • LLMs (parametric memory). Let $N$ be the number of parameters and suppose storing a fact $q$ in parameters consumes $|S_{q}|$ bits of effective capacity while avoiding $H_{\mathrm{derive}}(q\mid S_{O})$ bits of computation per access. A simple rule of thumb is to store $q$ parametrically only if
    $f_{q}\cdot H_{\mathrm{derive}}(q\mid S_{O})\;\gtrsim\;|S_{q}|.$
    If we model the “effective” per-parameter capacity at the scale of GPT-4 ($N\approx 1.7\times 10^{12}$) by $\log N\approx 40$ bits, this suggests that only facts whose derivation requires on the order of tens of bits of computation and that are accessed at least moderately often should be encoded in the weights. Rare or computationally easy facts are better left to retrieval or external tools.
  • Knowledge graphs (materialized views). For a join pattern $q$ over a graph with $|E|$ edges, let $|S_{q}|$ denote the storage needed for a materialized view and $H_{\mathrm{derive}}(q\mid S_{O})$ the cost of evaluating $q$ on demand. Using the same amortization logic, it is beneficial to materialize when $f_{q}>f_{c}\approx|S_{q}|/H_{\mathrm{derive}}(q\mid S_{O})$. In many practical graph workloads, both $|S_{q}|$ and $H_{\mathrm{derive}}(q\mid S_{O})$ scale only polylogarithmically in $|E|$, which can be crudely summarized as a threshold of the form
    $f_{c}\;\approx\;1+\frac{1}{\log|E|}.$
    For a large knowledge graph such as Freebase ($|E|\approx 10^{9}$), this gives $f_{c}\approx 1.05$, meaning that even slightly more-than-once multi-hop patterns can already justify precomputation.
  • Edge devices (caching). Suppose answering a query from the cloud costs $H_{\mathrm{cloud}}$ and answering it from a local cache costs $H_{\mathrm{local}}$, with a device-to-cloud speed ratio $\eta=H_{\mathrm{local}}/H_{\mathrm{cloud}}$. Caching $q$ locally with storage cost $|S_{q}|$ is beneficial when
    $f_{q}\cdot(H_{\mathrm{cloud}}-H_{\mathrm{local}})\;\gtrsim\;|S_{q}|,$
    i.e., when its expected number of accesses exceeds
    $f_{c}\;\approx\;\frac{|S_{q}|}{H_{\mathrm{cloud}}(1-\eta)}.$
    If we normalize units so that $|S_{q}|/H_{\mathrm{cloud}}\approx 1$ for a typical cached answer, this reduces to the simple rule $f_{q}\gtrsim 1/(1-\eta)$. For a 4G-like setting with $\eta\approx 0.1$, this gives $f_{c}\approx 1.1$ in idealized units; in practice, protocol overheads and limited cache sizes raise the effective threshold to a value on the order of
    $f_{c}\;\approx\;\frac{1}{\eta},$
    meaning only queries expected to be accessed more than $\sim 10$ times per session should reliably be cached locally.
At the system level, these rules are consistent with the global average-cost decomposition
$\mathbb{E}[\operatorname{Cost}]\;=\;\frac{|S|}{N}\;+\;\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr],$
and, in an information-theoretic view,
$\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]\;\approx\;H(Q\mid S_{O},S)\\ \;=\;H(Q\mid S_{O})-I(S;Q\mid S_{O}).$ (60)
In practice, precomputations and storage $S$ are valuable exactly to the extent that they increase $I(S;Q\mid S_{O})$—that is, they capture structure about the future query workload—while their one-time cost $|S|$ is amortized over the total number $N$ of queries.
Corollary 4 (Frequency-weighted extreme strategy costs).
Under the conditions of Theorem 5, fix a query $q$ in the information-rich regime with respect to $S_{O}$ and let $H_{q}:=H(q\mid S_{O})$, $f_{q}:=N\cdot P_{Q}(q)$, and $L_{S}:=\log(m_{S}+n_{S})$ as in (24)–(30). Then the following frequency-weighted cost benchmarks hold (up to additive $O(1)$ and multiplicative constant factors depending only on the universal Turing machine, proof system, and encoding scheme, but not on $S_{O}$, $q$, $S$, or $P_{Q}$):
  1. 1.
    Pure storage (answer caching). There exists a strategy $\Pi_{\mathrm{store}}$ that stores a near-optimally compressed representation of the answer to $q$ of length
    $|S|\;\leq\;H_{q}+O(1),$
    and answers $q$ from $S_{O}\cup S$ with $O(1)$ residual derivation cost:
    $H_{\mathrm{derive}}(q\mid S_{O}\cup S)\;=\;O(1).$
    The resulting amortized cost per access satisfies
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{store}};q)\\ \;=\;\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\;\leq\;\frac{H_{q}}{f_{q}}+O(1).$ (61)
    In particular, whenever $f_{q}\gtrsim L_{S}$ (i.e., $f_{q}$ exceeds the critical frequency scale $f_{c}\asymp L_{S}$ of Theorem 5), this pure-storage protocol matches, up to constant factors, the high-frequency bound
    $\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\lesssim\;C\,\frac{H_{q}}{L_{S}}$
    by making $\frac{H_{q}}{f_{q}}$ the dominant term.
  2. 2.
    Pure computation (no auxiliary storage). Consider the strategy $\Pi_{\mathrm{comp}}$ with $S=\varnothing$. Then, by Theorem 4 applied to $S_{O}$ alone, there exists a constant $C_{0}>0$ such that
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{comp}};q)\;=\;H_{\mathrm{derive}}(q\mid S_{O})\\ \;=\;\Theta\!\left(\frac{H_{q}}{\log\bigl(|\text{Atom}(S_{O})|+\text{Dd}(q\mid S_{O})\bigr)}\right)\\ \;\leq\;C_{0}\frac{H_{q}}{L_{S}}+O(1),$ (62)
    where the last inequality uses that $L_{S}=\log(m_{S}+n_{S})$ is of the same polynomial order as $\log\bigl(|\text{Atom}(S_{O})|+\text{Dd}(q\mid S_{O})\bigr)$ under the structural regularity assumptions (SA2). In particular, when $f_{q}\lesssim L_{S}$, this pure-computation protocol matches, up to constant factors, the low-frequency bound
    $\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\lesssim\;\frac{H_{q}}{f_{q}}.$
  3. 3.
    Near-critical hybrid strategies. For queries with access frequency $f_{q}$ on the order of the critical scale $f_{c}\asymp L_{S}$ (Theorem 5), there exist hybrid strategies $\Pi_{\mathrm{hyb}}=(S,C,R)$ that balance storage and computation so that
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{hyb}};q)\\ \;=\;\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\\ \;=\;H_{q}\Bigl(1-\Theta\!\bigl(\tfrac{1}{L_{S}}\bigr)\Bigr)+o(H_{q}),$ (63)
    i.e., they reduce the effective conditional entropy $H(q\mid S_{O},S)$ to a $1-\Theta(1/L_{S})$ fraction of $H_{q}$ while keeping $\frac{|S|}{f_{q}}$ of order $H_{q}/L_{S}$. Such strategies achieve (up to lower-order terms) the same asymptotic amortized cost as the upper bound in Theorem 5 at $f_{q}\asymp f_{c}$.
  4. 4.
    Frequency stratification (policy-level view). For a query distribution $P_{Q}$ over a finite query set $Q$, the frequency-weighted tradeoff of Theorem 5 induces an asymptotically optimal stratification of queries into three regimes:
    • High-frequency class ($f_{q}\gg f_{c}$): precompute and store (near-)complete answers or dense summaries (pure- or near-pure storage). The expected cost is dominated by
      $\mathbb{E}[\operatorname{Cost}(q)]\;\approx\;\mathbb{E}\Bigl[\tfrac{H_{q}}{f_{q}}\Bigr],$
      i.e., per-query costs decay inversely with frequency.
    • Low-frequency class ($f_{q}\ll f_{c}$): rely on on-demand derivation from $S_{O}$ (pure or near-pure computation). The expected cost is dominated by
      $\mathbb{E}[\operatorname{Cost}(q)]\\ \;\approx\;\mathbb{E}\Bigl[\frac{H_{q}}{\log\bigl(|\text{Atom}(S_{O})|+\text{Dd}(q\mid S_{O})\bigr)}\Bigr],$ (64)
      reflecting the logarithmic computational compression factor of Theorem 1.
    • Near-critical class ($f_{q}\asymp f_{c}$): allocate storage to intermediate facts or partial summaries for $q$ (hybrid strategies), achieving amortized costs of order
      $\mathbb{E}[\operatorname{Cost}(q)]\;\approx\;\mathbb{E}\bigl[H_{q}\bigr]\cdot\Bigl(1-\Theta\!\bigl(\tfrac{1}{L_{S}}\bigr)\Bigr),$
      i.e., constant-factor improvements over naive full recomputation or full storage, with diminishing returns as the system size/depth grows.
Remark 7 (Relationship to the non-frequency-weighted formulation).
The frequency-weighted Theorem 5 refines and generalizes the original storage–computation duality bound that did not explicitly account for access frequencies. In its simplest (single-execution) form, the original bound can be summarized schematically as
$\operatorname{Cost}_{\mathrm{sem}}(\Pi;q)\;:=\;|S|+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\\ \;\gtrsim\;H_{q}\left(1+\frac{1}{c\,\log|\text{Atom}(S_{O})|}\right)-o(H_{q}),$ (65)
for some constant $c>0$, where $\operatorname{Cost}_{\mathrm{sem}}$ is a one-shot semantic cost for answering $q$ (no frequency amortization), and the right-hand side reflects the fact that, in the information-rich regime, derivation depth per nat of semantic content cannot be compressed below a logarithmic factor in the size/depth of the underlying theory.
By contrast, Theorem 5 incorporates the access frequency $f_{q}=N\cdot P_{Q}(q)$ over a horizon of $N$ queries and analyzes the amortized cost per access
$\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;:=\;\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S).$ (66)
It shows that, up to constant factors and lower-order terms,
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\lesssim\;\max\!\left\{\frac{H_{q}}{f_{q}},\,C\frac{H_{q}}{L_{S}}\right\}+O(1),$ (67)
for some constant $C>0$, where the two terms correspond to:
  • a storage-dominated regime $\frac{H_{q}}{f_{q}}$, in which a near-optimal code for $q$ is stored once and reused across $f_{q}$ accesses;
  • a computation-dominated regime $CH_{q}/L_{S}$, in which the answer is derived from $S_{O}\cup S$ on demand, benefiting from the logarithmic compression factor $L_{S}=\log(m_{S}+n_{S})$ identified in Theorem 1.
The critical (break-even) frequency scale is therefore
$f_{c}\;\asymp\;L_{S},$
at which the two contributions $\frac{H_{q}}{f_{q}}$ and $CH_{q}/L_{S}$ are of the same order. Queries with $f_{q}\gg f_{c}$ are asymptotically storage-efficient (caching dominates), while queries with $f_{q}\ll f_{c}$ are asymptotically computation-efficient (on-demand derivation dominates).
The original non-frequency-weighted bound (65) corresponds effectively to the single-execution case $f_{q}=1$. In that regime, the amortized cost per access coincides with the one-shot cost up to constant factors, and the storage term $|S|$ is not meaningfully amortized. In the frequency-weighted setting, by contrast, the same total cost over $M$ executions of $q$,
$\mathbb{E}[\mathrm{Total\ Cost}]\;=\;|S|+M\cdot H_{\mathrm{derive}}(q\mid S_{O}\cup S),$
leads, upon division by $M$, to the amortized form
$\mathbb{E}[\operatorname{Cost}]\;=\;\frac{|S|}{M}+H_{\mathrm{derive}}(q\mid S_{O}\cup S),$
which recovers (66) when $M=f_{q}$. Thus, the frequency-weighted formulation is the natural extension of the original duality bound to long-running systems in which queries are issued multiple times under a distribution $P_{Q}$.
Conceptually, both views share the same structural insight:
  1. 1.
    storage $S$ is a one-time investment that can reduce the conditional algorithmic entropy $H(q\mid S_{O},S)$ (and hence $H_{\mathrm{derive}}$) only to the extent that it encodes information about $q$ beyond $S_{O}$, as quantified by the mutual information $I(S;Q\mid S_{O})$ (cf. Remark 5);
  2. 2.
    derivation cost scales, in the information-rich and polynomial-depth regime, like $H_{q}$ divided by a logarithmic factor in the size/depth of the theory, as captured by Theorem 1;
  3. 3.
    the introduction of $f_{q}$ merely clarifies when it is preferable to pay a larger storage cost to reduce per-access computation, namely when $f_{q}$ exceeds the critical scale $f_{c}\asymp L_{S}$.
In this sense, the non-frequency-weighted and frequency-weighted formulations are compatible limits of the same underlying storage–computation tradeoff: the former optimizes a single execution ($f_{q}=1$), while the latter optimizes repeated executions under a query distribution.
III-C Mutual Information and Submodularity
Theorem 6 (Storage Efficiency via Mutual Information).
The efficiency of storage set $S$ for query distribution $P_{Q}$ is:
$\eta(S)=\frac{I(S;Q)}{|S|}\in[0,1]$ (68)
with equality at 1 if and only if $S$ is a sufficient statistic for $Q$.
Theorem 7 (Submodularity of Mutual Information).
The mutual information $I(S;Q)$ is submodular in $S$:
$I(S_{1}\cup S_{2};Q)+I(S_{1}\cap S_{2};Q)\leq I(S_{1};Q)+I(S_{2};Q)$ (69)
This follows from the submodularity of entropy. For any sets $A,B$:
$H(A\cup B)+H(A\cap B)$ $\leq H(A)+H(B)$ (70)
$\Rightarrow I(A\cup B;Q)+I(A\cap B;Q)$ $\leq I(A;Q)+I(B;Q)$ (71)
by expanding mutual information as I(X;Q)=H(Q)H(Q|X)I(X;Q)=H(Q)-H(Q|X) and using the fact that conditional entropy is supermodular. ∎
Corollary 5 (Greedy Approximation).
A greedy algorithm for selecting storage items achieves a (11/e)(1-1/e) approximation ratio for maximizing $I(S;Q)$ subject to $|S|\leq M_{\max}$.
IV Thermodynamic Constraints and Physical Limits
This section establishes the physical constraints on information processing in knowledge systems, connecting our information-theoretic framework to thermodynamics via Landauer’s principle.
IV-A Landauer’s Principle and Energy Bounds
Theorem 8 (Landauer Energy Bound for Computation).
Under the assumption that each inference step constitutes an ideal logically irreversible operation, deriving formula $\phi$ from knowledge base $S_{O}$ requires minimum energy:
$E_{\text{compute}}(\varphi|S_{O})\geq\text{Dd}(\varphi|S_{O})\cdot k_{B}T\ln(2)$ (72)
where $k_{B}=1.38\times 10^{-23}$ J/K is Boltzmann’s constant and $T$ is the operating temperature.
Each inference step in deriving $\varphi$ from $S_{O}$ performs an irreversible logical operation. By Landauer’s principle [landauer1961irreversibility], each irreversible operation erasing or transforming 1 bit of information requires minimum energy $k_{B}T\ln(2)$.
Consider a derivation with $\text{Dd}(\varphi|S_{O})$ steps. Each step:
  1. 1.
    Applies an inference rule $r_{i}:\{\psi_{1},\ldots,\psi_{k}\}\vdash\psi^{\prime}$
  2. 2.
    Writes conclusion $\psi^{\prime}$ to working memory
  3. 3.
    Erases intermediate results (at least 1 bit per step)
The total bit erasures across all steps is at least $\text{Dd}(\varphi|S_{O})$. By Landauer’s principle, total energy:
$E_{\text{compute}}\geq\text{Dd}(\varphi|S_{O})\cdot k_{B}T\ln(2)$ (73)
Theorem 9 (Storage Energy Bound).
Maintaining storage $S$ containing $|S|$ bits for time $t$ requires energy:
$E_{\text{storage}}(S,t)\geq|S|\cdot k_{B}T\ln(2)\cdot\frac{t}{t_{\text{refresh}}}$ (74)
where $t_{\text{refresh}}$ is the thermal decoherence time.
Physical storage media suffer thermal noise causing bit flips at rate $\lambda_{\text{flip}}\sim\exp(-E_{\text{barrier}}/k_{B}T)$. The refresh period is $t_{\text{refresh}}=1/\lambda_{\text{flip}}$. Each refresh cycle requires re-writing $|S|$ bits, each costing $k_{B}T\ln(2)$ energy. Over time $t$, number of refreshes is $\lceil t/t_{\text{refresh}}\rceil$, yielding the stated bound. ∎
IV-B Energy-Time-Space Triality Bound
We now prove our main physical constraint theorem.
Theorem 10 (Fundamental Triality Bound).
For any hybrid strategy $\Pi=(S,C,R)$ answering query distribution $P_{Q}$ over knowledge base $S_{O}$ in the information-rich regime:
$E(\Pi)\cdot T(\Pi)\cdot M^{-1}(\Pi)\geq\Omega(H(Q|S_{O})\cdot k_{B}T\ln(2))$ (75)
This bound is asymptotically tight and achieved by the optimal hybrid strategy in the information-rich regime (see Section VI).
[Preamble] Since queries are in the information-rich regime:
$H(Q|S_{O})\gg\log_{2}(|\text{Atom}(S_{O})|),\log_{2}(Dd(Q|S_{O}))$ (76)
By Corollary LABEL:cor:Asymptotic_Equivalence, this implies:
$Dd(Q|S_{O})\cdot\ln(2)\approx H(Q|S_{O})\quad[\text{asymptotically}]$ (77)
We analyze three cases using this approximation.
Case 1: Pure Computation Strategy ($\Pi_{c}$)
Setup: - No precomputed storage: $S=\emptyset$, hence $M(\Pi_{c})\to 0$.
- Time cost: Average number of inference steps:
$T(\Pi_{c})=\mathbb{E}_{q\sim P_{Q}}[Dd(q|S_{O})]=Dd(Q|S_{O})$ (78)
- Energy cost: By Theorem 8,
$E(\Pi_{c})=\mathbb{E}_{q\sim P_{Q}}[E_{\text{compute}}(q|S_{O})]\geq Dd(Q|S_{O})\cdot k_{B}T\ln(2)$ (79)
Energy-Time Product: Since we are in the information-rich regime, we use the asymptotic approximation:
$Dd(Q|S_{O})\cdot\ln(2)\approx H(Q|S_{O})$ (80)
Thus,
$\begin{split}E(\Pi_{c})\cdot T(\Pi_{c})&\geq(Dd(Q|S_{O}))^{2}\cdot k_{B}T\ln(2)\\ &\geq\Omega(H(Q|S_{O})^{2}/\ln(2))\cdot k_{B}T\ln(2)\\ &=\Omega(H(Q|S_{O})^{2}\cdot k_{B}T)\end{split}$ (81)
Case 2: Pure Storage Strategy ($\Pi_{s}$)
Setup: - Precompute all answers: $S=\{(q,A(q,S_{O})):q\in\text{supp}(P_{Q})\}$. - Storage requirement:
$M(\Pi_{s})=H(Q|S_{O})$ (82)
- Maintenance energy: By Theorem 9, over average query inter-arrival time $t_{\text{avg}}=1/\lambda_{Q}$:
$E(\Pi_{s})=H(Q|S_{O})\cdot k_{B}T\ln(2)\cdot t_{\text{avg}}\cdot t_{\text{refresh}}$ (83)
Energy-Storage Product:
$E(\Pi_{s})\cdot M^{-1}(\Pi_{s})=k_{B}T\ln(2)\cdot t_{\text{avg}}\cdot t_{\text{refresh}}$ (84)
For reasonable systems where $t_{\text{avg}}\sim t_{\text{refresh}}$:
$E(\Pi_{s})\cdot M^{-1}(\Pi_{s})\geq\Omega(k_{B}T\ln(2))$ (85)
In the information-rich regime, this bound holds as stated.
Case 3: Hybrid Strategy ($\Pi_{h}$)
Setup: - We consider a hybrid strategy that balances both precomputation and on-demand computation.
Analysis: By the Storage-Computation Duality Theorem (Theorem LABEL:thm:duality):
$\min[\text{Storage}]+\min[\text{Computation}]\geq H(Q|S_{O})\cdot(1+\ln(2))$ (86)
In the information-rich regime, the total cost can be approximated as:
$E(\Pi_{h})\cdot T(\Pi_{h})\cdot M^{-1}(\Pi_{h})\geq\Omega(H(Q|S_{O})\cdot k_{B}T\ln(2))$ (87)
This completes the analysis for all three cases, showing consistency with the triality bound.
Corollary 6 (Critical Storage Capacity).
The critical storage capacity is:
$M_{\text{critical}}=\frac{H(Q)}{\log\left(\frac{E_{\text{budget}}}{H(Q|S_{O})\cdot k_{B}T\ln(2)}\right)}$ (88)
Below this capacity, storage is the bottleneck; above it, computation dominates.
IV-C Entropy Production and Irreversibility
Theorem 11 (Minimum Entropy Production).
Answering query $q$ produces minimum environmental entropy:
$\Delta S_{\text{env}}\geq\frac{I(S;q)}{T}$ (89)
By the second law of thermodynamics, total entropy change:
$\Delta S_{\text{total}}=\Delta S_{\text{system}}+\Delta S_{\text{env}}\geq 0$ (90)
The system extracts information $I(S;q)$ from storage $S$. This reduces system entropy by $I(S;q)/\ln(2)$ nats. To maintain non-negative total entropy:
$\Delta S_{\text{env}}\geq\frac{I(S;q)}{T}$ (91)
V Optimal Hybrid Strategies
V-A Threshold Policies for Hybrid Strategies
Theorem LABEL:thm:duality implies that the marginal benefit of storing additional information diminishes as the knowledge base grows. For a query $q$ with access frequency $f_{q}$, the expected cost of computing on-demand is
$\mathbb{E}[\text{Compute Cost}]=f_{q}\cdot H_{\mathrm{derive}}(q\mid S_{O})\\ =\Theta\!\left(f_{q}\cdot\frac{H_{q}}{\log|\text{Atom}(S_{O})|}\right),$ (92)
while the amortized cost of storing its answer is
$\mathbb{E}[\text{Store Cost}]=H_{q}+\lambda\cdot\text{(storage overhead)},$ (93)
where $\lambda$ penalizes storage capacity usage.
A query should be stored if and only if
$f_{q}\cdot\frac{H_{q}}{\log|\text{Atom}(S_{O})|}>H_{q}\quad\Longleftrightarrow\quad f_{q}>\Theta(\log|\text{Atom}(S_{O})|).$ (94)
Thus, we obtain a scale-aware threshold policy:
Policy 1 (Scale-Aware Threshold).
Store query $q$ if its normalized frequency satisfies
$f_{q}\cdot\frac{Dd(q\mid S_{O})}{H_{q}}>\tau,$ (95)
where the threshold $\tau=\Theta(\log|\text{Atom}(S_{O})|)$ grows logarithmically with knowledge base size.
This policy generalizes classical caching (e.g., LFU) by incorporating computational depth $Dd(q|S_{O})$ and semantic information $H_{q}$, and explains the phase transition in storage utility observed empirically.
V-B Phase Transition
Theorem 12 (Storage Utility Phase Transition).
Define $\alpha=E_{\text{storage}}/E_{\text{compute}}$. There exists critical point
$\alpha_{c}=\frac{H(Q)}{\mathbb{E}[\text{Dd}(Q|S_{O})]\cdot\ln(2)}$ (96)
such that: (i) for $\alpha<\alpha_{c}$, $\partial L/\partial M\sim-\exp(-\alpha/\alpha_{c})$; (ii) for $\alpha>\alpha_{c}$, $\partial L/\partial M\sim-1/M$.
VI Numerical Validation and Simulations
All datasets satisfy $H(q\mid S_{O})\gg\log|\mathrm{Atom}(S_{O})|$ (verified empirically), placing them firmly in the information-rich regime. We validate our theoretical predictions through extensive numerical experiments on synthetic knowledge systems. Knowledge bases range from 1K to 100K atoms. Query distributions include uniform and Zipfian ($\alpha\in\{1.0,1.2,1.5,1.8,2.0\}$). Physical parameters are: temperature T=300T=300K.
In Experiment 1, we validate Theorem LABEL:thm:duality (Frequency-Weighted Duality) by testing individual queries sampled from realistic distributions. For each query $q$, we compute its access frequency $f_{q}$ from a Zipfian distribution, then verify that the optimal amortized cost satisfies the theoretical bound $\text{Cost}_{\text{amort}}(q)\geq H_{q}(1+\frac{1}{c\log|K|}-\frac{1}{f_{q}})$. We test 50 randomly sampled queries from each of three knowledge bases: Small (1K atoms), Medium (10K atoms), Large (100K atoms), with average logical depths 3–5.
TABLE II: Single-Query Duality Theorem Validation (Theorem LABEL:thm:frequency_weighted_duality)
KB Size Entities # Queries Sat. Rate Avg. Margin
Small 1.00K 50 100.0% 6,785.7
Medium 10.00K 50 100.0% 280,828.3
Large 100.00K 50 100.0% 10,329,841.7
Note: Sat. Rate = fraction of queries satisfying $\mathrm{Cost}_{\mathrm{amort}}(q)\geq H_{q}\bigl(1+\frac{1}{c\log|K|}-\frac{1}{f_{q}}\bigr)$. Avg. Margin = average difference between observed cost and theoretical bound.
The results demonstrate perfect validation of Theorem 5: all 150 tested queries (50 per knowledge base) satisfy the frequency-weighted duality bound. The large positive margins (6.8K–10.3M nats) indicate that the bound is not tight in practice, which is expected since the theorem provides a fundamental lower limit rather than an achievable optimum. The increasing margin with knowledge base size reflects the $1/f_{q}$ term in the bound—larger knowledge bases have more queries with lower individual frequencies, leading to more conservative bounds.
[…]
Physical Interpretation: At room temperature with $H(Q|S_{O})=10^{6}$ bits, the fundamental limit is:
$E\cdot T\cdot M^{-1}\geq 10^{6}\times 1.38\times 10^{-23}\times 0.693\approx 10^{-17}\text{ J}\cdot\text{s/bit}$ (97)
In Experiment 2, we vary storage capacity from 0
Figure 1: Phase transition in query latency. The figure shows: (top-left) latency vs. storage capacity, (top-right) gradient $\partial L/\partial M$ marking the phase transition signature, (bottom-left) cache hit rate, and (bottom-right) latency vs. storage in bits on log scale. The red dashed vertical lines mark the transition point at approximately 10% storage capacity, where latency improvement shifts from rapid to gradual, demonstrating the phase transition phenomenon predicted by theory.
Results (Figure 1): The phase transition occurs at approximately 10% of the query set, identified by the inflection point in the latency gradient. Below this point, increasing storage yields rapid latency improvement (steep gradient); above it, further storage provides diminishing returns (gradual gradient). The transition validates the theoretical prediction that storage utility undergoes a qualitative shift at a critical capacity determined by information-theoretic and workload parameters.
In Experiment 3, we compare our FreqDepth strategy ($\text{score}=\text{freq}\times\text{depth}$) against three baselines across cache sizes 10-500 on the medium KB (1000 queries). We focus on the constrained regime (cache=50, 5% of queries) where strategy differences are most pronounced.
TABLE III: Baseline Comparison at Cache Size = 50 (5% of queries)
Strategy Hit Rate Latency Improv.
LRU 87.2% 1.54 baseline
LFU 87.9% 1.50 2.6%
TrueMI 51.6% 2.96 -92%
FreqDepth (Ours) 91.5% 1.36 11.6%
Figure 2: Baseline strategy comparison across cache sizes (10-500). The three panels show: (left) cache hit rate, (center) average query latency, and (right) total computation cost. Our FreqDepth strategy (freq $\times$ depth heuristic) consistently outperforms LRU and LFU across all metrics and cache sizes. TrueMI (mutual information maximization) performs poorly due to ignoring access frequency.
At the constrained cache size of 50 (Table 2), our FreqDepth strategy achieves 91.5% hit rate vs. 87.2% for LRU, with 11.6% latency improvement. Figure 2 shows performance across the full range (cache sizes 10-500): FreqDepth consistently outperforms LRU/LFU at all sizes. The theoretically optimal TrueMI performs poorly (51.6% hit rate at cache=50) because it greedily selects queries that maximize global mutual information without considering access frequency—it caches informationally rich but rarely accessed queries. In contrast, FreqDepth’s $\text{freq}\times\text{depth}$ heuristic balances information gain with access probability, achieving practical superiority. At larger cache sizes (e.g., cache $\geq$ 300), all classical strategies converge to $95\%$ hit rates as capacity constraints relax.
In Experiment 4, we test robustness via sensitivity analysis across: (i) five Zipf exponents $\alpha\in\{1.0,1.2,1.5,1.8,2.0\}$ modeling query frequency distributions; (ii) five average logical depths $\{2,3,5,7,10\}$; (iii) four knowledge base scales from 100 to 5000 entities:
Figure 3: Sensitivity analysis. Left pair: cost vs. storage fraction for different Zipf parameters $\alpha$ and optimal $\beta^{*}$ vs. $\alpha$. Right pair: cost vs. storage fraction for different logical depths and optimal $\beta^{*}$ vs. depth. Theoretical bounds hold across all configurations.
[…]
Figure 4 shows that $H(Q)$ scales logarithmically with knowledge base size as predicted by information theory. The framework’s bounds hold consistently across all tested configurations, validating theoretical predictions across diverse parameter regimes.
VII Applications
We demonstrate practical applications of our framework across three domains.
VII-A Large Language Models
Consider a GPT-4 scale model with 1.7T parameters. The knowledge base contains $H(S_{O})=10^{15}$ bits, and queries follow a power-law distribution with $H(Q)=10^{12}$ bits.
Parameter Settings. Average logical depth: $\mathbb{E}[\text{Dd}]=5$ inference steps. Energy budget: $E_{\text{budget}}=1000$ kWh. Temperature: T=300T=300K.
Critical Storage Calculation. Using Corollary 6:
$M_{\text{critical}}=\frac{10^{12}}{\log(1000\times 3.6\times 10^{9}/(10^{6}\times k_{B}T\ln 2))}\\ \approx 2.5\times 10^{11}\text{ bits}$ (98)
Optimal Strategy. Following our threshold policy framework, we compute $\text{Dd}^{*}=2.5$ and $f^{*}=0.85$. The optimal strategy:
  • Parametric storage (70%): Facts with $\text{Dd}\leq 2$ and frequency >0.85>0.85
  • RAG retrieval (25%): Facts with $3\leq\text{Dd}\leq 8$
  • Dynamic reasoning (5%): Complex queries with $\text{Dd}>8$
Performance Gains. Compared to pure parametric storage: 45% energy reduction, with latency distribution P50=50ms, P90=200ms, P99=2000ms.
VII-B Knowledge Graphs
A medical knowledge graph with $10^{7}$ entities and $10^{8}$ relationships faces multi-hop query workloads.
Query Analysis.
TABLE IV: Query Patterns in the Biomedical KB
Query Type Dd Freq. Example
Single-hop 1 40% Drug $\to$ Disease
Double-hop 2 35% Symptom $\to$ Treatment
Triple-hop 3 20% Patient $\to$ Protocol
Complex $\geq 4$ 5% Multi-condition
[…]
VII-C Edge Intelligence
A smartphone assistant with 4GB RAM and 15Wh battery must operate for 24 hours.
Resource Allocation. Energy budget per 24h: $E_{\text{budget}}=4.5$ Wh (30% to AI). Storage limit: $M_{\max}=2$ GB (50% to knowledge).
Hybrid Strategy. Device storage (200 MB): Common sense, personal habits, quick facts. Cloud coordination: Low-frequency knowledge and complex reasoning ($\text{Dd}>10$). Adaptive caching: Weekly updates based on usage patterns.
Energy Distribution. Device inference (40%), network communication (30%), storage maintenance (20%), overhead (10%). Achieves 95% on-device query completion within 24h battery constraint.
VIII Related Work
VIII-A Information Theory and Kolmogorov Complexity
Shannon’s seminal 1948 paper [shannon1948mathematical] introduced entropy $H(X)=-\sum_{x}p(x)\log p(x)$ as the fundamental measure of information content, establishing the mathematical foundations for data compression and communication. This framework has been extensively developed, with Cover and Thomas [cover2012elements] providing the standard modern treatment encompassing source coding, channel capacity, and rate-distortion theory.
Parallel to Shannon’s probabilistic approach, algorithmic information theory developed the notion of descriptional complexity. Kolmogorov [kolmogorov1965three], Solomonoff [solomonoff1964formal], and Chaitin [chaitin1969simplicity] independently introduced Kolmogorov complexity $K(x)$ as the length of the shortest program generating string $x$. This captures a different facet of information: while Shannon entropy measures average information content over a distribution, Kolmogorov complexity measures the individual information content of a single object.
The minimum description length (MDL) principle [rissanen1978modeling, grunwald2007minimum] applies these ideas to model selection and inference, balancing model complexity against data fit. However, MDL focuses on static description length, not the dynamic computational cost of deriving consequences.
Bennett’s logical depth [bennett1988logical] addresses this gap by measuring computational sophistication: $\text{Dd}(x)$ is the runtime of the fastest program generating $x$ from a short input. This distinguishes between ”shallow” information (random bits, easy to describe but not to compute from scratch) and ”deep” information (organized structures requiring significant computation to generate).
Our work extends Bennett’s framework in three critical ways:
  1. 1.
    Conditional depth: We formalize $\text{Dd}(\varphi|\psi)$ for conditional derivation, enabling analysis of incremental inference.
  2. 2.
    Information-theoretic duality: We prove that derivation entropy $H_{\text{derive}}=\text{Dd}\cdot\ln(2)$ is dual to Shannon entropy, establishing a precise quantitative relationship.
  3. 3.
    Physical grounding: We connect logical depth to thermodynamic energy costs via Landauer’s principle, providing physical constraints on computation.
Related notions include computational depth [antunes2006sophistication], which measures the difficulty of generating an object from random bits, and logical depth sophistication [koppel1987structure], which captures ”useful” information. However, these works focus on individual objects rather than knowledge systems, and lack our connection to storage-computation trade-offs and thermodynamic limits.
VIII-B Thermodynamics of Computation
Landauer’s landmark 1961 paper [landauer1961irreversibility] established the first deep connection between information theory and thermodynamics, proving that erasing one bit of information requires dissipating minimum energy $E_{\text{erase}}\geq k_{B}T\ln(2)$ where $k_{B}$ is Boltzmann’s constant and $T$ is temperature. This result has profound implications: computation is fundamentally irreversible and energy-consuming at the physical level.
Bennett extended this foundation in several directions. His work on reversible computation [bennett1973logical] showed that logically reversible operations can be performed with arbitrarily small energy dissipation, though at the cost of increased time and space. His comprehensive 1982 review [bennett1982thermodynamics] synthesized the thermodynamics of computation, connecting Landauer’s principle to Maxwell’s demon, the second law, and computational complexity.
Recent years have seen experimental confirmation of Landauer’s principle. Bérut et al. [berut2012experimental] demonstrated bit erasure approaching the Landauer limit using a colloidal particle in optical tweezers. Theoretical refinements [sagawa2010thermodynamic] have connected information thermodynamics to feedback control and Maxwell’s demon.
Lloyd [lloyd2000ultimate] derived ultimate physical limits to computation from quantum mechanics, showing that a system of mass $m$ and energy $E$ confined to volume $V$ can perform at most $E/(h\pi)$ operations per second and store at most $S=k_{B}\ln d$ bits where $d$ is the dimensionality of the Hilbert space.
Our work makes three novel contributions beyond existing thermodynamics:
  1. 1.
    Logical depth-energy connection: We prove that each inference step requires energy $k_{B}T\ln(2)$ by connecting logical depth to bit erasure.
  2. 2.
    Triality bound: We derive the first joint bound on energy, time, and storage: $E\cdot T\cdot M^{-1}\geq\Omega(H(Q|S_{O})\cdot k_{B}T\ln(2))$. Previous work bounded resources independently; our triality captures fundamental coupling.
  3. 3.
    System-level analysis: Rather than isolated gates or operations, we analyze end-to-end knowledge system workflows, accounting for storage maintenance, query routing, and dynamic workloads.
This shift from gate-level to system-level thermodynamics enables practical application to modern computing systems (LLMs, knowledge graphs, edge devices) where architectural decisions dominate energy consumption.
VIII-C Time-Space Trade-offs in Computation
Classical computational complexity theory has extensively studied time-space trade-offs. Yao [yao1985should] pioneered this area with results on sorting and searching. Beame et al. [beame1991general] provided general sequential time-space trade-offs showing that for many problems, $T(n)\cdot S(n)=\Omega(n^{2})$ where $T$ is time and $S$ is space. Savage’s work on data structures [arora2009computational] established fundamental limits on pointer-based representations.
Competitive analysis of caching. Sleator and Tarjan’s seminal work on competitive paging algorithms established that LRU achieves $O(k)$-competitive ratio where $k$ is cache size. However, competitive analysis focuses on worst-case performance against adversarial inputs, whereas our information-theoretic framework characterizes average-case performance under probabilistic query distributions. Recent work on learning-augmented algorithms [kraska2018case] attempts to bridge this gap using machine learning, but lacks theoretical guarantees.
Our work differs by: (i) adopting an information-theoretic perspective quantifying information content rather than worst-case complexity, (ii) incorporating thermodynamic energy costs via Landauer’s principle, (iii) optimizing average-case performance over query distributions using logical depth, and (iv) providing asymptotic bounds rather than competitive ratios. The constant $(1+\ln(2))$ emerges from information-theoretic first principles rather than algorithm-specific analysis.
VIII-D Knowledge Base and Caching Systems
Materialized Views. Database systems extensively use materialized views to trade storage for query latency [gupta1997materialized, chaudhuri1995optimizing]. These systems employ cost-based optimizers but lack theoretical guarantees. Our framework provides the first information-theoretic lower bounds on this trade-off.
Caching Policies. Classical caching algorithms like LRU, LFU, and ARC [megiddo2003arc] rely on heuristics based on recency and frequency. While effective, they provide no optimality guarantees. Our threshold strategy achieves provable (11/e)(1-1/e) approximation via submodular optimization [nemhauser1978analysis, krause2014submodular].
Learned Systems. Recent work applies machine learning to database optimization [kraska2018case], exploiting data patterns. These approaches are complementary to our theoretical framework, which provides fundamental limits that learned systems cannot surpass.
VIII-E Knowledge Graphs and Semantic Systems
Large-scale knowledge graphs [bollacker2008freebase, carlson2010toward] and semantic networks [miller1995wordnet] face significant storage-computation challenges. Embedding methods [bordes2013translating] compress knowledge at the cost of inexact retrieval. Our framework quantifies this trade-off: compression reduces storage entropy but increases logical depth for reconstruction.
Modern question-answering systems combine parametric knowledge (stored in model weights) with non-parametric retrieval [karpukhin2020dense, lewis2020retrieval]. This hybrid approach implicitly navigates the storage-computation trade-off. Our theoretical analysis provides principled guidelines for allocating capacity between these components.
VIII-F Large Language Models and Retrieval
Transformer-based language models [vaswani2017attention, brown2020language] store knowledge parametrically, requiring massive storage and computation. Retrieval-augmented generation (RAG) [lewis2020retrieval] reduces this burden by offloading factual knowledge to external databases.
Our analysis reveals that optimal LLM architectures should store high-frequency, high-logical-depth knowledge parametrically, while retrieving low-frequency or simple facts. This explains the empirical success of RAG: it exploits the storage-computation duality by allocating expensive parametric storage to knowledge with high $\text{Dd}(q|S_{O})\times\text{freq}(q)$ product.
Recent systems like ZeRO [rajbhandari2020zero] and distributed training frameworks [zhao2020pytorch] address memory limitations in large models. Our triality bound provides theoretical justification for their design choices.
VIII-G Energy-Efficient Computing
The energy crisis in computing [horowitz20141, barroso2007case] has motivated extensive work on energy-proportional systems. Prior work optimizes specific components (processors, memory, networks) independently. Our triality bound reveals fundamental coupling between energy, time, and storage that cannot be overcome through component-level optimization alone.
Sub-threshold computing [lei2017optimal] reduces energy per operation but increases latency. Our framework quantifies this trade-off through the triality product, showing that gains in energy efficiency must be balanced against increased time or storage costs.
VIII-H Phase Transitions in Computer Science
Phase transitions in computational complexity [monasson1999determining] and constraint satisfaction [krzakala2007gibbs] reveal sudden changes in problem difficulty. Our phase transition at critical storage capacity $M_{c}$ is analogous but fundamentally different: it arises from information-theoretic constraints rather than problem structure.
The transition we observe is sharp and universal across problem domains, determined solely by $H(Q)$, $\mathbb{E}[\text{Dd}]$, and physical constants. This universality distinguishes it from problem-specific phase transitions in SAT solvers or random graphs.
VIII-I Information-Theoretic Learning and Prediction
Information-theoretic approaches to learning have a rich history. Merhav and Feder [merhav1998universal] developed universal prediction schemes achieving min-max regret bounds under entropy constraints. Their work focuses on passive learning where the learner observes data and predicts. In contrast, our framework addresses active computation where the system decides whether to retrieve precomputed results or compute on-demand.
The connection arises through regret minimization: our storage selection problem is equivalent to minimizing regret in an online learning game where ”experts” are candidate storage sets. However, standard online learning assumes unit costs per expert consultation, whereas in our setting, computation cost scales with logical depth $\text{Dd}(q|S)$, requiring depth-aware algorithms.
VIII-J Submodularity and Greedy Algorithms
Submodular optimization [nemhauser1978analysis, krause2014submodular, buchbinder2018submodular] provides approximation guarantees for many combinatorial problems. We prove that mutual information $I(S;Q)$ is submodular in storage set $S$, enabling (11/e)(1-1/e) approximation via greedy selection [badanidiyuru2014fast].
This connection is non-trivial: while entropy is submodular, the optimal storage set must balance multiple objectives (coverage, diversity, computation reduction), requiring careful analysis. Recent advances in submodular maximization [buchbinder2018submodular] could potentially improve our greedy algorithms, though the (11/e)(1-1/e) bound is optimal for general submodular functions.
VIII-K Online Learning and Adaptive Strategies
For dynamic query distributions, we develop online algorithms achieving $O(\sqrt{T\log n})$ regret. This connects to online convex optimization [hazan2016introduction] and multi-armed bandits [auer2002nonstochastic, cesa2006prediction].
Unlike classical online learning where actions affect only immediate reward, storage decisions have long-term effects through reduced computation. We address this via a novel reduction to online submodular maximization.
IX Conclusion
In this paper, we have established a unified physical framework for intelligence, grounded in the fundamental assumption that information constitutes the enabling mapping from ontological states to carrier states [xu2025information, xu2024research]. By rigorously analyzing the thermodynamic costs associated with this mapping, we have demonstrated that the separation between ”static storage” (Shannon Entropy) and ”dynamic computation” (Logical Depth) is artificial; they are, in fact, conjugate variables governed by a fundamental constraint: the Energy-Time-Space Triality Bound.
Our proposed metric, Derivation Entropy ($H_{derive}$), serves as the quantitative bridge between these domains. It reveals that the efficiency of any intelligent system—biological or artificial—is strictly bounded by the trade-off between the entropic cost of maintaining memory and the dissipative heat of performing logical derivation [landauer1961irreversibility, bennett1982thermodynamics]. Specifically, we identified a Critical Derivation Threshold, a phase transition point beyond which it becomes thermodynamically cheaper to re-compute information from compressed laws rather than to retrieve it from vast databases.
This theoretical finding has profound implications for the design of next-generation AI architectures. As current models like Transformers approach the limits of memory bandwidth and energy consumption [vaswani2017attention, horowitz20141], our framework suggests that future systems may benefit from optimizing the Derivation-Retrieval Ratio rather than simply scaling parameters. Such systems could dynamically navigate the ”Logical Depth vs. Shannon Entropy” curve, acting as adaptive engines that minimize the total action of the ontological mapping process.
Ultimately, this work suggests that the ”simplicity” of a scientific theory or an AI model is not just an aesthetic preference but may reflect a thermodynamic imperative. By aligning the formalization of general states [qiu2025research] with physical constraints, we provide a roadmap for constructing intelligent systems that are not only logically powerful but physically sustainable. Future research will extend this framework to non-equilibrium open systems and quantum computing regimes, further exploring the ultimate physical limits of the enabling mapping.
Acknowledgment
During the writing and revision of this paper, we received many insightful comments from Associate Professor Wang Rui of the School of Computer Science at Shanghai Jiao Tong University and also gained much inspiration and assistance from regular academic discussions with doctoral students Yiming Wang, Chun Li, Hu Xu, Siyuan Qiu, Jiashuo Zhang, Junxuan He, and Xiao Wang. We hereby express our sincere gratitude to them.