### v4

# Information Physics of Intelligence: Unifying Derivation Depth and Entropy under Thermodynamic Constraints

Abstract
The rapid scaling of artificial intelligence models has revealed a fundamental tension between model capacity (storage) and inference efficiency (computation). While classical information theory focuses on transmission and storage limits, it lacks a unified physical framework to quantify the thermodynamic costs of generating information from compressed laws versus retrieving it from memory. In this paper, we propose a theoretical framework that treats information processing as an enabling mapping from ontological states to carrier states. We introduce a novel metric, Derivation Entropy, which quantifies the effective work required to compute a target state from a given derivation depth. By analyzing the interplay between Shannon entropy (storage) and computational complexity (time/energy), we demonstrate the existence of a critical phase transition point. Below this threshold, memory retrieval is thermodynamically favorable; above it, generative computation becomes the optimal strategy. This ”Energy-Time-Space” conservation law provides a physical explanation for the efficiency of generative models and offers a rigorous mathematical bound for designing next-generation, energy-efficient AI architectures. Our findings suggest that the minimization of Derivation Entropy is a governing principle for the evolution of both biological and artificial intelligence.
I Introduction
The fundamental nature of information has long been a subject of debate across physics, computer science, and philosophy. In our previous works, we established a formalized meta-framework positing that information is fundamentally the enabling mapping from an ontological state to a carrier state [xu2025information, xu2024research, qiu2025research]. Under this view, intelligence is not merely the accumulation of data, but an active physical process of mapping the infinite complexity of the external environment (ontology) onto the finite constraints of an internal representation (carrier). This mapping inherently necessitates a primal trade-off: the system must decide whether to store the mapping results statically (memory) or to regenerate them dynamically via algorithms (computation).
Classically, these two aspects have been treated separately. Shannon’s information theory [shannon1948mathematical] rigorously defines the limits of static storage and transmission, quantifying the irreducible randomness of the carrier state. However, Shannon entropy is agnostic to the causal history or the computational effort required to generate that state. Conversely, Algorithmic Information Theory (AIT), particularly through the lens of Kolmogorov complexity [kolmogorov1965three] and Bennett’s concept of Logical Depth [bennett1988logical], focuses on the computational time required to reconstruct a structure from its compressed description. While Derivation Depth captures the dynamic value of the ”mapping process,” it lacks a direct unification with the statistical mechanics governing the carrier’s physical existence.
This separation becomes problematic when we consider that information processing is strictly bound by physical laws. As Brillouin [brillouin1956science] and Landauer [landauer1961irreversibility] demonstrated, the manipulation of information—specifically the erasure of logical states—is inseparably linked to thermodynamic entropy and heat dissipation. Therefore, the ”enabling mapping” from ontology to carrier is not just a logical operation but a thermodynamic cycle. Current artificial intelligence systems, such as Large Language Models (LLMs) based on the Transformer architecture [vaswani2017attention, brown2020language], are rapidly approaching the physical limits of this cycle. The empirical scaling laws observed in these systems suggest a deep, underlying fundamental trade-off connecting parameter size (storage), training FLOPs (computation), and energy consumption, yet a unified theoretical explanation remains elusive.
In this paper, we propose the ”Information Physics of Intelligence,” a theoretical framework that unifies Derivation Depth and Shannon Entropy under rigorous thermodynamic constraints. We define Derivation Entropy ($H_{derive}$), which quantifies the total physical cost of the ontological-to-carrier mapping. By integrating the thermodynamic cost of storage with the algorithmic complexity of retrieval, we derive the Energy-Time-Space Triality Bound. This bound reveals that minimizing storage space (lowering Shannon entropy) inevitably necessitates a compensatory increase in computational time (increasing Derivation Depth) or energy flux, governed by Landauer’s limit.
Our contributions are threefold:
  1. 1.
    We formalize the ”ontological-to-carrier mapping” within a thermodynamic context, bridging the gap between abstract information measures [xu2025general] and physical complexity [bennett1982thermodynamics].
  2. 2.
    We mathematically derive the phase transition point where intelligent systems must switch from a retrieval-based strategy to a derivation-based strategy to maintain thermodynamic optimality.
  3. 3.
    We provide a first-principles explanation for the efficiency bottlenecks in modern AI models, offering a governing equation for designing next-generation, energy-efficient cognitive architectures.
II System Model and Preliminaries
We establish the formal framework for knowledge systems, introducing the concepts of enabling mappings, derivation depth, and their information-theoretic characterization.
II-A The Underlying Logical System $\mathcal{L}$
Throughout this work, we fix a formal logical system $\mathcal{L}$ that serves as the foundation for expressing semantic states and reasoning about information. Specifically, $\mathcal{L}$ is taken to be first-order logic with least fixed-point operators (FO(LFP)) [immerman1999descriptive], extended with multiple sorts, including at least $\mathsf{Obj}$ (for ontological entities), $\mathsf{Time}$ (for discrete time points), and $\mathsf{Carrier}$ (for physical carriers); function and relation symbols are typed accordingly (e.g., a state predicate may have signature $\mathsf{State}:\mathsf{Obj}\times\mathsf{Time}\to\mathsf{Bool}$). The language of $\mathcal{L}$ includes:
  • A countable set of sorted variables, constants, function symbols, and relation symbols;
  • Standard Boolean connectives and first-order quantifiers over each sort;
  • A least fixed-point operator $\mu$ enabling inductive definitions of relations (e.g., reachability, state transitions).
Crucially, we restrict attention to finite structures over a discrete, bounded time domain $T=\{t_{0},t_{1},\dots,t_{n}\}$ for some $n\in\mathbb{N}$. Under this restriction, $L$ enjoys the following essential properties:
  • Consistency: The proof system of $\mathcal{L}$ is sound and free of contradiction.
  • Effective model checking: For any finite structure $\mathcal{M}$ and formula $\varphi\in\mathcal{L}$, the satisfaction relation $\mathcal{M}\models\varphi$ is decidable.
  • Expressive adequacy: By the Immerman–Vardi theorem [immerman1986relational, vardi1982complexity], FO(LFP) captures exactly the class of polynomial-time computable properties over ordered finite structures. This makes $\mathcal{L}$ sufficiently expressive to model the state evolution of a wide range of real-world information systems under finite-horizon autoregressive decoding.
II-B Enabling Mappings and Information
Definition 1 (State Expressibility).
Let $\mathcal{L}$ be the logical system defined above. The state set $S(X,T)$ of an object set $X$ over a time set $T$ is said to be expressible in $\mathcal{L}$ if it can be represented by assigning $X$, $T$, and their subsets and elements as interpretations of a set of well-formed formulas (WFFs) in $\mathcal{L}$.
In all subsequent discussions of this paper, we consider only states expressible in the logical system $\mathcal{L}$[qiu2025research, xu2025information].
Definition 2 (Enabling Mapping).
Let $O$ be an ontological space, $C$ a carrier space, and $T_{o},T_{r}\subseteq\mathbb{R}^{+}$ be time domains. A correspondence
$\mathcal{E}:S_{O}(O,T_{o})\rightrightarrows S_{C}(C,T_{r})$ (1)
is called an enabling mapping if the following conditions hold:
(E1) Realizability:
For every ontological state $s_{o}\in S_{O}(O,T_{o})$, the image $\mathcal{E}(s_{o})$ is a non-empty subset of $S_{C}(C,T_{r})$. That is, there exists at least one valid carrier state $s_{c}\in\mathcal{E}(s_{o})$ that can physically represent $s_{o}$.
(E2) Causality:
For every $s_{o}$ and any realization $s_{c}\in\mathcal{E}(s_{o})$, we have $s_{o}\prec s_{c}$, where $\prec$ denotes a causal dependency relation, implying: (i) The occurrence time of $s_{o}$ precedes or is concurrent with the realization of $s_{c}$; (ii) The ontological state $s_{o}$ is a necessary condition for the intentional generation of $s_{c}$.
(E3) Physicality:
The realization of any mapping instance $s_{o}\mapsto s_{c}$ (where $s_{c}\in\mathcal{E}(s_{o})$) consumes non-zero physical resources (e.g., energy, time).
Remark 1.
[…]
Definition 3 (Information).
Information is an enabling mapping [xu2024research]:
$\mathcal{I}:S_{O}(O,T_{o})\rightrightarrows S_{C}(C,T_{r})$ (2)
where
  • $O$: ontology (objective or subjective reality)
  • $T_{o}\subseteq\mathbb{R}^{+}$: occurrence time
  • $S_{O}=S_{O}(O,T_{o})$: state set of ontology
  • $C$: carrier (objective reality)
  • $T_{r}\subseteq\mathbb{R}^{+}$: reflection time
  • $S_{C}=S_{C}(C,T_{r})$: state set of carrier
Equivalently, information can be represented as a sextuple $\mathcal{I}=\langle O,T_{o},S_{O},C,T_{r},S_{C}\rangle$. It is also assumed to be effectively computable on semantic inputs
Definition 4 (Ideal Information).
Let $S_{O}$ and $S_{C}$ be defined as above. Ideal Information $\mathcal{I}$ is an enabling mapping satisfying two additional strict conditions:
Completeness (Surjectivity on Domain):
Every semantic state $s_{o}\in S_{O}$ has a physical representation. Formally, $\forall s_{o}\in S_{O},\mathcal{I}(s_{o})\neq\emptyset$. (Inherited from E1).
Reliability (Unambiguous Decoding):
While a semantic state may map to multiple physical carriers (synonyms), any specific physical carrier maps back to at most one semantic state. Formally, if $s_{c}\in\mathcal{I}(s_{o_{1}})$ and $s_{c}\in\mathcal{I}(s_{o_{2}})$, then $s_{o_{1}}=s_{o_{2}}$.
Remark 2.
The set $\mathcal{I}(s_{o})\subseteq S_{C}$ constitutes the synonymous carrier set of the semantic state $s_{o}\in S_{O}$. Therefore, the semantic characteristics of an Ideal Information instance $\mathcal{I}=\langle O,T_{o},S_{O},C,T_{r},S_{C}\rangle$ can be fully characterized by directly studying the semantic state space $S_{O}$, with each $s_{o}\in S_{O}$ being equipped with its associated synonymous carrier set $\mathcal{I}(s_{o})$.
II-C Information Cost of Energy and Time
According to Definitions 23, state transitions of an information carrier are constrained by physical conditions. Building on fundamental results such as the energy requirements of high-frequency harmonic oscillators in classical physics [Nyquist1928Thermal] and the Margolus–Levitin bound in quantum mechanics [Margolus1998maximum], we adopt the following abstract physical principle.
Assumption 1 (Physical Cost Principle).
For a physical system evolving from a state $s_{1}$ to a distinguishable state $s_{2}$, the required energy $E$ and time $\tau$ satisfy
$E\cdot\tau\;\geq\;\kappa,$ (3)
where $\kappa$ is a universal constant with the dimension of action.
Intuitively, Assumption 1 states that every physically distinguishable transition of the carrier system incurs a non-zero resource cost that cannot be made arbitrarily small: the product of the available energy scale and the time scale must exceed a fixed action threshold $\kappa$. In particular, a physical process that implements an information mapping cannot realize arbitrarily many mutually distinguishable semantic alternatives without paying at least a proportional cost in the joint resource $E\cdot\tau$.
In the context of Ideal Information (Definition 4), the semantic state space $S_{O}(O,T_{o})$ has finite cardinality, and the corresponding amount of information is measured in bits as
$I\;=\;\log_{2}|S_{O}(O,T_{o})|.$ (4)
Because Ideal Information is complete and reliably decodable, any physical implementation must, in effect, be able to distinguish among all $|S_{O}(O,T_{o})|=2^{I}$ possible semantic inputs via mutually distinguishable carrier states. This leads to the following resource lower bound.
Theorem 1 (Resource Bound for Ideal Enabling Mapping).
Let $\mathcal{I}:S_{O}(O,T_{o})\rightrightarrows S_{C}(C,T_{r})$ be an Ideal Information instance, and let $I=\log_{2}|S_{O}(O,T_{o})|$ denote the amount of information transmitted (in bits) as in (4). Then any physical implementation of $\mathcal{I}$ satisfies
$E\cdot\tau\;\geq\;\kappa\cdot I,$ (5)
where $E$ is the energy resource, $\tau$ is the implementation time, and $\kappa$ is the universal constant from Assumption 1.
Since $\mathcal{I}$ is Ideal Information in the sense of Definition 4, the mapping $\mathcal{I}:S_{O}(O,T_{o})\rightrightarrows S_{C}(C,T_{r})$ is complete and reliable: every semantic state $s_{o}\in S_{O}(O,T_{o})$ has at least one physical realization $s_{c}\in S_{C}(C,T_{r})$, and distinct semantic states are not confounded at the carrier level. Hence any physical implementation of $\mathcal{I}$ must realize at least $|S_{O}(O,T_{o})|$ pairwise distinguishable carrier states, one (or a synonymous set) for each semantic state.
Equivalently, the implementation process must be capable of distinguishing among $|S_{O}(O,T_{o})|=2^{I}$ alternative semantic inputs. This corresponds to an information content of $I=\log_{2}|S_{O}(O,T_{o})|$ bits.
Consider now the physical evolution of the carrier system during the implementation of $\mathcal{I}$. Each time the system evolves from one distinguishable physical state to another in such a way that it increases or commits a new semantic distinction (for instance, by encoding one more bit of choice among alternative semantic states), Assumption 1 applies: the associated energy and time resources $E_{i}$ and $\tau_{i}$ must satisfy
$E_{i}\cdot\tau_{i}\;\geq\;\kappa.$
Let $N$ denote the number of such information-relevant, distinguishable state transitions required by the implementation. Since each such transition can encode at most a constant number of bits of new distinguishable semantic information, we have $N\geq I$ up to a fixed multiplicative constant. Absorbing this constant into $\kappa$ if necessary, we may assume without loss of generality that $N\geq I$.
We analyze two implementation paradigms:
Case 1: Sequential execution. The total time is $\tau=\sum_{i=1}^{N}\tau_{i}$, and energy can be reused across steps, so the effective energy is $E=\max_{i}E_{i}$. By Assumption 1, $E\cdot\tau_{i}\geq\kappa$ for each $i$, hence
$E\cdot\tau=E\cdot\sum_{i=1}^{N}\tau_{i}\geq N\kappa\geq\kappa I.$
Case 2: Parallel execution. The total time is $\tau=\max_{i}\tau_{i}$, and energy must be summed, so $E=\sum_{i=1}^{N}E_{i}$. By Assumption 1, $E_{i}\cdot\tau\geq\kappa$ for each $i$, hence
$E\cdot\tau=\tau\cdot\sum_{i=1}^{N}E_{i}\geq N\kappa\geq\kappa I.$
In both cases, $E\cdot\tau\geq\kappa I$. ∎
In the following, we interpret the constant $\kappa$ in Assumption 1 in the usual thermodynamic setting of information processing. Specifically, for systems operating in a thermal equilibrium environment at temperature $T$ and undergoing quasistatic transformations, the minimal resource cost per bit can be identified with the Landauer bound $k_{B}T\ln 2$ measured in terms of dissipated heat. Under this interpretation, Theorem 1 specializes to a Landauer-type inequality for Ideal Information.
Corollary 1 (Landauer Principle for Ideal Information).
In a thermal equilibrium environment at temperature $T$, an Ideal Information instance implemented via a quasistatic process satisfies the following lower bound on energy dissipation:
$E_{\mathrm{diss}}\;\geq\;k_{B}T\ln 2\cdot I,$ (6)
where $I=\log_{2}|S_{O}(O,T_{O})|$ is the amount of information transmitted losslessly (in bits).
In a thermal environment at temperature $T$, Landauer’s principle[landauer1961irreversibility, berut2012experimental] states that any logically irreversible operation that erases or thermodynamically commits one bit of information leads, in the quasistatic limit, to a minimal heat dissipation of at least $k_{B}T\ln 2$. In the setting of Theorem 1, implementing an ideal information instance that reliably encodes $I=\log_{2}|S_{O}(O,T_{o})|$ bits requires at least $I$ effective information-bearing operations, each of which is thermodynamically equivalent to committing one bit of semantic information to a physical carrier state.
Identifying the per-bit resource constant in Assumption 1 with $\kappa_{T}=k_{B}T\ln 2$ in this thermal, quasistatic regime, Theorem 1 implies that the total dissipated energy satisfies
$E_{\mathrm{diss}}\;\geq\;\kappa_{T}\cdot I\;=\;k_{B}T\ln 2\cdot I,$
which is exactly the bound in (6). This yields the claimed Landauer-type lower bound for any lossless implementation of Ideal Information. ∎
II-D Atomic Information
In an Ideal Information instance $\mathcal{I}=\langle O,T_{o},S_{O},C,T_{r},S_{C}\rangle$ (Definition 4), the ontological state space $S_{O}$ is equipped with a predecessor function
$P_{O}:S_{O}\to 2^{S_{O}},$ (7)
which captures the semantic causal structure among states. Intuitively, $P_{O}(s)$ is the set of semantic states that can directly lead to $s$ within the ontological dynamics of $O$. We impose the following axioms on $P_{O}(\cdot)$.
Axiom 1 (Strict Temporal Precedence).
For any $s_{o}(o,t),s^{\prime}_{o^{\prime}}(o^{\prime},t^{\prime})\in S_{O}$, if $s^{\prime}_{o^{\prime}}\in P_{O}(s_{o})$, then
$t^{\prime}<t.$
Axiom 2 (Finite and Computable Predecessors).
For any $s\in S_{O}$, the predecessor set $P_{O}(s)$ is finite, and there exists an effective procedure that, given $s$, computes $P_{O}(s)$.
Axiom 3 (Lower-Bounded and Discrete Time).
The time domain $T_{o}$ is bounded from below and discrete, i.e., there exists $t_{\min}\in T_{o}$ and $\Delta t>0$ such that for any strictly decreasing sequence $\{t_{n}\}$ in $T_{o}$ we have $t_{n}\geq t_{\min}$ and $t_{n}-t_{n+1}\geq\Delta t$.
Axiom 1 states that predecessors must occur strictly earlier in ontological time; Axiom 2 excludes infinitely branching causal histories and ensures that local causal structure is effectively inspectable; and Axiom 3 guarantees that there is no infinite regress of ever earlier states in finite time. Together, these axioms imply that the backward causal unfolding of any semantic state is a finite, effectively traversable structure. This leads to the following result.
Theorem 2 (Existence and Computability of Semantic Atomic States).
For any target semantic state $s_{\text{target}}\in S_{O}$, there exists a finite, non-empty set of semantic atomic states $A(s_{\text{target}})\subseteq S_{O}$ such that:
  1. (i)
    Each $s_{\text{atom}}\in A(s_{\text{target}})$ is predecessor-free:
    $P_{O}(s_{\text{atom}})=\varnothing.$
  2. (ii)
    Every $s_{\text{atom}}\in A(s_{\text{target}})$ can reach $s_{\text{target}}$ along the causal structure induced by $P_{O}(\cdot)$, i.e., there exists a finite sequence
    $s_{\text{atom}}=s_{0},s_{1},\dots,s_{n}=s_{\text{target}}$
    such that $s_{i}\in P_{O}(s_{i+1})$ for all $i=0,\dots,n-1$.
  3. (iii)
    The mapping $s_{\text{target}}\mapsto A(s_{\text{target}})$ is computable.
Fix any $s_{\text{target}}\in S_{O}$. Consider the backward expansion of $s_{\text{target}}$ along the predecessor function $P_{O}(\cdot)$: start from $s_{\text{target}}$, enumerate $P_{O}(s_{\text{target}})$, then for each $s^{(1)}\in P_{O}(s_{\text{target}})$ enumerate $P_{O}(s^{(1)})$, and so on.
By Axiom 1, every time we move from a state to one of its predecessors, the time-stamp strictly decreases: if $s^{\prime}_{o^{\prime}}(o^{\prime},t^{\prime})\in P_{O}(s_{o}(o,t))$, then $t^{\prime}<t$. Since $T_{o}$ is lower-bounded and discrete (Axiom 3), there is no infinite strictly decreasing temporal sequence. Therefore, any backward traversal starting from $s_{\text{target}}$ must terminate in finitely many steps. Equivalently, the backward expansion tree is of finite depth.
By Axiom 2, each node has a finite predecessor set, so the branching factor of the backward expansion tree is finite. Thus the entire backward tree rooted at $s_{\text{target}}$ consists of finitely many nodes. The leaves of this finite tree are exactly those states with no predecessors, i.e., all $s\in S_{O}$ such that $P_{O}(s)=\varnothing$ and that can reach $s_{\text{target}}$ along the tree.
Define $A(s_{\text{target}})$ to be the set of all such leaves. Since the tree is finite and non-empty (it contains at least $s_{\text{target}}$), $A(s_{\text{target}})$ is finite and non-empty, and every element has no predecessors, establishing (i) and (ii).
For (iii), consider the following recursive procedure
TRACEATOMS(s)(s):
  • If $P_{O}(s)=\varnothing$, output $\{s\}$.
[…]
Theorem 2 shows that, for any semantic state of interest, there exists a finite and computable collection of causally irreducible semantic states that underpin it. We now upgrade this semantic view to the level of physically realizable information.
Definition 5 (Atomic Information).
Let $\mathcal{I}=\langle O,T_{o},S_{O},C,T_{r},S_{C}\rangle$ be an Ideal Information instance and let $\alpha\in S_{O}$ be any semantic state. The Atomic Information associated with $\alpha$, denoted by $\mathcal{I}_{\alpha}$, is defined as the restriction of $\mathcal{I}$ to the singleton $\{\alpha\}$:
$\mathcal{I}_{\alpha}\;=\;\langle O,\,T_{o},\,\{\alpha\},\,C,\,T_{r},\,\mathcal{I}(\alpha)\rangle,$ (8)
where
$\mathcal{I}(\alpha)=\left\{\,s_{c}\in S_{C}\mid\begin{array}[]{@{}l@{}}\text{the realization of }s_{c}\text{ yields}\\ \text{the semantic state }\alpha\end{array}\right\}$
is the set of carrier states synonymous with $\alpha$ under the enabling mapping $\mathcal{I}:S_{O}\rightrightarrows S_{C}$.
Because $\mathcal{I}$ is Ideal Information (Definition 4), it satisfies completeness and unambiguous decoding. In particular, for every semantic state $\alpha\in S_{O}$ we have $\mathcal{I}(\alpha)\neq\varnothing$, so there always exists at least one carrier state $s_{c}\in S_{C}$ that physically realizes $\alpha$. Thus each $\mathcal{I}_{\alpha}$ defined in (8) is a physically realizable information entity that tightly couples an abstract semantic state with the set of its synonymous physical representations.
Combining Theorem 2 with the above definition yields the following corollary.
Corollary 2 (Existence and Computability of Atomic Information).
Let $\mathcal{I}=\langle O,T_{o},S_{O},C,T_{r},S_{C}\rangle$ be an Ideal Information instance, and let $s_{\text{target}}\in S_{O}$ be any semantic state. Then:
  1. 1.
    There exists a finite, non-empty set of semantic atomic states $A(s_{\text{target}})\subseteq S_{O}$ such that each $\alpha\in A(s_{\text{target}})$ is predecessor-free and causally contributes to $s_{\text{target}}$ as in Theorem 2.
  2. 2.
    For every $\alpha\in A(s_{\text{target}})$, the corresponding Atomic Information
    $\mathcal{I}_{\alpha}\;=\;\langle O,T_{o},\{\alpha\},C,T_{r},\mathcal{I}(\alpha)\rangle$
    is well-defined and physically realizable, and the finite set $\{\mathcal{I}_{\alpha}\mid\alpha\in A(s_{\text{target}})\}$ is computable from $s_{\text{target}}$.
Item (1) is exactly Theorem 2.
For item (2), since $\mathcal{I}$ is Ideal Information, its completeness property ensures $\mathcal{I}(\alpha)\neq\varnothing$ for every $\alpha\in S_{O}$; in particular, for each $\alpha\in A(s_{\text{target}})$, $\mathcal{I}(\alpha)$ is a non-empty set of synonymous carrier states, so $\mathcal{I}_{\alpha}$ in (8) is well-defined and physically realizable.
Moreover, Theorem 2 provides a total, effective procedure TRACEATOMS that, given $s_{\text{target}}$, computes $A(s_{\text{target}})$. Under the system model adopted in this paper, the enabling mapping $\mathcal{I}(\cdot)$ is also assumed to be computable for each input semantic state. Therefore, for every $\alpha\in A(s_{\text{target}})$ we can effectively enumerate $\mathcal{I}(\alpha)$ and construct the corresponding Atomic Information $\mathcal{I}_{\alpha}$. Since $A(s_{\text{target}})$ is finite, the set $\{\mathcal{I}_{\alpha}\mid\alpha\in A(s_{\text{target}})\}$ is finite and computable, as claimed. ∎
II-E Semantic Inference Structure
The predecessor function $P_{O}(\cdot)$ introduced above captures the semantic causal structure among ontological states. In order to connect this causal view with the proof-theoretic notion of derivation in the underlying logical system $\mathcal{L}$, we make the following alignment assumption.
Assumption 2 (Alignment of Causality and Inference).
For every semantic state $s\in S_{O}$, the predecessor set $P_{O}(s)$ coincides with the set of its immediate inference premises in $\mathcal{L}$. More precisely:
  • Soundness of predecessors: For every $s^{\prime}\in P_{O}(s)$, there exists a single inference step in $\mathcal{L}$ that derives $s$ from $s^{\prime}$ (possibly together with elements of $P_{O}(s)\setminus\{s^{\prime}\}$ as side premises).
  • Completeness of predecessors: Conversely, whenever there is a single-step inference in $\mathcal{L}$ whose conclusion is $s$, all semantic premises that this inference depends on are contained in $P_{O}(s)$.
In other words, the directed graph induced by $P_{O}(\cdot)$ is (up to inessential bureaucratic details of the proof system) identical to the derivation graph generated by single-step inference rules of $\mathcal{L}$ over $S_{O}$.
On this basis, we define the global set of semantic atomic states as
$\text{Atom}(S_{O})\coloneqq\{\,s\in S_{O}\mid P_{O}(s)=\emptyset\,\}.$ (9)
Intuitively, $\text{Atom}(S_{O})$ collects all ontological states that have no semantic predecessors and therefore serve as the “primitive” or “axiomatic” basis for subsequent derivations in $\mathcal{L}$. By Axioms 1–3, $\text{Atom}(S_{O})$ is well-defined; it may be finite or infinite, but any particular target state $s_{\text{target}}$ depends only on a finite subset of $\text{Atom}(S_{O})$ identified in Theorem 2.
II-F Derivation Depth
With the atomic basis and the semantic inference structure in place, we can now define the complexity of any derived state in terms of its distance from the atomic layer along the causal/inference graph [Buss1987Boolean][Krajicek1995Bounded].
Definition 6 (Derivation Depth).
Under Assumption 2, the predecessor function $P_{O}(\cdot)$ induces the immediate semantic inference structure of $\mathcal{L}$ over $S_{O}$. For any state $s\in S_{O}$, the Derivation Depth of $s$, denoted $\text{Dd}(s)$, is defined inductively as follows:
$\text{Dd}(s)\coloneqq\begin{cases}0,&\text{if }P_{O}(s)=\emptyset,\\[2.15277pt] 1+\displaystyle\min_{s^{\prime}\in P_{O}(s)}\text{Dd}(s^{\prime}),&\text{otherwise.}\end{cases}$ (10)
Equivalently, $\text{Dd}(s)$ is the length of the shortest finite derivation chain
$s_{0},s_{1},\dots,s_{n}=s$ (11)
such that
  1. 1.
    $s_{0}\in\text{Atom}(S_{O})$ (i.e., $P_{O}(s_{0})=\emptyset$);
  2. 2.
    $s_{i}\in P_{O}(s_{i+1})$ for all $i=0,\dots,n-1$.
In this case we write
[…]
In particular, for any atomic state $\alpha\in\text{Atom}(S_{O})$, we have $\text{Dd}(\alpha)=0$.
Theorem 3 (Computability of Derivation Depth).
For any state $s\in S_{O}$, the derivation depth $\text{Dd}(s)$ defined in (10) is a unique, finite, and computable integer.
We view the directed graph over $S_{O}$ induced by the predecessor function $P_{O}(\cdot)$. By Axiom 1 (Strict Temporal Precedence), every edge $s^{\prime}\in P_{O}(s)$ leads from $s$ to a strictly earlier time-stamp $t^{\prime}<t$. By Axiom 3 (Lower-Bounded and Discrete Time), there is no infinite strictly decreasing sequence of time-stamps. Hence, every backward path
$s=s_{n},s_{n-1},\dots,s_{0}$
following predecessors $s_{i-1}\in P_{O}(s_{i})$ has finite length. Equivalently, the backward expansion tree rooted at any $s\in S_{O}$ has finite depth.
By Axiom 2 (Finite and Computable Predecessors), each state $s$ has a finite, effectively computable predecessor set $P_{O}(s)$. Thus, the backward expansion from $s$ yields a finite, effectively traversable directed acyclic graph (DAG) whose edges are given by the predecessor relation.
Now consider the recursive definition (10):
$\text{Dd}(s)=\begin{cases}0,&\text{if }P_{O}(s)=\emptyset,\\[2.15277pt] 1+\displaystyle\min_{s^{\prime}\in P_{O}(s)}\text{Dd}(s^{\prime}),&\text{otherwise.}\end{cases}$
Well-foundedness. Every recursive call from $s$ to $s^{\prime}\in P_{O}(s)$ strictly decreases the associated time-stamp (by Axiom 1). Since time is bounded below and discrete (Axiom 3), there is no infinite descending chain of such calls. Therefore, the recursion is well-founded.
Termination. The recursion depth is bounded by the maximal length of a backward predecessor chain starting from $s$, which is finite as argued above. Hence, the evaluation of $\text{Dd}(s)$ terminates after finitely many steps.
Uniqueness and computability. For any non-atomic state $s$ with $P_{O}(s)\neq\emptyset$, the set $\{\text{Dd}(s^{\prime})\mid s^{\prime}\in P_{O}(s)\}$ consists of finitely many well-defined integers (by the well-founded recursion). The minimum over this finite set is unique. Moreover, because $P_{O}(s)$ is effectively computable for every $s$ (Axiom 2), we can compute $\text{Dd}(s)$ by a bottom-up dynamic-programming traversal of the finite backward DAG: first assign depth 0 to all atomic states ($P_{O}(s)=\emptyset$), then iteratively compute $\text{Dd}(s)$ for states whose predecessors’ depths are already known. Thus, $\text{Dd}(s)$ is a unique, finite, and computable integer for every $s\in S_{O}$. ∎
Proposition 1 (Derivation depth as Bennett-style logical depth).
Let $\mathcal{L}$ be the underlying logical system and let $P$ be a fixed sound and complete proof system for $\mathcal{L}$. Let $\Gamma$ be a finite base theory, and let $\varphi$ be any formula such that $\Gamma\vdash_{P}\varphi$. We write $Dd(\varphi\mid\Gamma)$ for the minimal derivation depth of $\varphi$ with respect to the semantic/inference structure induced by $\Gamma$ in the sense of Definition 6.
Fix a universal Turing machine $U$, and fix a computable encoding scheme $\langle\cdot\rangle$ that encodes formulas, proofs, and finite sets of formulas as finite strings. Assume that the following two conditions hold.
  1. (EPS)
    Efficient proof simulation. There exists a constant $c_{1}>0$ such that, for any proof $\pi$ in $P$ of length [^1] Here the length of $\pi$ is measured as the number of inference steps in $P$. $n$ whose conclusion is $\varphi$, the universal machine $U$, given as input $\langle\mathrm{Atom}(\Gamma)\rangle$ and $\langle\pi\rangle$, can verify $\pi$ and output $\langle\varphi\rangle$ within time at most $c_{1}\cdot n$.
  2. (NPC)
    No super-linear proof compression into time. There exists a constant $c_{2}>0$ such that, for any program $p$ for $U$ with oracle access to $\mathrm{Atom}(\Gamma)$ that outputs $\langle\varphi\rangle$ in time $T$, there is a proof $\pi$ in $P$ of $\varphi$ from $\Gamma$ of length at most $c_{2}\cdot T$.
Then there exist constants $c,C>0$ (depending only on $U$, $P$ and the encoding scheme, but not on $\Gamma$ or $\varphi$) such that for all $\varphi$ with $\Gamma\vdash_{P}\varphi$ we have
$\begin{split}c\cdot\mathrm{Dd}(\varphi\mid\Gamma)\;\leq\;&\mathrm{depth}_{C\cdot\mathrm{Dd}(\varphi\mid\Gamma)}\bigl(\langle\varphi\rangle\mid\langle\mathrm{Atom}(\Gamma)\rangle\bigr)\\ \;\leq\;&C\cdot\mathrm{Dd}(\varphi\mid\Gamma),\end{split}$ (12)
where $\mathrm{depth}_{t}(x\mid y)$ denotes Bennett’s logical depth [bennett1988logical] of string $x$ at significance level $t$, conditional on $y$, with respect to the universal machine $U$.
In particular, under (EPS) and (NPC) we have
$\mathrm{depth}_{O(Dd(\varphi\mid\Gamma))}\bigl(\langle\varphi\rangle\mid\langle\mathrm{Atom}(\Gamma)\rangle\bigr)\;=\;\Theta\bigl(Dd(\varphi\mid\Gamma)\bigr),$ (13)
where the implicit constants in the $O(\cdot)$ and $\Theta(\cdot)$ notation depend only on $U$, $P$, and the chosen encodings.
III Derivation Depth as an Information Metric
We now establish that derivation depth is an information-theoretic quantity dual to Shannon entropy [shannon1948mathematical], bridging the gap between logical reasoning and information processing.
III-A Derivation Entropy and Information Content
Definition 7 (Derivation Entropy – General Form).
The derivation entropy from $\psi$ to $\phi$ is defined as:
$H_{\mathrm{derive}}(\phi|\psi)=\mathrm{Dd}(\phi|\psi)\cdot\lambda,$ (14)
where $\mathrm{Dd}(\phi|\psi)$ is the derivation depth (Definition 6) and $\lambda>0$ is the information cost per inference step.
Remark 3 (Physical Interpretation of $\lambda$).
The parameter $\lambda$ quantifies the information-theoretic cost per inference step:
  1. 1.
    Classical thermodynamics: Each logically irreversible step erases $\geq 1$ bit. By Landauer’s principle, $\lambda=\ln 2$ nats [landauer1961irreversibility].
  2. 2.
    Quantum mechanics: By the Margolus–Levitin bound, $\lambda=\pi/2$ in units of $\hbar$ [Margolus1998maximum].
  3. 3.
    Proof system structure: Selecting from $|\mathcal{R}|$ rules per step suggests $\lambda\gtrsim\ln|\mathcal{R}|$ nats, as a structural lower bound on the information cost of inference.
TABLE I: Analogy Between Thermodynamic and Information-Theoretic Costs
Domain Discrete Unit Continuous Unit Conversion Factor
Thermodynamics 1 bit erasure Energy (J) $k_{B}T\ln 2$
Information Physics 1 inference step Information (nats) $\ln 2$
Assumption 3 (Classical Thermodynamic Regime).
The knowledge system operates in the classical thermodynamic regime: (1) Inference steps are logically irreversible at thermal equilibrium; (2) Each step erases $O(1)$ bits; (3) Landauer erasure dominates other costs. Under this assumption, $\lambda=\ln 2$ nats.
Throughout the remainder of this paper, unless otherwise specified, we adopt the classical thermodynamic regime with $\lambda=\ln 2$ nats, so that derivation entropy is measured in the same units (nats) as Shannon entropy.
Corollary 3 (Derivation Entropy in Classical Regime).
Under Assumption 3: $H_{\mathrm{derive}}(\phi|\psi)=\mathrm{Dd}(\phi|\psi)\cdot\ln 2$.
Lemma 1 (Derivation Encoding).
Let $\mathcal{P}$ be a proof system with rule set $R$ and maximum rule arity $k$. Any derivation $\pi$ of length $n$ from $m$ premises satisfies:
$|\mathrm{Enc}(\pi)|\leq n\cdot c_{L}\cdot\log(m+n)+O(\log(m+n)),$ (15)
where $c_{L}=\max\{\lceil\log_{2}|R|\rceil,k\}$.
A derivation $\pi=((r_{1},\vec{p}_{1}),\ldots,(r_{n},\vec{p}_{n}))$ consists of $n$ steps. Each step $(r_{i},\vec{p}_{i})$ encodes:
  • Rule: $r_{i}\in R$ in $\lceil\log_{2}|R|\rceil$ bits.
  • Premises: $\vec{p}_{i}=(p_{i,1},\ldots,p_{i,k_{i}})$ with $k_{i}\leq k$ pointers, each in $\{1,\ldots,m+i-1\}$, requiring $\leq k\cdot\lceil\log_{2}(m+n)\rceil$ bits total.
Per-step encoding: $\lceil\log_{2}|R|\rceil+k\cdot\lceil\log_{2}(m+n)\rceil\leq c_{L}\cdot(1+\log_{2}(m+n))$.
Over $n$ steps: $n\cdot c_{L}\cdot(1+\log_{2}(m+n))$.
Adding $O(\log(m+n))$ header bits for $m,n$, and noting that $n\cdot c_{L}\leq n\cdot c_{L}\cdot\log(m+n)$ when $\log(m+n)\geq 1$:
$|\mathrm{Enc}(\pi)|\leq n\cdot c_{L}\cdot\log(m+n)+O(\log(m+n)).\qed$
Definition 8 (Information-Rich Regime).
Let $S_{O}$ be a knowledge base and let $q$ be a query (logical formula) over $S_{O}$. We define the conditional algorithmic entropy of $q$ given $S_{O}$ as
$H(q\mid S_{O})\;:=\;K\bigl(q\mid\langle\text{Atom}(S_{O})\rangle\bigr),$ (16)
where $K(\cdot\mid\cdot)$ denotes conditional Kolmogorov complexity with respect to a fixed universal Turing machine and a computable encoding scheme $\langle\cdot\rangle$ [li2008introduction].
We say that $q$ lies in the information-rich regime (with respect to $S_{O}$) if
$H(q\mid S_{O})\;\geq\;\omega\bigl(\log|\text{Atom}(S_{O})|\bigr),$ (17)
that is, equivalently,
$\frac{H(q\mid S_{O})}{\log|\text{Atom}(S_{O})|}\;\longrightarrow\;\infty\qquad\text{as }|\text{Atom}(S_{O})|\to\infty.$ (18)
For query classes satisfying the polynomial-depth condition
$Dd(q\mid S_{O})\;\leq\;\mathrm{poly}\bigl(|\text{Atom}(S_{O})|\bigr),$ (19)
(e.g., Datalog [ceri1989you], SPARQL [perez2009semantics]), and template-based natural-language queries), conditions (17)–(18) are asymptotically equivalent to
$H(q\mid S_{O})\;\gg\;\log\max\bigl\{|\text{Atom}(S_{O})|,\,Dd(q\mid S_{O})\bigr\},$ (20)
as $|\text{Atom}(S_{O})|\to\infty$.
Lemma 2 (Derivation Incompressibility in the Information-Rich Regime).
Let $S_{O}$ be a knowledge base and $q$ a logical formula such that:
  1. (i)
    $n=Dd(q\mid S_{O})\geq 1$,
  2. (ii)
    $m=|\mathrm{Atom}(S_{O})|$,
  3. (iii)
    $q$ lies in the information-rich regime (Definition 8),
  4. (iv)
    $n=\mathrm{poly}(m)$ (e.g., for bounded-depth reasoning systems).
Then there exists a constant $c_{2}>0$, depending only on the proof system and encoding scheme, such that the conditional Kolmogorov complexity of $q$ given $S_{O}$ satisfies
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)\geq c_{2}\cdot n\log(m+n)-O(\log m).$
By Lemma 1, any derivation of $q$ from $\mathrm{Atom}(S_{O})$ of length $n$ can be encoded in at most
$L_{\text{enc}}=c_{1}\cdot n\log(m+n)+O(\log m)$
bits for some constant $c_{1}>0$. This establishes an upper bound on $K(q\mid\langle\mathrm{Atom}(S_{O})\rangle)$.
Now suppose, for contradiction, that no such lower bound holds. Then for every $\varepsilon>0$, there exist arbitrarily large $m$ and corresponding $q$ such that
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)<\varepsilon\cdot n\log(m+n).$
Since $n=\mathrm{poly}(m)$, we have $\log(m+n)=\Theta(\log m)$, so the above implies
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)=o\bigl(n\log m\bigr).$
Consequently, the conditional algorithmic entropy satisfies
$H(q\mid S_{O})=K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)\cdot\ln 2=o\bigl(n\log m\bigr).$
But because $n\geq 1$ and $n=\mathrm{poly}(m)$, we have $n\log m\geq\log m$, and in fact $n\log m\gg\log m$ as $m\to\infty$. Hence
$H(q\mid S_{O})=o\bigl(n\log m\bigr)\not\gg\log m,$
which contradicts condition (iii)—that $q$ lies in the information-rich regime. Therefore, there must exist a constant $c_{2}>0$ such that
$K\bigl(q\mid\langle\mathrm{Atom}(S_{O})\rangle\bigr)\geq c_{2}\cdot n\log(m+n)-O(\log m),$
as claimed. ∎
Throughout this subsection, all asymptotic notation ($O(\cdot)$, $\Theta(\cdot)$, $\omega(\cdot)$, $\gg$, etc.) is taken with respect to
$m\;:=\;|\text{Atom}(S_{O})|\;\longrightarrow\;\infty,$
under the standing assumption that
$n\;:=\;Dd(q\mid S_{O})\;=\;\mathrm{poly}(m).$
Theorem 4 (Derivation Depth as an Information Metric).
Let $S_{O}$ be a knowledge base and let $q$ be a query satisfying conditions (i)(iv) of Lemma 2 (in particular, $q$ lies in the information-rich regime of Definition 8 and $n:=Dd(q\mid S_{O})=\mathrm{poly}(m)$ with $m:=|\text{Atom}(S_{O})|$). Define
$H(q\mid S_{O})\;:=\;K\bigl(q\mid\langle\text{Atom}(S_{O})\rangle\bigr)\cdot\ln 2,$ (21)
and, under the classical thermodynamic regime (Assumption 3), define the derivation entropy as
$H_{\mathrm{derive}}(q\mid S_{O})\;:=\;Dd(q\mid S_{O})\cdot\ln 2.$ (22)
Then
$H_{\mathrm{derive}}(q\!\mid\!S_{O})\!=\!\Theta\!\left(\frac{H(q\mid S_{O})}{\!\log\!\big(|\text{Atom}(S_{O})|\!+\!\mathrm{Dd}(q\mid S_{O})\big)}\right)\!,$ (23)
where the implicit constants in the $\Theta(\cdot)$ notation depend only on the choice of universal Turing machine, proof system, and encoding scheme, but not on $S_{O}$ or $q$.
Let $m=|\text{Atom}(S_{O})|$ and $n=Dd(q\mid S_{O})$. By Lemma 1 and 2, there exist constants $c_{1}$, $c_{2}>0$, or more precisely, $c_{1}$ and $c_{2}$ depend only on the universal Turing machine, proof system $\mathcal{P}$, and encoding scheme, but not on the polynomial degree in the assumption $n={\rm poly}(m)$, such that
$c_{2}\,n\log(m+n)-O(\log m)\;\leq\;K\bigl(q\mid\langle\text{Atom}(S_{O})\rangle\bigr)\\ \;\leq\;c_{1}\,n\log(m+n)+O(\log m).$
Multiplying both sides by $\ln 2$ yields
$c_{2}^{\prime}\,n\log(m+n)-O(\log m)\;\leq\;H(q\mid S_{O})\\ \;\leq\;c_{1}^{\prime}\,n\log(m+n)+O(\log m),$
for suitable constants $c_{1}^{\prime},c_{2}^{\prime}>0$ with $c_{i}^{\prime}=c_{i}\ln 2$. Because $n=\mathrm{poly}(m)$, we have $\log(m+n)=\Theta(\log m)$, and the $O(\log m)$ term is asymptotically dominated by $n\log(m+n)$. Hence
$H_{\mathrm{derive}}(q\!\mid\!S_{O})\!=\!\Theta\!\left(\!\frac{H(q\mid S_{O})}{\log\!\big(|\text{Atom}(S_{O})|\!+\!\mathrm{Dd}(q\mid S_{O})\big)}\!\right)\!,$
Rearranging gives
$n\;=\;\Theta\!\left(\frac{H(q\mid S_{O})}{\log(m+n)}\right).$
Finally, substituting $H_{\mathrm{derive}}(q\mid S_{O})=n\cdot\ln 2$ and recalling that $m=|\text{Atom}(S_{O})|$ and $n=Dd(q\mid S_{O})$, we obtain
$H_{\mathrm{derive}}(q\!\mid\!S_{O})\!=\!\Theta\!\left(\frac{H(q\mid S_{O})}{\!\log\!\big(|\text{Atom}(S_{O})|\!+\!\mathrm{Dd}(q\mid S_{O})\big)}\right)\!,$
which is exactly (23). This completes the proof. ∎
III-B Frequency-weighted Storage–Computation Tradeoff for a Single Query
In this subsection we combine the semantic information metrics to obtain a quantitative tradeoff between storage and computation under a given query distribution.
Definition 9 (Access Frequency).
For a query distribution $P_{Q}$ over query set $Q$ and a total of $N$ queries, the access frequency (expected access count) of query $q$ is:
$f_{q}:=N\cdot P_{Q}(q).$ (24)
Thus $f_{q}\in[0,N]$ with $\sum_{q\in Q}f_{q}=N$.
Throughout this subsection we assume Definition 9 has been fixed, and we keep its notation. In particular, we let $P_{Q}$ denote the query distribution over a set $Q$ of admissible queries, and
$f_{q}:=N\cdot P_{Q}(q)$
denote the expected access count of query $q$ over $N$ total queries. [^2] As in Definition 9, $f_{q}$ is a dimensionless quantity: it is a pure count (expected number of accesses) without an intrinsic time unit. When we speak of amortized cost per access, we will divide a total bit-length (storage) or total semantic derivation cost (in nats) by this dimensionless $f_{q}$ to obtain a per-access cost in the corresponding information unit.
We now state the central theorem of this section.
Theorem 5 (Frequency-weighted storage–computation tradeoff).
Let $S_{O}$ be a knowledge base with atomic semantic state set $\text{Atom}(S_{O})$, and let $q$ be a query in the information-rich regime with respect to $S_{O}$ in the sense of Definition 8. Let $P_{Q}$ be a query distribution over a query set $Q$, and define $f_{q}:=N\cdot P_{Q}(q)$ as in Definition 9.
Consider a hybrid strategy $\Pi=(S,C,R)$ where:
  1. 1.
    $S$ is an auxiliary storage structure (e.g., materialized facts, indices, intermediate derivations) that is effectively encodable as a binary string $\text{enc}{S}$ of length $\text{abs}{S}$ bits.
  2. 2.
    $C$ and $R$ describe the residual computation (inference) carried out over the combined knowledge base $S_{O}\cup S$.
We impose the following assumptions on $S$ and on inference over $S_{O}\cup S$:
  1. 1.
    (SA1) Information-theoretic measurement of storage. The bit length $|S|$ of the encoding $\text{enc}{S}$ satisfies
    $H(S)\;\leq\;|S|\;\leq\;H(S)+O(1),$ (25)
    where $H(S)$ is the Shannon entropy of $S$ for a suitable prior over admissible storage states. In particular,
    $I(q;S\mid S_{O})\;\leq\;H(S)\;\leq\;|S|.$ (26)
  2. 2.
    (SA2) Structural regularity of the extended knowledge base. The combined knowledge base $S_{O}\cup S$ admits a well-defined atomic semantic state set $\text{Atom}(S_{O}\cup S)$, a predecessor function satisfying Axioms 1-3, and hence a derivation depth function $Dd(\cdot\mid S_{O}\cup S)$ as in Definition 6. We write
    $m_{S}:=|\text{Atom}(S_{O}\cup S)|,n_{S}:=Dd(q\mid S_{O}\cup S).$ (27)
    We assume $n_{S}=\mathrm{poly}(m_{S})$ (polynomial-depth regime) and that $q$ remains in the information-rich regime with respect to $S_{O}\cup S$ (in the sense of Definition 8, replacing $S_{O}$ by $S_{O}\cup S$). Moreover, the size/depth parameters of $S_{O}\cup S$ are of the same polynomial order as those of $S_{O}$.
  3. 3.
    (SA3) Applicability of the derivation-depth bound. The proof-system/encoding conditions (EPS) and (NPC) of Proposition 1 hold not only for the base theory induced by $S_{O}$, but also for the theory induced by $S_{O}\cup S$. Consequently, the asymptotic relation of Theorem 4 applies to inference from $S_{O}\cup S$: there exist constants $c_{1},c_{2}>0$ such that
    $c_{1}\,\frac{H(q\mid S_{O}\cup S)}{\log(m_{S}+n_{S})}-O(1)\!\;\leq\;\!H_{\mathrm{derive}}(q\mid S_{O}\cup S)\\ \;\leq\;c_{2}\,\frac{H(q\mid S_{O}\cup S)}{\log(m_{S}+n_{S})}+O(1),$ (28)
    where
    $H_{\mathrm{derive}}(q\mid S_{O}\cup S):=Dd(q\mid S_{O}\cup S)\,\lambda$
    is the derivation entropy (Definition 7) with per-step cost $\lambda>0$, and all entropies are measured in nats.
Under these assumptions, define:
  • The semantic information content of $q$ (Definition 8):
    $H_{q}\!:=\!H(q\mid S_{O})\!:=\!K\big(\text{enc}{q}\mid\text{enc}{\text{Atom}(S_{O})}\big)\,\ln 2.$ (29)
  • The effective derivation cost factor associated with $S$:
    $L_{S}:=\log(m_{S}+n_{S}),$ (30)
    which captures the logarithmic compression advantage available through derivations over $S_{O}\cup S$.
  • The amortized cost per access of answering $q$ using strategy $\Pi$:
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi;q):=\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S).$ (31)
Then there exist positive constants $c,C>0$ (depending only on the universal Turing machine, proof system, and encoding conventions, but not on $S_{O}$, $q$, $S$, or $P_{Q}$) such that
$\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;\min\left\{\frac{H_{q}}{f_{q}},\;C\,\frac{H_{q}}{L_{S}}\right\}+O(1).$ (32)
In particular, up to additive $O(1)$ and multiplicative constant factors, there exists a strategy $\Pi$:
  1. (1)
    Low-frequency regime ($f_{q}\leq L_{S}$, on-demand computation dominated):
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;C\,\frac{H_{q}}{L_{S}}+O(1).$ (33)
  2. (2)
    High-frequency regime ($f_{q}>L_{S}$, precomputation/storage dominated):
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;\frac{H_{q}}{f_{q}}+O(1).$ (34)
The break-even frequency scale is given (up to constant factors) by
$f_{c}\asymp L_{S}=\log(m_{S}+n_{S}),$ (35)
where, for the baseline case $S=\emptyset$, we have $f_{c}=\log(|\mathrm{Atom}(S_{O})|+D_{d}(q|S_{O}))$; see Remark 4 for the general case.
We sketch the main steps.
Step 0: Information-theoretic preliminaries By (SA1) and the data-processing inequality,
$I(q;S\mid S_{O})\;\leq\;H(S)\;\leq\;|S|.$ (36)
Using the mutual-information identity
$H(q\mid S_{O})=H(q\mid S_{O},S)+I(q;S\mid S_{O}),$ (37)
we obtain
$H(q\mid S_{O},S)=H_{q}-I(q;S\mid S_{O})\geq H_{q}-|S|.$ (38)
Thus $S$ can reduce the conditional entropy of $q$ (given $S_{O}$) by at most $|S|$ nats.
Step 1: Applying Theorem 4. By (SA2) and (SA3), Theorem 4 applies to $S_{O}\cup S$ and yields
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)\;\leq\;c_{2}\,\frac{H(q\mid S_{O}\cup S)}{L_{S}}+O(1).$ (39)
Using the bounds
$H_{q}-|S|\;\leq\;H(q\mid S_{O}\cup S)\;\leq\;H_{q},$ (40)
we obtain
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)\;\leq\;c_{2}\,\frac{H_{q}}{L_{S}}+O\!\left(\frac{|S|}{L_{S}}\right)+O(1).$ (41)
Step 2: Amortized cost. By definition,
$\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)=\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S).$ (42)
Substituting the previous inequality,
$\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;\leq\;\frac{|S|}{f_{q}}+c_{2}\,\frac{H_{q}}{L_{S}}+O\!\left(\frac{|S|}{L_{S}}\right)+O(1).$ (43)
This decomposition makes explicit the storage-amortization term $|S|/f_{q}$ and the derivation-entropy term $\sim H_{q}/L_{S}$.
Step 3 Extremal strategies. To obtain a clean upper bound in terms of $(H_{q},f_{q},L_{S})$ alone, we compare against two constructive extremes.
(a) Pure computation strategy $\Pi_{\mathrm{comp}}$: take $S=\varnothing$. Then, by Theorem 4 applied to $S_{O}$,
$\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{comp}};q)=H_{\mathrm{derive}}(q\mid S_{O})\\ \;\asymp\;\frac{H_{q}}{\log\big(|\text{Atom}(S_{O})|+Dd(q\mid S_{O})\big)}\\ \;\leq\;C_{0}\,\frac{H_{q}}{L_{S}}+O(1)$ (44)
for some constant $C_{0}>0$, where the last inequality uses that $L_{S}=\log(m_{S}+n_{S})$ is of the same asymptotic order as $\log(|\text{Atom}(S_{O})|+Dd(q\mid S_{O}))$ under the polynomial-depth assumptions.
(b) Pure storage strategy $\Pi_{\mathrm{store}}$: by the coding theorem and the definition of $H_{q}$, there exists a conditional code for the answers to $q$ (given $S_{O}$) of length at most $H_{q}+O(1)$ nats. We can store such a near-optimal representation in $S$, so that $|S|\leq H_{q}+O(1)$, and design $C$ and $R$ so that answering $q$ from $S_{O}\cup S$ requires only constant derivation cost:
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)=O(1).$ (45)
Hence
$\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{store}};q)=\frac{|S|}{f_{q}}+O(1)\;\leq\;\frac{H_{q}}{f_{q}}+O(1).$ (46)
Since we are interested in achievable costs, the optimal amortized cost
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q):=\inf_{\Pi}\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)$ (47)
is bounded above by the minimum of these two explicit strategies:
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\\ \;\leq\;\min\big\{\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{comp}};q),\,\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{store}};q)\big\}\\ \;\leq\;\min\left\{C_{0}\,\frac{H_{q}}{L_{S}},\;\frac{H_{q}}{f_{q}}\right\}+O(1).$ (48)
Equivalently, there exists $C>0$ such that
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\leq\;\min\left\{\frac{H_{q}}{f_{q}},\;C\,\frac{H_{q}}{L_{S}}\right\}+O(1),$ (49)
establishing the claimed bound.
Step 4: Frequency regimes. From the last inequality we obtain two asymptotic regimes:
  • If $f_{q}\leq L_{S}$, then $\frac{H_{q}}{f_{q}}\geq\frac{H_{q}}{L_{S}}$ and
    $\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\leq\;\frac{H_{q}}{f_{q}}+O(1).$ (50)
  • If $f_{q}>L_{S}$, then $\frac{H_{q}}{f_{q}}<\frac{H_{q}}{L_{S}}$ and
    $\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\leq\;C\,\frac{H_{q}}{L_{S}}+O(1).$ (51)
The transition occurs when $f_{q}$ is on the order of $L_{S}$, i.e., $f_{c}\asymp L_{S}=\log(m_{S}+n_{S})$. ∎
Remark 4 (Operational interpretation of critical frequency).
The critical frequency $f_{c}$ admits the following operational interpretation:
  1. 1.
    Baseline case ($S=\emptyset$): When no auxiliary storage is used, the critical frequency is determined solely by the base knowledge:
    $f_{c}^{(0)}:=\log\bigl(|\mathrm{Atom}(S_{O})|+D_{d}(q|S_{O})\bigr).$
    This is the threshold for deciding whether any storage investment for query $q$ is worthwhile.
  2. 2.
    Post-allocation case: For a given storage strategy $\Pi=(S,C,R)$, the effective critical frequency shifts to
    $f_{c}(S):=L_{S}=\log(m_{S}+n_{S}),$
    where $m_{S}=|\mathrm{Atom}(S_{O}\cup S)|$ and $n_{S}=D_{d}(q|S_{O}\cup S)$.
  3. 3.
    Decision rule: The initial decision of whether to allocate storage for $q$ should compare $f_{q}$ against the baseline $f_{c}^{(0)}$, not against $f_{c}(S)$, to avoid circular reasoning. Once $S$ is chosen, $f_{c}(S)$ governs the marginal benefit of additional storage.
Remark 5 (Physical and engineering interpretation).
Theorem 5 admits a natural physical and engineering interpretation.
(1) Units and dimensions. By construction, $H_{q}$ and $H_{\mathrm{derive}}$ are measured in nats (or bits up to a constant factor), while $|S|$ is measured in bits. The frequency $f_{q}$ is a dimensionless expected count of accesses. Therefore:
  • $|S|/f_{q}$ has the same information dimension as $|S|$ (bits), but represents an average storage cost per access;
  • $H_{\mathrm{derive}}$ is a semantic “work” or “action” measured in nats;
  • the bound in Theorem 5 compares two quantities of the same information-theoretic dimension: per-access storage and per-access semantic computation.
In a physical implementation, both can be converted into expected energy dissipation using Landauer’s principle or related models of irreversible computation.
(2) Two limiting protocols. The proof explicitly constructs two limiting protocols:
  1. 1.
    A pure computation protocol, where no auxiliary storage is used ($S=\varnothing$) and the cost is determined entirely by derivation from $S_{O}$; this yields an amortized cost of order $H_{q}/\log(\cdot)$.
  2. 2.
    A pure storage protocol, where we store a near-optimally compressed representation of the answer to $q$, making the derivation cost $O(1)$ and the amortized cost dominated by $H_{q}/f_{q}$.
Theorem 5 states that, up to constant factors and lower-order terms, no hybrid protocol can asymptotically outperform the better of these two extremes, once $H_{q}$, $f_{q}$, and $L_{S}$ are fixed.
(3) Break-even frequency and caching. The critical scale $f_{c}\asymp L_{S}=\log(m_{S}+n_{S})$ plays the role of a break-even frequency between caching and on-demand computation:
  • If $f_{q}\ll f_{c}$, the term $H_{q}/f_{q}$ dominates; it is more efficient to answer $q$ by recomputation each time than to allocate storage to cache its answers.
  • If $f_{q}\gg f_{c}$, the term $CH_{q}/L_{S}$ dominates; investing storage to reduce $H(q\mid S_{O}\cup S)$ (and hence $H_{\mathrm{derive}}$) becomes asymptotically advantageous.
This mirrors classical engineering heuristics for cache and index design: high-frequency queries justify aggressive precomputation and indexing, while low-frequency queries are better handled by direct inference.
(4) Role of $L_{S}$. The factor $L_{S}=\log(m_{S}+n_{S})$ reflects the same logarithmic compression phenomenon as in Theorem 4: in the information-rich regime, the derivation cost per nat of semantic information scales like $1/\log(\text{size}+\text{depth})$. A larger, deeper, but still well-structured theory $S_{O}\cup S$ allows the system to answer semantically complex queries with fewer effective derivation steps per nat.
(5) Robustness. The qualitative form of the tradeoff—and the scale $f_{c}\asymp L_{S}$ at which caching becomes beneficial—is robust to:
  • changes in the underlying universal Turing machine (which affect $H_{q}$ only by an additive constant);
  • variations in the proof system, as long as (EPS) and (NPC) hold and Theorem 4 applies;
  • different encodings of $S$, as long as (SA1) holds, i.e., the code is within $O(1)$ of the Shannon-optimal length.
Thus the storage–computation tradeoff is controlled mainly by the semantic and structural properties of $S_{O}$ and $S$, rather than by low-level representational details.
Remark 6 (Amortized storage cost and query frequencies).
In Theorem 5 we analyzed the amortized cost of serving a single fixed query $q$ that is expected to be accessed repeatedly. Recall that for a query distribution $P_{Q}$ over a finite query set $Q$ and a total horizon of $N$ queries, we define in Definition 9 the expected number of accesses of $q$ by
$f_{q}\;\coloneqq\;N\cdot P_{Q}(q).$
Here $f_{q}$ is a frequency (an expected count), not a probability. Under this convention, the one-time storage investment $|S|$ is amortized over the $f_{q}$ accesses of $q$, leading to the single-query amortized cost
$\operatorname{Cost}_{\mathrm{amort}}(q)\;=\;\frac{|S|}{f_{q}}\;+\;H_{\mathrm{derive}}(q\mid S_{O}\cup S)\\ \;=\;\frac{|S|}{N\cdot P_{Q}(q)}\;+\;H_{\mathrm{derive}}(q\mid S_{O}\cup S).$ (52)
Here $|S|$ is the one-time storage cost (in bits) needed to maintain $S$, and $H_{\mathrm{derive}}(q\mid S_{O}\cup S)$ is the per-access inference cost (in bits) to derive $q$ from $(S_{O},S)$.
A tempting but incorrect way to “average” this cost across query types is to write
$\mathbb{E}_{q\sim P_{Q}}\Bigl[\frac{|S|}{P_{Q}(q)}+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\Bigr]\\ \;=\;\sum_{q_{i}\in Q}P_{Q}(q_{i})\Bigl(\frac{|S|}{P_{Q}(q_{i})}+H_{\mathrm{derive}}(q_{i}\mid S_{O}\cup S)\Bigr).$ (53)
By direct calculation, this yields
$\sum_{q_{i}}P_{Q}(q_{i})\,\frac{|S|}{P_{Q}(q_{i})}\;=\;|S|\cdot|Q|,$
and hence
$\mathbb{E}_{q\sim P_{Q}}\Bigl[\frac{|S|}{P_{Q}(q)}+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\Bigr]\\ \;=\;|S|\cdot|Q|\;+\;\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr].$
This expression is conceptually wrong: it treats $|S|$ as if it were incurred separately for each query type, effectively charging $|S|$ a total of $|Q|$ times. This contradicts the amortization model underlying Theorem 5 and Definition 9, where $|S|$ is a single one-time investment that is shared across all queries.
To see the problem numerically, consider an example where $Q=\{q_{1},q_{2},q_{3}\}$,
$P_{Q}(q_{1})=0.5,\quad P_{Q}(q_{2})=0.3,\quad P_{Q}(q_{3})=0.2,$
the storage cost is $|S|=100$ bits, and
$H_{\mathrm{derive}}(q_{1}\mid S_{O}\cup S)=10,$
$H_{\mathrm{derive}}(q_{2}\mid S_{O}\cup S)=20,$
$H_{\mathrm{derive}}(q_{3}\mid S_{O}\cup S)=30.$
(bits per access). Let $N=1000$ be the total number of queries. Then the expected total inference cost is
$N\cdot\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]$
= $1000\cdot(0.5\cdot 10+0.3\cdot 20+0.2\cdot 30)$
= $17{,}000\text{ bits},$
and the total cost over all queries is
$|S|+17{,}000\;=\;17{,}100\text{ bits}.$
The correct average cost per query is therefore
$\frac{|S|}{N}\;+\;\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]\\ \;=\;\frac{100}{1000}+17\;=\;17.1\text{ bits per query}.$ (54)
In contrast, the naive averaging in (53) yields
$|S|\cdot|Q|+17\;=\;100\cdot 3+17\;=\;317\text{ bits per query},$
which is clearly nonsensical: the one-time cost $|S|$ has been multiplied by the number of query types $|Q|$ instead of being spread over the actual number of query occurrences $N$.
In summary, the correct way to account for storage in a multi-query setting is:
  • treat $|S|$ as a single, one-time storage investment;
  • amortize it over the total number $N$ of query occurrences, giving a per-query storage contribution of $|S|/N$;
  • add the expected per-query inference cost $\mathbb{E}_{q\sim P_{Q}}[H_{\mathrm{derive}}(q\mid S_{O}\cup S)]$.
This yields the expected cost per query as in (54):
$\mathbb{E}[\operatorname{Cost}]\;=\;\frac{|S|}{N}\;+\;\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr].$ (55)
If we now view $Q$ as a random variable with distribution $P_{Q}$, and interpret $H_{\mathrm{derive}}(q\mid S_{O}\cup S)$ as a per-instance algorithmic cost compatible with a Shannon conditional entropy $H(Q\mid S_{O},S)$, then
$\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]\;\approx\;H(Q\mid S_{O},S).$
Consequently, the expected Shannon cost term can be decomposed as
$H(Q\mid S_{O},S)\;=\;H(Q\mid S_{O})-I(S;Q\mid S_{O}),$ (56)
where $I(S;Q\mid S_{O})$ is the conditional mutual information between $S$ and $Q$ given $S_{O}$. Equation (56) shows that the storage $S$ is beneficial to the extent that it captures information about the future queries $Q$ beyond what is already in $S_{O}$: maximizing $I(S;Q\mid S_{O})$ minimizes the residual cost $H(Q\mid S_{O},S)$, and thus reduces the expected inference burden in (55).
In summary, Theorem 5 provides a frequency-weighted amortized cost bound at the level of a single query type, while Remark 6 clarifies how the one-time storage cost $|S|$ should be correctly amortized. This completes the frequency-weighted analysis at the level of a single query type.
III-C System-Wide Storage Allocation under Amortized Cost
We now move from the single-query amortization view to a system-wide analysis over a full query distribution.
Remark 7 (From single-query amortization to system-wide averages).
Remark 6 clarifies that Theorem 5 naturally applies first to a single fixed query $q$ that is accessed $f_{q}=N\cdot P_{Q}(q)$ times. In that setting, the amortized cost per access of $q$ is
$\operatorname{Cost}_{\mathrm{amort}}(q)\;=\;\frac{|S|}{f_{q}}\;+\;H_{\mathrm{derive}}(q\mid S_{O}\cup S),$ (57)
where $|S|/f_{q}$ is the per-access share of the one-time storage cost $|S|$. This formula is meaningful only when $f_{q}$ is understood as a number of accesses (frequency), in accordance with Definition 9.
When we move from this single-query view to a multi-query workload with distribution $P_{Q}$ and horizon $N$, the correct aggregation is:
$\text{total storage cost}\;=\;|S|,$
$\text{total inference cost}\;=\;\sum_{q\in Q}f_{q}\cdot H_{\mathrm{derive}}(q\mid S_{O}\cup S)$
$\;=\;N\cdot\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr],$
$\text{total number of queries}\;=\;\sum_{q\in Q}f_{q}\;=\;N.$
Hence the system-wide average cost per query is exactly
$\mathbb{E}[\operatorname{Cost}]\;=\;\frac{|S|}{N}\;+\;\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr],$ (58)
which is consistent with (55). In particular, the one-time storage cost $|S|$ appears only once in the numerator and is amortized over the $N$ actual query occurrences, rather than over the number of query types $|Q|$.
If, incorrectly, one were to start from the single-query formula (57), replace $f_{q}$ by $N\cdot P_{Q}(q)$, and then naively average over $q$ as in
$\sum_{q_{i}\in Q}P_{Q}(q_{i})\Bigl(\frac{|S|}{P_{Q}(q_{i})}+H_{\mathrm{derive}}(q_{i}\mid S_{O}\cup S)\Bigr),$ (59)
one would again obtain the spurious term $|S|\cdot|Q|$ discussed in Remark 6. This mistake arises from double-counting the storage cost $|S|$ as if it were separately incurred for each query type. The correct interpretation is that $|S|$ is a fixed capital cost that supports the entire query workload, and therefore must be amortized over the full horizon $N$ of query instances.
From an information-theoretic perspective, if we interpret
$\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]\;\approx\;H(Q\mid S_{O},S),$
with $H(\cdot\mid\cdot)$ now denoting the Shannon conditional entropy of the random variable $Q$, then the second term in (58) can be written as
$H(Q\mid S_{O},S)\;=\;H(Q\mid S_{O})-I(S;Q\mid S_{O}),$
where $I(S;Q\mid S_{O})$ is the conditional mutual information between $S$ and $Q$ given $S_{O}$. This identity shows that, at the system level, the storage $S$ is beneficial to the extent that it reduces the Shannon uncertainty of future queries $Q$ beyond what $S_{O}$ already provides. In other words, $S$ is “useful” precisely when it encodes structure about the query distribution $P_{Q}$ so as to increase $I(S;Q\mid S_{O})$ and thereby decrease the expected inference cost term in (58).
We remark that the per-query algorithmic complexity $H_{\mathrm{derive}}(q\mid S_{O}\cup S)$ used in Theorems 15 and the Shannon entropy $H(Q\mid S_{O},S)$ appearing here live at different levels: the former describes instance-wise description length or inference difficulty, while the latter describes average-case uncertainty over the random variable $Q$. They are, however, compatible in the sense that suitable averaging of algorithmic costs over $q\sim P_{Q}$ recovers the Shannon quantities, which is why both viewpoints lead to the same amortized-cost decomposition.
Practical implication: The critical frequency $f_{c}$ provides a quantitative decision rule for capacity allocation in knowledge systems. For a query type $q$ with expected frequency $f_{q}$ over a horizon of $N$ queries, precomputing and storing its answer (or relevant structure) is worthwhile when the one-time storage cost per query $\frac{|S_{q}|}{f_{q}}$ is smaller than the expected per-query inference cost $H_{\mathrm{derive}}(q\mid S_{O})$ (or, in a Shannon view, its expected uncertainty $H(Q\mid S_{O})$). Equivalently, the storage investment is justified when $f_{q}>f_{c}$, where the threshold $f_{c}$ depends on the available capacity and the inference difficulty of $q$.
Concretely, this yields the following design rules in different settings:
  • LLMs (parametric memory). Let $N$ be the number of parameters and suppose storing a fact $q$ in parameters consumes $|S_{q}|$ bits of effective capacity while avoiding $H_{\mathrm{derive}}(q\mid S_{O})$ bits of computation per access. A simple rule of thumb is to store $q$ parametrically only if
    $f_{q}\cdot H_{\mathrm{derive}}(q\mid S_{O})\;\gtrsim\;|S_{q}|.$
    If we model the “effective” per-parameter capacity at the scale of GPT-4 ($N\approx 1.7\times 10^{12}$[openai2023gpt4] by $\log N\approx 40$ bits, this suggests that only facts whose derivation requires on the order of tens of bits of computation and that are accessed at least moderately often should be encoded in the weights. Rare or computationally easy facts are better left to retrieval or external tools.
  • Knowledge graphs (materialized views). For a join pattern $q$ over a graph with $|E|$ edges, let $|S_{q}|$ denote the storage needed for a materialized view and $H_{\mathrm{derive}}(q\mid S_{O})$ the cost of evaluating $q$ on demand. Using the same amortization logic, it is beneficial to materialize when $f_{q}>f_{c}\approx|S_{q}|/H_{\mathrm{derive}}(q\mid S_{O})$. In many practical graph workloads, both $|S_{q}|$ and $H_{\mathrm{derive}}(q\mid S_{O})$ scale only polylogarithmically in $|E|$, which can be crudely summarized as a threshold of the form
    $f_{c}\;\approx\;1+\frac{1}{\log|E|}.$
    For a large knowledge graph such as Freebase ($|E|\approx 10^{9}$[bollacker2008freebase], this gives $f_{c}\approx 1.05$, meaning that even slightly more-than-once multi-hop patterns can already justify precomputation.
  • Edge devices (caching). Suppose answering a query from the cloud costs $H_{\mathrm{cloud}}$ and answering it from a local cache costs $H_{\mathrm{local}}$, with a device-to-cloud speed ratio $\eta=H_{\mathrm{local}}/H_{\mathrm{cloud}}$. Caching $q$ locally with storage cost $|S_{q}|$ is beneficial when
    $f_{q}\cdot(H_{\mathrm{cloud}}-H_{\mathrm{local}})\;\gtrsim\;|S_{q}|,$
    i.e., when its expected number of accesses exceeds
    $f_{c}\;\approx\;\frac{|S_{q}|}{H_{\mathrm{cloud}}(1-\eta)}.$
    If we normalize units so that $|S_{q}|/H_{\mathrm{cloud}}\approx 1$ for a typical cached answer, this reduces to the simple rule $f_{q}\gtrsim 1/(1-\eta)$. For a 4G-like setting with $\eta\approx 0.1$, this gives $f_{c}\approx 1.1$ in idealized units; in practice, protocol overheads and limited cache sizes raise the effective threshold to a value on the order of
    $f_{c}\;\approx\;\frac{1}{\eta},$
    meaning only queries expected to be accessed more than $\sim 10$ times per session should reliably be cached locally.
At the system level, these rules are consistent with the global average-cost decomposition
$\mathbb{E}[\operatorname{Cost}]\;=\;\frac{|S|}{N}\;+\;\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr],$
and, in an information-theoretic view,
$\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]\;\approx\;H(Q\mid S_{O},S)\\ \;=\;H(Q\mid S_{O})-I(S;Q\mid S_{O}).$ (60)
In practice, precomputations and storage $S$ are valuable exactly to the extent that they increase $I(S;Q\mid S_{O})$—that is, they capture structure about the future query workload—while their one-time cost $|S|$ is amortized over the total number $N$ of queries.
Corollary 4 (Frequency-weighted extreme strategy costs).
Under the conditions of Theorem 5, fix a query $q$ in the information-rich regime with respect to $S_{O}$ and let $H_{q}:=H(q\mid S_{O})$, $f_{q}:=N\cdot P_{Q}(q)$, and $L_{S}:=\log(m_{S}+n_{S})$ as in (24)–(30). Then the following frequency-weighted cost benchmarks hold (up to additive $O(1)$ and multiplicative constant factors depending only on the universal Turing machine, proof system, and encoding scheme, but not on $S_{O}$, $q$, $S$, or $P_{Q}$):
  1. 1.
    Pure storage (answer caching). There exists a strategy $\Pi_{\mathrm{store}}$ that stores a near-optimally compressed representation of the answer to $q$ of length
    $|S|\;\leq\;H_{q}+O(1),$
    and answers $q$ from $S_{O}\cup S$ with $O(1)$ residual derivation cost:
    $H_{\mathrm{derive}}(q\mid S_{O}\cup S)\;=\;O(1).$
    The resulting amortized cost per access satisfies
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{store}};q)\\ \;=\;\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\;\leq\;\frac{H_{q}}{f_{q}}+O(1).$ (61)
    In particular, whenever $f_{q}\gtrsim L_{S}$ (i.e., $f_{q}$ exceeds the critical frequency scale $f_{c}\asymp L_{S}$ of Theorem 5), this pure-storage protocol matches, up to constant factors, the high-frequency bound
    $\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\lesssim\;C\,\frac{H_{q}}{L_{S}}$
    by making $\frac{H_{q}}{f_{q}}$ the dominant term.
  2. 2.
    Pure computation (no auxiliary storage). Consider the strategy $\Pi_{\mathrm{comp}}$ with $S=\varnothing$. Then, by Theorem 4 applied to $S_{O}$ alone, there exists a constant $C_{0}>0$ such that
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{comp}};q)\;=\;H_{\mathrm{derive}}(q\mid S_{O})\\ \;=\;\Theta\!\left(\frac{H_{q}}{\log\bigl(|\text{Atom}(S_{O})|+\text{Dd}(q\mid S_{O})\bigr)}\right)\\ \;\leq\;C_{0}\frac{H_{q}}{L_{S}}+O(1),$ (62)
    where the last inequality uses that $L_{S}=\log(m_{S}+n_{S})$ is of the same polynomial order as $\log\bigl(|\text{Atom}(S_{O})|+\text{Dd}(q\mid S_{O})\bigr)$ under the structural regularity assumptions (SA2). In particular, when $f_{q}\lesssim L_{S}$, this pure-computation protocol matches, up to constant factors, the low-frequency bound
    $\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\lesssim\;\frac{H_{q}}{f_{q}}.$
  3. 3.
    Near-critical hybrid strategies. For queries with access frequency $f_{q}$ on the order of the critical scale $f_{c}\asymp L_{S}$ (Theorem 5), there exist hybrid strategies $\Pi_{\mathrm{hyb}}=(S,C,R)$ that balance storage and computation so that
    $\operatorname{Cost}_{\mathrm{amort}}(\Pi_{\mathrm{hyb}};q)\\ \;=\;\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\\ \;=\;H_{q}\Bigl(1-\Theta\!\bigl(\tfrac{1}{L_{S}}\bigr)\Bigr)+o(H_{q}),$ (63)
    i.e., they reduce the effective conditional entropy $H(q\mid S_{O},S)$ to a $1-\Theta(1/L_{S})$ fraction of $H_{q}$ while keeping $\frac{|S|}{f_{q}}$ of order $H_{q}/L_{S}$. Such strategies achieve (up to lower-order terms) the same asymptotic amortized cost as the upper bound in Theorem 5 at $f_{q}\asymp f_{c}$.
  4. 4.
    Frequency stratification (policy-level view). For a query distribution $P_{Q}$ over a finite query set $Q$, the frequency-weighted tradeoff of Theorem 5 induces an asymptotically optimal stratification of queries into three regimes:
    • High-frequency class ($f_{q}\gg f_{c}$): precompute and store (near-)complete answers or dense summaries (pure- or near-pure storage). The expected cost is dominated by
      $\mathbb{E}[\operatorname{Cost}(q)]\;\approx\;\mathbb{E}\Bigl[\tfrac{H_{q}}{f_{q}}\Bigr],$
      i.e., per-query costs decay inversely with frequency.
    • Low-frequency class ($f_{q}\ll f_{c}$): rely on on-demand derivation from $S_{O}$ (pure or near-pure computation). The expected cost is dominated by
      $\mathbb{E}[\operatorname{Cost}(q)]\\ \;\approx\;\mathbb{E}\Bigl[\frac{H_{q}}{\log\bigl(|\text{Atom}(S_{O})|+\text{Dd}(q\mid S_{O})\bigr)}\Bigr],$ (64)
      reflecting the logarithmic computational compression factor of Theorem 1.
    • Near-critical class ($f_{q}\asymp f_{c}$): allocate storage to intermediate facts or partial summaries for $q$ (hybrid strategies), achieving amortized costs of order
      $\mathbb{E}[\operatorname{Cost}(q)]\;\approx\;\mathbb{E}\bigl[H_{q}\bigr]\cdot\Bigl(1-\Theta\!\bigl(\tfrac{1}{L_{S}}\bigr)\Bigr),$
      i.e., constant-factor improvements over naive full recomputation or full storage, with diminishing returns as the system size/depth grows.
Remark 8 (Relationship to the non-frequency-weighted formulation).
The frequency-weighted Theorem 5 refines and generalizes the original storage–computation duality bound that did not explicitly account for access frequencies. In its simplest (single-execution) form, the original bound can be summarized schematically as
$\operatorname{Cost}_{\mathrm{sem}}(\Pi;q)\;:=\;|S|+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\\ \;\gtrsim\;H_{q}\left(1+\frac{1}{c\,\log|\text{Atom}(S_{O})|}\right)-o(H_{q}),$ (65)
for some constant $c>0$, where $\operatorname{Cost}_{\mathrm{sem}}$ is a one-shot semantic cost for answering $q$ (no frequency amortization), and the right-hand side reflects the fact that, in the information-rich regime, derivation depth per nat of semantic content cannot be compressed below a logarithmic factor in the size/depth of the underlying theory.
By contrast, Theorem 5 incorporates the access frequency $f_{q}=N\cdot P_{Q}(q)$ over a horizon of $N$ queries and analyzes the amortized cost per access
$\operatorname{Cost}_{\mathrm{amort}}(\Pi;q)\;:=\;\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S).$ (66)
It shows that, up to constant factors and lower-order terms,
$\operatorname{Cost}_{\mathrm{amort}}^{\star}(q)\;\lesssim\;\max\!\left\{\frac{H_{q}}{f_{q}},\,C\frac{H_{q}}{L_{S}}\right\}+O(1),$ (67)
for some constant $C>0$, where the two terms correspond to:
  • a storage-dominated regime $\frac{H_{q}}{f_{q}}$, in which a near-optimal code for $q$ is stored once and reused across $f_{q}$ accesses;
  • a computation-dominated regime $CH_{q}/L_{S}$, in which the answer is derived from $S_{O}\cup S$ on demand, benefiting from the logarithmic compression factor $L_{S}=\log(m_{S}+n_{S})$ identified in Theorem 1.
The critical (break-even) frequency scale is therefore
$f_{c}\;\asymp\;L_{S},$
at which the two contributions $\frac{H_{q}}{f_{q}}$ and $CH_{q}/L_{S}$ are of the same order. Queries with $f_{q}\gg f_{c}$ are asymptotically storage-efficient (caching dominates), while queries with $f_{q}\ll f_{c}$ are asymptotically computation-efficient (on-demand derivation dominates).
The original non-frequency-weighted bound (65) corresponds effectively to the single-execution case $f_{q}=1$. In that regime, the amortized cost per access coincides with the one-shot cost up to constant factors, and the storage term $|S|$ is not meaningfully amortized. In the frequency-weighted setting, by contrast, the same total cost over $M$ executions of $q$,
$\mathbb{E}[\mathrm{Total\ Cost}]\;=\;|S|+M\cdot H_{\mathrm{derive}}(q\mid S_{O}\cup S),$
leads, upon division by $M$, to the amortized form
$\mathbb{E}[\operatorname{Cost}]\;=\;\frac{|S|}{M}+H_{\mathrm{derive}}(q\mid S_{O}\cup S),$
which recovers (66) when $M=f_{q}$. Thus, the frequency-weighted formulation is the natural extension of the original duality bound to long-running systems in which queries are issued multiple times under a distribution $P_{Q}$.
Conceptually, both views share the same structural insight:
  1. 1.
    storage $S$ is a one-time investment that can reduce the conditional algorithmic entropy $H(q\mid S_{O},S)$ (and hence $H_{\mathrm{derive}}$) only to the extent that it encodes information about $q$ beyond $S_{O}$, as quantified by the mutual information $I(S;Q\mid S_{O})$ (cf. Remark 6);
  2. 2.
    derivation cost scales, in the information-rich and polynomial-depth regime, like $H_{q}$ divided by a logarithmic factor in the size/depth of the theory, as captured by Theorem 1;
  3. 3.
    the introduction of $f_{q}$ merely clarifies when it is preferable to pay a larger storage cost to reduce per-access computation, namely when $f_{q}$ exceeds the critical scale $f_{c}\asymp L_{S}$.
In this sense, the non-frequency-weighted and frequency-weighted formulations are compatible limits of the same underlying storage–computation tradeoff: the former optimizes a single execution ($f_{q}=1$), while the latter optimizes repeated executions under a query distribution.
III-D Mutual Information and Submodularity
The system-wide cost decomposition in Section III.C shows that, for a query workload $Q\sim P_{Q}$ and a storage set $S$, the expected per-query cost can be written as
$\mathbb{E}[\text{Cost}]\;=\;\frac{|S|}{N}\;+\;\mathbb{E}_{q\sim P_{Q}}\!\big[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\big]\\ \;\approx\;\frac{|S|}{N}+H(Q\mid S_{O},S),$ (68)
where $H(\cdot\mid\cdot)$ denotes Shannon conditional entropy. Using the identity
$H(Q\mid S_{O},S)\;=\;H(Q\mid S_{O})-I(S;Q\mid S_{O}),$ (69)
we see that, under a fixed base knowledge $S_{O}$ and query distribution $P_{Q}$, choosing an effective storage set $S$ amounts to maximizing the conditional mutual information $I(S;Q\mid S_{O})$ under storage constraints. Intuitively, $I(S;Q\mid S_{O})$ measures how much additional structure about the future query workload $Q$ is captured by $S$ beyond what is already encoded in $S_{O}$. In this subsection we study mutual information as a set function of $S$, and show that it is submodular, which enables efficient approximate optimization of storage allocation.
We first formalize an efficiency measure that quantifies the information-theoretic benefit of storage per unit capacity consumed.
Definition 10 (Storage efficiency via conditional mutual information).
Fix a base knowledge $S_{O}$ and a query distribution $P_{Q}$ over a discrete query space $\mathcal{Q}$. For any storage set $S$ (viewed as a random object induced by a storage policy), we define its storage efficiency as
$\eta(S)\;:=\;\frac{I(S;Q\mid S_{O})}{|S|},$ (70)
where $I(S;Q\mid S_{O})$ is the conditional mutual information between $S$ and $Q$ given $S_{O}$, and $|S|$ is the bit-length of an effective encoding of $S$ as in assumption (SA1).
The upper bound $\eta(S)\leq 1$ follows from the data-processing inequality and assumption (SA1), which ensures $|S|\leq H(S)+O(1)$. In particular, $\eta(S)=1$ if and only if $S$ is a sufficient statistic for $Q$ given $S_{O}$ and the encoding of $S$ is Shannon-optimal.
The numerator $I(S;Q\mid S_{O})$ in (70) measures how much the storage $S$ reduces the Shannon uncertainty of the query workload beyond what is already known from $S_{O}$:
$I(S;Q\mid S_{O})\;=\;H(Q\mid S_{O})-H(Q\mid S_{O},S).$ (71)
A larger value of $I(S;Q\mid S_{O})$ means that the residual entropy $H(Q\mid S_{O},S)$—and hence the expected inference burden in (68)—is smaller.
The denominator $|S|$ corresponds to the one-time storage cost of $S$ (in bits). Thus $\eta(S)$ can be interpreted as the information gain per unit storage: a larger $\eta(S)$ indicates that each bit of storage allocated to $S$ carries more predictive information about the query workload $Q$ beyond the base knowledge $S_{O}$.
We can make this interpretation explicit in the idealized limit where $S$ serves as a sufficient statistic for $Q$ relative to $S_{O}$.
Proposition 2 (Efficiency and sufficiency).
Fix $S_{O}$ and a query distribution $P_{Q}$. Suppose $S$ is a sufficient statistic for $Q$ given $S_{O}$ in the sense that
$H(Q\mid S_{O},S)=0.$ (72)
Then
$I(S;Q\mid S_{O})=H(Q\mid S_{O}),$ (73)
and hence the storage efficiency
$\eta(S)\;=\;\frac{H(Q\mid S_{O})}{|S|}$ (74)
attains its maximal possible value under the given encoding of $S$.
By the definition of conditional mutual information we have
$I(S;Q\mid S_{O})\;=\;H(Q\mid S_{O})-H(Q\mid S_{O},S).$ (75)
If $S$ is sufficient for $Q$ given $S_{O}$ in the sense of (72), then $H(Q\mid S_{O},S)=0$, and therefore
$I(S;Q\mid S_{O})\;=\;H(Q\mid S_{O}).$ (76)
Since, for any encoding of $S$, we have the generic upper bound
$I(S;Q\mid S_{O})\;\leq\;H(S\mid S_{O})\;\leq\;H(S)\;\leq\;|S|,$ (77)
it follows that, for a fixed encoding of $S$, the ratio $I(S;Q\mid S_{O})/|S|$ is maximized if and only if $I(S;Q\mid S_{O})=H(Q\mid S_{O})$. Under (72), this yields
$\eta(S)\;=\;\frac{I(S;Q\mid S_{O})}{|S|}\;=\;\frac{H(Q\mid S_{O})}{|S|},$ (78)
which is the maximal efficiency consistent with the chosen encoding and storage budget. This proves the claim. ∎
Proposition 2 shows that, in the ideal case where $S$ captures all the uncertainty of $Q$ beyond $S_{O}$, the storage efficiency $\eta(S)$ reaches its conceptual maximum, with each stored bit contributing as much predictive information as the encoding allows. In realistic systems, $S$ will only capture part of the structure of $P_{Q}$, leading to intermediate values of $\eta(S)$ that quantify the tradeoff between storage capacity and workload-specific informativeness.
From the system-level viewpoint of Section III.C, a natural storage allocation problem is:
Given a storage budget $M_{\max}$ (in bits), select a storage set $S$ so as to minimize the expected per-query cost $\mathbb{E}[\text{Cost}]$ in (68), equivalently to maximize the conditional mutual information $I(S;Q\mid S_{O})$ subject to $|S|\leq M_{\max}$.
To analyze this optimization problem we now show that $I(S;Q\mid S_{O})$, viewed as a set function of $S$, is submodular. This structure underlies the greedy approximation guarantee stated later.
Throughout this subsection we treat $S$ as a finite subset of some ground set $\mathcal{U}$ indexing elementary storage items (e.g., candidate facts, cached answers, or intermediate derivations). For any $S\subseteq\mathcal{U}$, we write $I(S;Q\mid S_{O})$ as a shorthand for $I(X_{S};Q\mid S_{O})$, where $X_{S}$ denotes the random vector of storage variables corresponding to the items in $S$.
Proposition 3 (Submodularity of conditional mutual information).
Fix a base knowledge $S_{O}$ and a query random variable $Q$. For any two storage sets $S_{1},S_{2}\subseteq\mathcal{U}$, the conditional mutual information satisfies
$I(S_{1}\cup S_{2};Q\mid S_{O})\;+\;I(S_{1}\cap S_{2};Q\mid S_{O})\\ \;\leq\;I(S_{1};Q\mid S_{O})\;+\;I(S_{2};Q\mid S_{O}).$ (79)
In particular, the set function
$f(S):=I(S;Q\mid S_{O})$ (80)
is submodular in $S$.
Recall that, for any random variables $X,Q$ and conditioning variable $Z$, the conditional mutual information is
$I(X;Q\mid Z)\;=\;H(Q\mid Z)-H(Q\mid Z,X),$ (81)
where $H(\cdot\mid\cdot)$ denotes conditional entropy.
Let us denote, for convenience,
$A:=S_{1},\qquad B:=S_{2}.$ (82)
Then
$I(A\cup B;Q\mid S_{O})$ $=H(Q\mid S_{O})-H(Q\mid S_{O},A\cup B),$ (83)
$I(A\cap B;Q\mid S_{O})$ $=H(Q\mid S_{O})-H(Q\mid S_{O},A\cap B),$ (84)
$I(A;Q\mid S_{O})$ $=H(Q\mid S_{O})-H(Q\mid S_{O},A),$ (85)
$I(B;Q\mid S_{O})$ $=H(Q\mid S_{O})-H(Q\mid S_{O},B).$ (86)
Summing the left-hand side of (79) and using (83)–(84) yields
$I(A\cup B;Q\mid S_{O})+I(A\cap B;Q\mid S_{O})$
$\qquad=2H(Q\mid S_{O})-H(Q\mid S_{O},A\cup B)$
$\qquad-H(Q\mid S_{O},A\cap B).$ (87)
Similarly, summing the right-hand side of (79) and using (85)–(86) gives
$I(A;Q\mid S_{O})+I(B;Q\mid S_{O})$
$\qquad=2H(Q\mid S_{O})-H(Q\mid S_{O},A)-H(Q\mid S_{O},B).$ (88)
Therefore, the inequality (79) is equivalent to
$H(Q\mid S_{O},A\cup B)\;+\;H(Q\mid S_{O},A\cap B)$
$\qquad\;\geq\;H(Q\mid S_{O},A)\;+\;H(Q\mid S_{O},B).$ (89)
That is, we need to show that the mapping
$S\;\mapsto\;H(Q\mid S_{O},S)$ (90)
is supermodular in $S$.
This is a standard property of conditional entropy. Intuitively, as we condition on a larger set of variables, the marginal benefit of conditioning on any additional variable decreases, which is precisely the notion of supermodularity for $H(Q\mid S_{O},S)$ and, via (81), submodularity for $I(S;Q\mid S_{O})$.
More formally, for any $A\subseteq B\subseteq\mathcal{U}$ and any element $x\in\mathcal{U}\setminus B$, the marginal reduction in conditional entropy due to conditioning on $x$ satisfies
$H(Q\mid S_{O},A)-H(Q\mid S_{O},A\cup\{x\})$
$\quad=I(x;Q\mid S_{O},A),$ (91)
and
$H(Q\mid S_{O},B)-H(Q\mid S_{O},B\cup\{x\})$
$\quad=I(x;Q\mid S_{O},B).$ (92)
Since conditioning cannot increase mutual information, we have
$I(x;Q\mid S_{O},B)\;\leq\;I(x;Q\mid S_{O},A),$ (93)
which implies that the marginal entropy reduction $H(Q\mid S_{O},S)-H(Q\mid S_{O},S\cup\{x\})$ is decreasing as $S$ grows. This is exactly the defining property of supermodularity of $H(Q\mid S_{O},S)$ in $S$, which in turn is equivalent to the submodularity [fujishige1978polymatroidal] of $I(S;Q\mid S_{O})$ via (81)–(89). This proves (79) and establishes that $f(S)$ in (80) is submodular. ∎
Proposition 3 shows that the incremental utility of adding a storage item to SS—measured by the gain in conditional mutual information $\Delta f:=I(S\cup\{x\};Q\mid S_{O})-I(S;Q\mid S_{O})$—exhibits diminishing returns as $S$ grows. This property is fundamentally important for storage allocation, because it implies that simple greedy procedures enjoy strong approximation guarantees when maximizing $I(S;Q\mid S_{O})$ under a budget constraint.
We now state a direct consequence of Proposition 3 for the optimization problem induced by the system-level cost decomposition (68)–(69).
Corollary 5 (Greedy approximation for storage selection).
Consider the problem of selecting a storage set $S\subseteq\mathcal{U}$ under a budget constraint
$|S|\;\leq\;M_{\max},$ (94)
so as to maximize the conditional mutual information
$f(S):=I(S;Q\mid S_{O}),$ (95)
where $S_{O}$ and $P_{Q}$ are fixed. Assume that the ground set $\mathcal{U}$ is finite, that $f(\varnothing)=0$, and that $f(S)$ is nonnegative and monotone non-decreasing in $S$. Then the standard greedy algorithm that iteratively adds the storage item with the largest marginal gain in $f(S)$ subject to the remaining budget achieves a (11/e)(1-1/e)-approximation to the optimal value:
$f(S_{\mathrm{greedy}})\;\geq\;(1-1/e)\,f(S^{\star}),$ (96)
where $S^{\star}$ is an optimal solution with $|S^{\star}|\leq M_{\max}$, and $S_{\mathrm{greedy}}$ is the solution returned by the greedy procedure.
By Proposition 3, the function $f(S)=I(S;Q\mid S_{O})$ is submodular in $S$. Nonnegativity and monotonicity follow from basic properties of mutual information:
$0\;\leq\;I(S;Q\mid S_{O})\;\leq\;I(S\cup T;Q\mid S_{O})$ (97)
whenever $S\subseteq S\cup T$. Therefore, $f(S)$ satisfies the standard conditions of the classical result on maximizing a nonnegative monotone submodular function under a cardinality or knapsack constraint. By the well-known theorem of Nemhauser et al. [nemhauser1978analysis], the greedy algorithm that repeatedly adds the element with largest marginal gain in $f(S)$ subject to the remaining budget achieves a (11/e)(1-1/e)-approximation to the optimal value. Applying this general result to the specific choice $f(S)=I(S;Q\mid S_{O})$ yields the claimed inequality. ∎
Corollary 5 provides a principled algorithmic prescription for system-wide storage allocation: under the amortized cost model of Section III.C, one should select storage items greedily in order of their marginal contribution to $I(S;Q\mid S_{O})$ per unit of storage cost, subject to the overall budget $M_{\max}$. Up to the universal (11/e)(1-1/e) factor, this strategy is provably as effective as any other storage allocation policy in terms of reducing the expected inference burden $\mathbb{E}_{q\sim P_{Q}}[H_{\mathrm{derive}}(q\mid S_{O}\cup S)]$ and, equivalently, the Shannon term $H(Q\mid S_{O},S)$ in (68).
IV THERMODYNAMIC CONSTRAINTS AND PHYSICAL LIMITS
This section establishes the physical constraints on information processing in knowledge systems, connecting our information-theoretic framework to thermodynamics via Landauer’s principle and the physical cost assumptions of Section II.
IV-A Landauer’s Principle and Energy Bounds
We now relate derivation depth and storage size to minimal energy requirements under the classical thermodynamic regime of Assumption 3.
Recall that, in this regime, each logically irreversible operation that commits or erases one bit of information dissipates at least $k_{B}T\ln 2$ joules of heat, where $k_{B}$ is Boltzmann’s constant and $T$ is the operating temperature. In Section III-A we fixed the per-step information cost to be $\lambda\;=\;\ln 2\ \text{nats}$, so that each inference step in a derivation is modeled as committing or erasing $\Theta(1)$ bit of information, and, in particular, we normalize units so that each step incurs a Landauer cost of at least $k_{B}T\ln 2$.
Proposition 4 (Landauer energy bound for computation).
Under the assumption that each inference step constitutes a logically irreversible operation in the classical thermodynamic regime (Assumption 3), deriving a formula $\varphi$ from a knowledge base $S_{O}$ requires at least the following energy:
$E_{\mathrm{compute}}(\varphi\mid S_{O})\;\geq\;Dd(\varphi\mid S_{O})\,\cdot\,k_{B}T\ln 2,$ (98)
where $Dd(\varphi\mid S_{O})$ is the derivation depth of $\varphi$ from $S_{O}$ in the sense of Definition 6.
By Assumption 3 and the discussion above, we model each inference step as a logically irreversible operation that commits or erases at least one bit of information. By Landauer’s principle [landauer1961irreversibility], each such operation incurs an energy cost of at least $k_{B}T\ln 2$.
Consider a derivation of $\varphi$ from $S_{O}$ of minimal derivation depth $Dd(\varphi\mid S_{O})$. This derivation consists of exactly $Dd(\varphi\mid S_{O})$ inference steps. Since each step is modeled as performing at least one bit’s worth of logically irreversible information processing, the total number of bit erasures (or committed bits) across the derivation is lower bounded by
$\text{bit-erasures}\;\geq\;Dd(\varphi\mid S_{O}).$
By Landauer’s principle, the total energy dissipated by these operations thus satisfies
$E_{\mathrm{compute}}(\varphi\mid S_{O})\;\geq\;Dd(\varphi\mid S_{O})\,k_{B}T\ln 2,$
which is exactly (98). ∎
In terms of the derivation entropy defined in Section III-A, we immediately obtain a compact form.
Remark 9 (Derivation entropy as thermodynamic cost).
Under Assumption 3 we have, by Corollary 3,
$H_{\mathrm{derive}}(\varphi\mid S_{O})\;=\;Dd(\varphi\mid S_{O})\,\ln 2\quad\text{(in nats)}.$
Combining this with Proposition 4 yields
$E_{\mathrm{compute}}(\varphi\mid S_{O})\;\geq\;k_{B}T\,H_{\mathrm{derive}}(\varphi\mid S_{O}),$ (99)
so that derivation entropy directly lower-bounds the minimal thermodynamic energy required to carry out the corresponding inference.
We now complement the dynamic computation cost with an energy bound for maintaining stored information over time.
Proposition 5 (Energy bound for maintaining storage).
In a simple Arrhenius model of thermally noisy storage, maintaining a stored state $S$ with $|S|$ effective bits over a time horizon $t$ requires at least
$E_{\mathrm{storage}}(S,t)\;\gtrsim\;|S|\,k_{B}T\ln 2\cdot\frac{t}{t_{\mathrm{refresh}}},$ (100)
where $t_{\mathrm{refresh}}$ denotes the characteristic refresh interval implied by the thermal flip rate of the storage medium.
We adopt a coarse-grained Arrhenius model for thermal bit flips. Let $E_{\mathrm{barrier}}$ denote the energy barrier separating the two logical states of a physical bit cell. Classical transition-state theory then suggests that the bit-flip rate satisfies
$\gamma_{\mathrm{flip}}\;\approx\;\gamma_{0}\exp\!\Bigl(-\frac{E_{\mathrm{barrier}}}{k_{B}T}\Bigr),$
for some attempt frequency $\gamma_{0}$ [atkins2014physical]. The characteristic time between spontaneous flips is thus
$t_{\mathrm{refresh}}\;\approx\;\frac{1}{\gamma_{\mathrm{flip}}},$
which we interpret as the required refresh interval to reliably maintain the stored logical state against thermal noise.
In a simple refresh model, each refresh cycle rewrites all $|S|$ bits in $S$ to their intended logical values. By Landauer’s principle, each rewritten bit incurs at least $k_{B}T\ln 2$ of energy dissipation, so a single refresh of $S$ requires at least
$E_{\mathrm{per\;refresh}}\;\geq\;|S|\,k_{B}T\ln 2.$
Over a total time horizon $t$, the number of such refresh cycles is on the order of $t/t_{\mathrm{refresh}}$. Hence the total energy dissipated in maintaining $S$ satisfies
$E_{\mathrm{storage}}(S,t)\;\gtrsim\;\frac{t}{t_{\mathrm{refresh}}}\,|S|\,k_{B}T\ln 2,$
which is (100). Here the symbol $\gtrsim$ emphasizes that we are working with a lower bound under simplifying assumptions: (i) independent bit flips with Arrhenius rate, and (ii) full rewriting of all $|S|$ bits at each refresh cycle. More sophisticated coding or hardware techniques may improve constant factors but cannot avoid the linear dependence on $|S|$ and $t/t_{\mathrm{refresh}}$ in this regime. ∎
Remark 10 (Link to amortized storage cost).
The storage size $|S|$ already appears in the information-theoretic amortized-cost decomposition of Section III-C, where the expected per-query information cost is
$E[\mathrm{Cost}]\;=\;\frac{|S|}{N}\;+\;E_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr],$
for a workload of $N$ queries drawn from $P_{Q}$. Proposition 5 shows that this bit-length $|S|$ also controls a minimal thermodynamic cost for maintaining the corresponding physical memory:
$\frac{E_{\mathrm{storage}}(S,t)}{N}\;\gtrsim\;\frac{|S|}{N}\,k_{B}T\ln 2\cdot\frac{t}{t_{\mathrm{refresh}}},$
which is directly analogous to the $\frac{|S|}{N}$ term above. Together with (99), this yields a chain of correspondences
$\text{derivation depth }Dd(\cdot)$ $\;\longrightarrow\;H_{\mathrm{derive}}(\cdot)\ \text{(nats)}$
$\;\longrightarrow\;E_{\mathrm{compute}}(\cdot)\ \text{(J)},$
and
$\text{storage size }|S|\ \text{(bits)}$ $\;\longrightarrow\;\frac{|S|}{N}\ \text{(bits/query)}$
$\;\longrightarrow\;\frac{E_{\mathrm{storage}}(S,t)}{N}\ \text{(J/query)}.$
Thus, the storage–computation tradeoffs derived in Section III not only constrain algorithmic and information-theoretic costs, but also translate into minimal thermodynamic costs in physically implemented knowledge systems.
IV-B Storage-Computation Energy Tradeoff
Having established the information-theoretic foundations of mixed strategies in Section III, we now derive fundamental lower bounds on the total energy cost of semantic information processing. These results connect the abstract information measures to physical thermodynamic constraints.
We consider a knowledge system operating over a time horizon during which $N$ queries are processed. Let $P_{Q}$ be a computable probability distribution over queries, and let $Q$ denote a random query drawn from $P_{Q}$. We assume the system operates in the classical thermodynamic regime (Assumption 3).
Definition 11 (Amortized Energy Cost).
For a mixed strategy $\Pi=(S,C,R)$ processing $N$ queries over time $t$, the amortized energy cost per query is defined as:
$\bar{E}(\Pi):=\frac{E_{\mathrm{storage}}(S,t)}{N}+\mathbb{E}_{q\sim P_{Q}}\left[E_{\mathrm{compute}}(q|S_{O}\cup S)\right]$ (101)
where $E_{\mathrm{storage}}(S,t)$ is the total energy required to maintain storage $S$ for time $t$, and $E_{\mathrm{compute}}(q|S_{O}\cup S)$ is the energy required to derive the answer to process query $q$ from the augmented knowledge base $S_{O}\cup S$.
Throughout this section, we assume the system operates in the standard refresh regime, meaning that the time horizon $t$ and query count $N$ satisfy $t/(N\cdot t_{refresh})=\Theta(1)$, i.e., the storage is refreshed a constant number of times per query on average.
Definition 12 (Normalized Information Cost).
For a mixed strategy $\Pi=(S,C,R)$, the normalized information cost per query is:
$\bar{I}(\Pi):=\frac{|S|}{N}+\mathbb{E}_{q\sim P_{Q}}\left[H_{\mathrm{derive}}(q|S_{O}\cup S)\right]$ (102)
where $|S|$ denotes the size of auxiliary storage in bits, and $H_{\mathrm{derive}}(q|S_{O}\cup S)=D_{d}(q|S_{O}\cup S)\cdot\ln 2$ is the derivation entropy.
The normalized information cost $\bar{I}(\Pi)$ serves as an information-theoretic proxy for energy consumption, with the conversion to physical energy units given by Propositions 4 and 5.
We now state and prove the main theorem of this section, which establishes a fundamental lower bound on the resources required by any mixed strategy.
Assumption 4 (Uniform Polynomial-Depth Regime).
Let $P_{Q}$ be a query distribution over a query set $\mathcal{Q}$ with respect to knowledge base $S_{O}$. We say that $P_{Q}$ satisfies the uniform polynomial-depth regime with respect to a mixed strategy $\Pi=(S,C,R)$ if there exists a polynomial $p(\cdot)$ such that
$D_{d}^{\sup}(P_{Q}\mid S_{O}\cup S):=\\ \sup_{q\in\mathrm{supp}(P_{Q})}D_{d}(q\mid S_{O}\cup S)\leq p\bigl(|\mathrm{Atom}(S_{O}\cup S)|\bigr).$ (103)
In particular, writing $m_{S}:=|\mathrm{Atom}(S_{O}\cup S)|$, we have $\log D_{d}^{\sup}=O(\log m_{S})$ and hence
$L_{S}:=\log(m_{S}+D_{d}^{\sup})=\Theta(\log m_{S}).$ (104)
Remark 11 (Scope of Assumption 4).
The uniform polynomial-depth regime is satisfied by most practical query classes, including:
  • Datalog queries with bounded recursion depth [ceri1989you];
  • SPARQL queries over RDF graphs with bounded property path length [perez2009semantics];
  • Template-based natural language queries compiled to first-order logic;
  • Bounded-depth inference in description logics (e.g., $\mathcal{ALC}$, $\mathcal{SHIQ}$).
The key requirement is that the inference depth does not grow super-polynomially with the size of the knowledge base—a condition that holds for essentially all tractable reasoning tasks. More precisely, for these query classes, the derivation depth is bounded by a fixed polynomial in the number of base facts, independent of the query complexity.
Theorem 6 (Storage-Computation Information Tradeoff).
Let $S_{O}$ be a knowledge base and let $P_{Q}$ be a computable query distribution satisfying the information-rich regime (Definition 8). Let $\Pi=(S,C,R)$ be any mixed strategy satisfying assumptions (SA1)–(SA3) and the uniform polynomial-depth regime (Assumption 4).
Define the effective complexity parameter:
$L_{S}:=\log\Bigl(|\mathrm{Atom}(S_{O}\cup S)|+D_{d}^{\sup}(P_{Q}\mid S_{O}\cup S)\Bigr),$ (105)
where $D_{d}^{\sup}(P_{Q}|S_{O}\cup S)$ is as defined in Assumption 4.
If the total number of queries satisfies $N=\Omega(L_{S})$, then there exists a constant $c>0$ depending only on the proof system and encoding scheme such that:
$\bar{I}(\Pi)\geq c\cdot\frac{H(Q\mid S_{O})}{L_{S}}-O(1),$ (106)
where $H(Q\mid S_{O})$ denotes the conditional Shannon entropy of $Q$ given the knowledge base $S_{O}$.
The proof proceeds in five steps.
Step 1: Lower bound on derivation entropy. By the lower bound in Theorem 4, for any query $q\in\text{supp}(P_{Q})$ in the information-rich regime, the derivation entropy satisfies:
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)=D_{d}(q\mid S_{O}\cup S)\cdot\ln 2\\ \geq c_{1}\frac{H(q\mid S_{O}\cup S)}{L(q)}-O(1),$ (107)
where $c_{1}>0$ is the constant from Theorem 4, $H(q\mid S_{O}\cup S):=K(q\mid\langle\mathrm{Atom}(S_{O}\cup S)\rangle)\cdot\ln 2$ is the conditional algorithmic entropy, and
$L(q):=\log\bigl(|\mathrm{Atom}(S_{O}\cup S)|+D_{d}(q\mid S_{O}\cup S)\bigr).$ (108)
Step 2: Uniformization via the polynomial-depth assumption. Under Assumption 4, we have for all $q\in\mathrm{supp}(P_{Q})$:
$D_{d}(q\mid S_{O}\cup S)\leq D_{d}^{\sup}(P_{Q}\mid S_{O}\cup S),$ (109)
and consequently,
$L(q)\leq L_{S}\quad\text{for all }q\in\mathrm{supp}(P_{Q}).$ (110)
Moreover, since both $|\mathrm{Atom}(S_{O}\cup S)|$ and $D_{d}^{\sup}$ are polynomially related under Assumption 4, we have:
$L(q)=\Theta(L_{S})\quad\text{for all }q\in\mathrm{supp}(P_{Q}).$ (111)
Therefore, there exists a constant $c_{1}^{\prime}>0$ (absorbing the $\Theta(\cdot)$ factor into $c_{1}$) such that for all $q\in\mathrm{supp}(P_{Q})$:
$H_{\mathrm{derive}}(q\mid S_{O}\cup S)\geq c_{1}^{\prime}\frac{H(q\mid S_{O}\cup S)}{L_{S}}-O(1).$ (112)
Taking expectations over $q\sim P_{Q}$:
$\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]\\ \geq c_{1}^{\prime}\frac{\mathbb{E}_{q\sim P_{Q}}\bigl[H(q\mid S_{O}\cup S)\bigr]}{L_{S}}-O(1),$ (113)
where the $O(1)$ term remains $O(1)$ after expectation since it is uniform over all $q\in\mathrm{supp}(P_{Q})$.
Step 3: Relating Kolmogorov complexity to Shannon entropy. For a computable distribution $P_{Q}$, the expected Kolmogorov complexity relates to Shannon entropy via the classical correspondence (see, e.g., [LiVitanyi2019Kolmogorov, cover2006elements]):
$\mathbb{E}_{q\sim P_{Q}}\bigl[K(q\mid S_{O}\cup S)\bigr]\geq H(Q\mid S_{O},S)-O(1),$ (114)
where $H(Q\mid S_{O},S)$ is the conditional Shannon entropy of $Q$ given both $S_{O}$ and $S$, and the $O(1)$ term accounts for the description complexity of the distribution $P_{Q}$ (which is assumed computable and hence has finite Kolmogorov complexity).
By the chain rule for Shannon entropy:
$H(Q\mid S_{O},S)=H(Q\mid S_{O})-I(S;Q\mid S_{O}),$ (115)
where $I(S;Q\mid S_{O})$ is the conditional mutual information between the auxiliary storage $S$ and the query $Q$ given $S_{O}$.
Combining with (113):
$\mathbb{E}\bigl[H_{\mathrm{derive}}(Q\mid S_{O}\cup S)\bigr]\\ \geq c_{1}^{\prime}\frac{H(Q\mid S_{O})-I(S;Q\mid S_{O})}{L_{S}}-O(1).$ (116)
Step 4: Bounding the mutual information. By assumption (SA1), the auxiliary storage $S$ is a valid logical encoding, which implies:
$I(S;Q\mid S_{O})\leq H(S\mid S_{O})\leq H(S)\leq|S|,$ (117)
where the first inequality is the standard property that mutual information is bounded by marginal entropy, the second follows from conditioning reducing entropy, and the third follows from the fact that Shannon entropy is bounded by the description length.
Step 5: Combining storage and computation costs. From the definition of $\bar{I}(\Pi)$ and inequalities (116) and (117):
$\bar{I}(\Pi)=\frac{|S|}{N}+\mathbb{E}[H_{\mathrm{derive}}(Q|S_{O}\cup S)]\\ \geq\frac{|S|}{N}+c_{1}^{\prime}\frac{H(Q|S_{O})-|S|}{L_{S}}-O(1).$ (118)
We analyze this expression by considering the extremal cases.
Case (a): Pure computation ($|S|=0$). In this case,
$\bar{I}(\Pi)\geq c_{1}^{\prime}\frac{H(Q|S_{O})}{L_{S}}-O(1).$
Case (b): Pure storage ($|S|=H(Q|S_{O})+O(1)$). By assumption (SA1), any useful storage satisfies $|S|\leq H(Q|S_{O})+O(\log|S_{O}|)$. Taking $|S|=H(Q|S_{O})$, we obtain
$\bar{I}(\Pi)\geq\frac{H(Q|S_{O})}{N}+0-O(1)=\frac{H(Q|S_{O})}{N}-O(1).$
Under $N=\Omega(L_{S})$, this yields $\bar{I}(\Pi)\geq\Omega(H(Q|S_{O})/L_{S})-O(1)$.
General case. For intermediate values of $|S|$, the right-hand side of (118) is a linear function of $|S|$ with slope $\frac{1}{N}-\frac{c_{1}^{\prime}}{L_{S}}$. Since $N=\Omega(L_{S})$, this slope is $O(1/L_{S})$ in magnitude. In either case (slope positive or negative), the minimum of the bound over $|S|\in[0,H(Q|S_{O})]$ is achieved at one of the endpoints, which we have already analyzed.
Combining Cases (a) and (b), there exists a constant $c>0$ such that
$\bar{I}(\Pi)\geq c\cdot\frac{H(Q|S_{O})}{L_{S}}-O(1),$
completing the proof. ∎
Remark 12 (Role of the Supremum in $L_{S}$).
The use of $D_{d}^{\sup}(P_{Q}\mid S_{O}\cup S)$ rather than, say, the expected value $\mathbb{E}[D_{d}(Q\mid S_{O}\cup S)]$ in the definition of $L_{S}$ serves two purposes:
  1. 1.
    Uniformity: The lower bound (106) holds for all queries in the support of $P_{Q}$, not just on average. This is essential for worst-case guarantees.
  2. 2.
    Asymptotic equivalence: Under the polynomial-depth regime (Assumption 4), the logarithmic factor satisfies
    $\log\bigl(m+\mathbb{E}[D_{d}]\bigr)=\Theta\bigl(\log(m+D_{d}^{\sup})\bigr)=\Theta(\log m),$
    where $m=|\mathrm{Atom}(S_{O}\cup S)|$. Thus, the choice between supremum and expectation affects only the implicit constants, not the asymptotic behavior.
In practice, for query distributions concentrated on a polynomial-depth class, the distinction is immaterial.
Remark 13 (Tightness of the Bound).
The lower bound in Theorem 6 is tight up to constant factors. By Theorem 5, there exists a strategy $\Pi^{*}$ achieving:
$\bar{I}(\Pi^{*})\leq C\cdot\frac{H(Q|S_{O})}{L_{S^{*}}}+O(1)$ (119)
for some constant $C>0$, where $L_{S^{*}}$ is the value of $L_{S}$ corresponding to the optimal strategy $\Pi^{*}$. Thus, the optimal normalized information cost satisfies:
$\bar{I}^{*}=\Theta\left(\frac{H(Q|S_{O})}{L^{*}}\right)$ (120)
where $L^{*}=\Theta(L_{S})$ for the optimal strategy.
We now translate the information-theoretic bound into physical energy units using the thermodynamic correspondences established in Section IV-A.
Corollary 6 (Thermodynamic Energy Lower Bound).
Under the conditions of Theorem 6, additionally assuming the classical thermodynamic regime (Assumption 3), the amortized energy cost per query satisfies:
$\bar{E}(\Pi)\geq c\cdot\frac{H(Q|S_{O})}{L_{S}}\cdot k_{B}T\ln 2-O(k_{B}T)$ (121)
where $k_{B}$ is Boltzmann’s constant and $T$ is the operating temperature.
By Proposition 4, each derivation step requires at least $k_{B}T\ln 2$ of energy dissipation, giving:
$E_{\mathrm{compute}}(q|S_{O}\cup S)\geq D_{d}(q|S_{O}\cup S)\cdot k_{B}T\ln 2\\ =H_{derive}(q\mid S_{O}\cup S)\cdot k_{B}T$ (122)
By Proposition 5, maintaining $|S|$ bits for time $t$ requires:
$E_{\mathrm{storage}}(S,t)\geq|S|\cdot k_{B}T\ln 2\cdot\frac{t}{t_{\mathrm{refresh}}}$ (123)
Under the assumption that the system operates in the standard refresh regime, where the average number of refresh cycles per query satisfies $\frac{t}{N\cdot t_{\mathrm{refresh}}}=\Theta(1)$, we have:
$\frac{E_{\mathrm{storage}}(S,t)}{N}=\Theta\left(\frac{|S|}{N}\right)\cdot k_{B}T\ln 2$ (124)
Combining these bounds:
$\bar{E}(\Pi)$
$=\frac{E_{\mathrm{storage}}(S,t)}{N}+\mathbb{E}\left[E_{\mathrm{compute}}(Q|S_{O}\cup S)\right]$
$\geq\Theta(1)\cdot\left(\frac{|S|}{N}+\mathbb{E}\left[H_{\mathrm{derive}}(Q|S_{O}\cup S)\right]\right)\cdot k_{B}T\ln 2$
$=\Theta(1)\cdot\bar{I}(\Pi)\cdot k_{B}T\ln 2$
$\geq c\cdot\frac{H(Q|S_{O})}{L_{S}}\cdot k_{B}T\ln 2-O(k_{B}T)$ (125)
where the last inequality applies Theorem 6. ∎
Corollary 7 (Per-Nit Energy Cost).
The minimum energy required to process one nit (natural unit) of semantic information content is:
$\frac{\bar{E}^{*}}{H(Q|S_{O})}=\Theta\left(\frac{k_{B}T}{L^{*}}\right)$ (126)
where $L^{*}=\Theta(\log|\mathrm{Atom}(S_{O})|)$ under the polynomial-depth regime (Assumption 4).
This follows directly from Corollary 6 and Remark 13 by dividing both sides by $H(Q|S_{O})$. ∎
Theorem 6 and its corollaries establish several fundamental insights:
The Logarithmic Efficiency Factor.
The denominator $L_{S}=\log(|\mathrm{Atom}(S_{O}\cup S)|+D_{d}^{\sup}(P_{Q}|S_{O}\cup S))$ represents a compression factor arising from the logical structure of the knowledge base. This factor quantifies the efficiency gain from exploiting logical relationships rather than storing or computing information independently. In large knowledge systems where $|\mathrm{Atom}(S_{O})|=2^{\Omega(n)}$, this factor can be substantial, yielding energy costs that scale as $O(n)$ rather than $O(2^{n})$.
Comparison with Landauer’s Principle.
The classical Landauer bound states that erasing one bit requires at least $k_{B}T\ln 2$ of energy. Corollary 7 shows that processing one nit of semantic information requires $\Theta(k_{B}T/\log|\mathrm{Atom}|)$ of energy. The additional $1/\log|\mathrm{Atom}|$ factor reflects the efficiency gained through logical inference: derivation compresses information processing by exploiting the structure of the knowledge base.
The Storage-Computation Boundary.
The lower bound (IV-B) treats storage and computation symmetrically through their information-theoretic costs. This symmetry breaks down in the energy interpretation (Corollary 6) only through the refresh rate parameter $t_{\mathrm{refresh}}$, which is a technology-dependent quantity. The information-theoretic formulation thus captures the fundamental tradeoff independent of implementation details.
Regime Conditions.
The assumption $N=\Omega(L_{S})$ ensures that the system processes enough queries to amortize storage costs. When $N=o(L_{S})$, storage costs dominate and the bound must be modified. This corresponds to the intuition that caching is only beneficial when queries are sufficiently frequent relative to the system’s logical complexity.
IV-C Entropy Production and Irreversibility
The energy bounds established in this section can be recast in the language of entropy production, providing a complementary perspective rooted directly in the second law of thermodynamics.
IV-C1 Entropy Production in Query Answering
Consider a knowledge system operating in thermal equilibrium at temperature $T$. When the system answers a query $q$ by deriving it from a knowledge base $\mathcal{S}_{O}$, the computation constitutes an irreversible process that produces entropy in the environment.
Proposition 6 (Entropy Production Bound for Query Answering).
Let $\Pi=(S,C,R)$ be a mixed strategy answering a query $q$ from knowledge base $\mathcal{S}_{O}$ in the classical thermodynamic regime (Assumption 3). The minimum entropy produced in the environment satisfies:
$\Delta S_{\mathrm{env}}(q)\geq k_{B}\ln 2\cdot D_{d}(q\mid\mathcal{S}_{O}\cup S),$ (127)
or equivalently, in terms of derivation entropy:
$\Delta S_{\mathrm{env}}(q)\geq k_{B}\cdot H_{\mathrm{derive}}(q\mid\mathcal{S}_{O}\cup S).$ (128)
By Proposition 4, answering query $q$ requires energy $E_{\mathrm{compute}}(q)\geq D_{d}(q\mid\mathcal{S}_{O}\cup S)\cdot k_{B}T\ln 2$. In the quasistatic limit of the classical thermodynamic regime, this energy is dissipated as heat $Q_{\mathrm{diss}}=E_{\mathrm{compute}}$ into the thermal environment. The corresponding entropy increase in the environment is
$\Delta S_{\mathrm{env}}=\frac{Q_{\mathrm{diss}}}{T}\geq\frac{D_{d}(q\mid\mathcal{S}_{O}\cup S)\cdot k_{B}T\ln 2}{T}\\ =D_{d}(q\mid\mathcal{S}_{O}\cup S)\cdot k_{B}\ln 2.$ (129)
Substituting $H_{\mathrm{derive}}(q\mid\mathcal{S}_{O}\cup S)=D_{d}(q\mid\mathcal{S}_{O}\cup S)\cdot\ln 2$ from Corollary 3 yields (128). ∎
Remark 14 (Duality with Proposition 4).
Proposition 6 is the entropic reformulation of Proposition 4: the energy bound $E\geq D_{d}\cdot k_{B}T\ln 2$ and the entropy bound $\Delta S_{\mathrm{env}}\geq D_{d}\cdot k_{B}\ln 2$ are related by the fundamental thermodynamic identity $\Delta S_{\mathrm{env}}=E/T$ in the quasistatic regime. This duality reflects the interchangeability of energy and entropy as thermodynamic costs at fixed temperature.
IV-C2 Expected Entropy Production and Mutual Information
For a query distribution $P_{Q}$, we can relate the expected entropy production to the conditional mutual information $I(S;Q\mid\mathcal{S}_{O})$ introduced in Section III-D.
Proposition 7 (Expected Entropy Production Bound).
Under the conditions of Proposition 6, let $P_{Q}$ be a query distribution satisfying the information-rich regime (Definition 8) and the uniform polynomial-depth regime (Assumption 4). The expected environmental entropy production satisfies:
$\mathbb{E}_{q\sim P_{Q}}\bigl[\Delta S_{\mathrm{env}}(q)\bigr]\\ \geq\frac{c\cdot k_{B}}{L_{S}}\cdot\bigl(H(Q\mid\mathcal{S}_{O})-I(S;Q\mid\mathcal{S}_{O})\bigr)-O(k_{B}),$ (130)
where $c>0$ is the constant from Theorem 4, $L_{S}=\log(|\mathrm{Atom}(\mathcal{S}_{O}\cup S)|+D_{d}^{\sup})$ is the effective complexity parameter, and $I(S;Q\mid\mathcal{S}_{O})$ is the conditional mutual information between storage $S$ and query $Q$ given $\mathcal{S}_{O}$.
By Proposition 6, $\mathbb{E}[\Delta S_{\mathrm{env}}(q)]\geq k_{B}\cdot\mathbb{E}[H_{\mathrm{derive}}(q\mid\mathcal{S}_{O}\cup S)]$. Applying inequality (116) from the proof of Theorem 6 yields the claimed bound. ∎
This result provides a thermodynamic interpretation of storage utility: each nat of mutual information $I(S;Q\mid\mathcal{S}_{O})$ reduces expected entropy production by $c\cdot k_{B}/L_{S}$. Storage that anticipates future queries thus reduces not only computational cost but also environmental entropy production.
IV-C3 Irreversibility and the Limits of Reversible Computation
Remark 15 (Reversible Computation).
The entropy production bounds above apply to logically irreversible computation. Bennett’s reversible simulation theorem [bennett1973logical] shows that any irreversible computation can be made reversible with polynomial overhead in time and logarithmic overhead in space. For knowledge systems, this implies that the Landauer cost could in principle be avoided by storing $O(D_{d}\log D_{d})$ bits of auxiliary information, assuming the original derivation uses time and space proportional to $D_{d}$. However, this shifts cost from computation to storage without eliminating the fundamental information-theoretic constraints of Theorem 5. The practical relevance of reversible computation for knowledge systems remains limited by engineering challenges and the overhead of reversible simulation.
V Design Rules for Hybrid Storage–Computation Strategies
We now move from the instance-wise and information-theoretic analysis of Section III to higher-level design rules for hybrid strategies that combine storage and computation. Rather than introducing new formal theorems, our goal here is to distill the implications of Theorem 5, the mutual-information analysis of Section III-D, and the thermodynamic constraints of Section IV into simple decision guidelines.
V-A Frequency-Dependent Choice of Storage vs. Computation
The frequency-weighted tradeoff in Theorem 5 provides a natural criterion for deciding, for each query type $q$, whether it should be answered purely by on-demand derivation, purely by storage (caching), or by a hybrid strategy that stores partial information.
Recall the notation from Section III-B:
  • $H_{q}:=H(q\mid S_{O})=K(\text{enc}{q}\mid\text{enc}{\text{Atom}(S_{O})})\ln 2$ is the conditional algorithmic entropy of $q$ given $S_{O}$.
  • $f_{q}:=N\cdot P_{Q}(q)$ is the expected number of accesses of query $q$ over a horizon of $N$ queries (Definition 9).
  • For a storage structure $S$, we denote $m_{S}:=|\text{Atom}(S_{O}\cup S)|$ and $n_{S}:=Dd(q\mid S_{O}\cup S)$, and define
    $L_{S}:=\log(m_{S}+n_{S}),$ (131)
    which plays the role of an effective logarithmic compression factor in Theorem 4 and 5.
Under the assumptions of Theorem 9, there exist constants $C,C>0$ such that for any admissible hybrid strategy $\Pi=(S,C,R)$ serving a fixed query type $q$ we have
$\mathrm{Cost}_{\mathrm{amort}}(\Pi;q)\;=\;\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\\ \;\leq\;\min\left\{\frac{H_{q}}{f_{q}},\;C\,\frac{H_{q}}{L_{S}}\right\}+O(1),$ (132)
where all costs are measured in nats (or bits up to a constant factor).
The two competing terms inside the minimum correspond to two extremal—and constructive—strategies:
  1. 1.
    Pure storage (caching) strategy: store a near-optimally compressed representation of the answer to $q$ of length at most $H_{q}+O(1)$ nats (or equivalent bits) so that $H_{\mathrm{derive}}(q\mid S_{O}\cup S)=O(1)$. This yields an amortized per-access cost $\mathrm{Cost}_{\mathrm{amort}}\lesssim H_{q}/f_{q}+O(1)$.
  2. 2.
    Pure computation (no auxiliary storage) strategy: take $S=\varnothing$ and derive $q$ from $S_{O}$ on demand. By Theorem 4, this yields an amortized cost $\mathrm{Cost}_{\mathrm{amort}}\lesssim CH_{q}/L_{S^{(0)}}+O(1)$, where
    $L_{S^{(0)}}=\log\bigl(|\text{Atom}(S_{O})|+Dd(q\mid S_{O})\bigr)$
    is the baseline compression factor for inference from $S_{O}$ alone.
Comparing these two extremes leads to a natural critical frequency scale. In the baseline case $S=\varnothing$ we define
$f_{c}^{(0)}(q)\;\asymp\;L_{S^{(0)}}\;=\;\log\bigl(|\text{Atom}(S_{O})|+Dd(q\mid S_{O})\bigr).$ (133)
When $f_{q}$ is much smaller than $f_{c}^{(0)}(q)$, the term $H_{q}/f_{q}$ is larger than $CH_{q}/L_{S^{(0)}}$, so it is more efficient to answer $q$ by on-demand derivation from $S_{O}$ (pure computation). Conversely, when $f_{q}$ is much larger than $f_{c}^{(0)}(q)$, the caching term $H_{q}/f_{q}$ dominates the pure computation cost, so precomputing and storing (some representation of) the answer to $q$ becomes more attractive.
This motivates the following rule-of-thumb.
Policy 1 (Frequency-aware caching rule (informal)).
Fix a base knowledge $S_{O}$, a query type $q$, and its expected access frequency $f_{q}=N\cdot P_{Q}(q)$ over a horizon of $N$ queries. Let
$f_{c}^{(0)}(q)\;\asymp\;\log\bigl(|\text{Atom}(S_{O})|+Dd(q\mid S_{O})\bigr)$
be the baseline critical frequency as in (133). Then, up to multiplicative constants and lower-order terms:
  • If $f_{q}\ll f_{c}^{(0)}(q)$, the pure computation strategy achieves amortized cost $\Theta(H_{q}/L_{S}^{(0)})$, which is better than the pure storage cost $\Theta(H_{q}/f_{q})$.
  • If $f_{q}\gg f_{c}^{(0)}(q)$, it is typically more efficient to cache the answer to $q$ (or suitable intermediate results), so that most of the cost is paid once in storage and the marginal per-access cost becomes $O(1)$.
  • If $f_{q}\approx f_{c}^{(0)}(q)$, hybrid strategies that store partial information about $q$ (e.g., indices, intermediate derivations) can provide a good balance between storage footprint and computation cost.
Remark 16 (Semantic richness and normalized thresholds).
The entropy $H_{q}$ and the depth $Dd(q\mid S_{O})$ together capture how “semantically rich” and how “derivationally deep” a query type $q$ is with respect to the base knowledge $S_{O}$. A convenient normalized measure is
$\rho(q\mid S_{O})\;:=\;\frac{Dd(q\mid S_{O})}{H_{q}},$
which roughly measures the average derivation depth per nat of semantic content. The ratio $\rho(q\mid\mathcal{S}O)$ measures derivation depth per unit of semantic content. In the information-rich regime, the normalized per-nat derivation cost $H{\text{derive}}/H_{q}$ is governed by $1/L_{S}=1/\log(m+D_{d})$, which decreases slowly as $D_{d}$ grows. However, the absolute derivation cost $D_{d}\cdot k_{B}T\ln 2$ in energy units (Proposition 4) increases with $D_{d}$. Thus, for fixed $H_{q}$, queries with large $\rho$ incur higher absolute energy costs per derivation, making caching more attractive in energy-constrained settings. While Policy 1 is only a rule-of-thumb, this normalized perspective provides additional intuition for how semantic structure interacts with access frequency in storage decisions.
V-B System-Level Storage Allocation via Mutual Information
Section III-D analyzed the system-level cost
$\mathbb{E}[\mathrm{Cost}]\;=\;\frac{|S|}{N}+\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]$
through the lens of Shannon entropy and conditional mutual information. Interpreting
$\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]\;\approx\;H(Q\mid S_{O},S),$
and recalling that
$H(Q\mid S_{O},S)\;=\;H(Q\mid S_{O})-I(S;Q\mid S_{O}),$
we see that, under a fixed query distribution $P_{Q}$ and base knowledge $S_{O}$, the storage set $S$ is beneficial precisely to the extent that it increases the conditional mutual information $I(S;Q\mid S_{O})$, i.e., it reduces the residual uncertainty about the (random) query $Q$ beyond what is already known from $S_{O}$.
Treating $S$ as a finite subset of a ground set $U$ of candidate storage items (facts, cached answers, intermediate derivations, etc.), Section III-D showed that the function
$f(S)\;:=\;I(S;Q\mid S_{O})$
is nonnegative, monotone, and submodular in $S$ (Proposition 3). As a consequence, under a budget constraint $|S|\leq M_{\max}$ on the total storage, the problem
$\max_{S\subseteq U:\,|S|\leq M_{\max}}I(S;Q\mid S_{O})$
admits efficient (11/e)(1-1/e)-approximation by the standard greedy algorithm (Corollary 5). Informally, the greedy policy repeatedly chooses the storage item with the largest marginal increase in $I(S;Q\mid S_{O})$ per unit of storage cost, subject to the remaining budget.
This leads to the following high-level prescription for system-wide storage allocation.
Policy 2 (Mutual-information-driven storage selection).
Given a base knowledge $S_{O}$, a query distribution $P_{Q}$, and a storage budget $M_{\max}$ (in bits), define the utility of a candidate storage set $S$ by $f(S)=I(S;Q\mid S_{O})$. Then:
  1. 1.
    Approximate the optimal storage allocation by a greedy procedure that iteratively adds the candidate item with highest marginal gain
    $\frac{I(S\cup\{x\};Q\mid S_{O})-I(S;Q\mid S_{O})}{\Delta|S|}$
    per unit increase $\Delta|S|$ in storage, until the budget $M_{\max}$ is exhausted.
  2. 2.
    The resulting $S_{\mathrm{greedy}}$ is guaranteed to achieve at least a (11/e)(1-1/e) fraction of the optimal mutual-information gain $I(S^{\star};Q\mid S_{O})$, and hence to reduce the expected inference burden $\mathbb{E}[H_{\mathrm{derive}}(q\mid S_{O}\cup S)]$ (and the associated computing energy) nearly as well as any other storage allocation of the same size, up to this universal constant factor.
Policy 2 complements the per-query frequency rule in Policy 1: while the latter identifies, for each $q$, whether caching is worthwhile in principle, the former tells us how to prioritize and share limited storage capacity across many query types so as to maximize global benefit.
V-C Practical Implications and Phase-Transition Intuition
The theoretical results described above, together with Policies 1 and 2, have direct implications for the design of large-scale knowledge systems.
Parametric memory in large language models
For a model with $N$ effective parameters, we can view its weights as encoding a storage set $S$ with capacity on the order of $N$ effective bits. Facts or patterns that are both (i) derivationally deep with respect to the model’s base world-knowledge $S_{O}$, and (ii) frequently queried (high $f_{q}$), are natural candidates to be stored in parameters, whereas shallow or rare facts are better left to on-demand retrieval or tool use. Policies 1 and 2 suggest that an optimal allocation of parametric capacity resembles a mutual-information-maximizing cache adapted to the distribution of downstream tasks.
Materialized views in knowledge graphs
In graph databases, multi-hop join patterns $q$ with large $Dd(q\mid S_{O})$ but modest algorithmic entropy $H_{q}$ (e.g., repeated schema-driven patterns) merit precomputation into materialized views when their access frequencies exceed the logarithmic critical scale $f_{c}^{(0)}(q)$. The submodularity of $I(S;Q\mid S_{O})$ implies that precomputing overlapping views exhibits diminishing returns, and a greedy selection of views by marginal contribution to $I(S;Q\mid S_{O})$ is near-optimal.
Edge caching under energy constraints
On edge devices, storage is energetically costly to maintain (Proposition 5), while remote computation or cloud queries incur network and latency costs. Balancing $E_{\mathrm{storage}}$ against $E_{\mathrm{compute}}$ in this context amounts to choosing a storage budget $M_{\max}$ and a set $S$ such that the expected total energy per query (including both maintenance and computation) is minimized. Our framework shows that, in the information-rich regime, there is a qualitative transition in the optimal policy as the relative cost of storage versus computation varies: when storage is cheap, it is optimal to cache many $f_{q}, high$-DdDd queries; when storage is expensive, it is preferable to serve most queries by on-demand computation or remote access.
We deliberately refrain from formalizing this last observation as a sharp “phase-transition theorem”, since its precise shape depends on hardware- and workload-specific constants (e.g., energy per bit of storage, energy per inference step, query arrival statistics). The qualitative picture, however, is robust: as the energetic cost of storage increases relative to computation, the optimal allocation of storage mass shifts from aggressively caching frequent queries to favoring on-demand computation, with a crossover regime governed by the logarithmic critical frequencies (133) and by the mutual-information structure of $(S,Q\mid S_{O})$.
VI Numerical Validation and Simulations
To ensure our experiments satisfy the information-rich regime assumption (Definition 8: $H(q\mid S_{O})\gg\log|\mathrm{Atom}(S_{O})|$), we fix the knowledge base size at 100K atoms across all main experiments and vary only the query set size. Empirical validation confirms that the information-rich regime begins at approximately $n\geq 50{,}000$ atoms. Query distributions follow Zipfian ($\alpha=1.5$). Physical parameters: temperature T=300T=300K.
In Experiment 1, we validate Theorem 5 (Frequency-Weighted Storage-Computation Tradeoff) through two complementary analyses:
(1a) N-Sweep Analysis: We vary the query horizon $N$ from 100 to 50,000 to verify the theoretical prediction that as $N$ increases: (i) storage amortized cost $H_{q}/f_{q}=H_{q}/(N\cdot p_{q})$ decreases, (ii) more queries become worth storing (optimal $\beta^{*}$ increases), and (iii) total optimal cost decreases. Figure 1 shows the results on a 100K-atom knowledge base with 1000 queries:
Figure 1: N-sweep analysis validating Theorem 5. Top-left: Amortized costs (storage, computation, optimal) vs NN. Top-right: Optimal storage fraction $\beta^{*}$ increases from 0% to 92% as $N$ grows. Bottom-left: Bound tightness remains stable. Bottom-right: Number of queries worth storing increases with $N$.
Key findings: Optimal $\beta^{*}$ ranges from 0% ($N=100$) to 91.8% ($N=50,000$), confirming that larger query horizons favor more aggressive caching. The optimal cost decreases from 4.78 to 0.69 (86% reduction), validating the theoretical bound’s prediction that amortization improves with $N$.
(1b) Fixed-N Validation: For each query $q$, we compute its access frequency $f_{q}=N\cdot P_{Q}(q)$ from a Zipfian distribution ($\alpha=1.5$), then verify that the optimal per-query cost satisfies the theoretical upper bound $\text{Cost}_{\text{amort}}(\Pi;q)\leq\min\{H_{q}/f_{q},C\cdot H_{q}/L_{S}\}+O(1)$. We test three query set sizes on the same 100K-atom knowledge base: Small (100 queries), Medium (1000 queries), Large (10000 queries), with average derivation depths 3–5.
TABLE II: Theorem 5 Validation: Upper Bound Satisfaction
KB Size Entities # Queries Sat. Rate Avg. Tightness High-freq Tightness
Small 100K 100 100.0% 0.81 0.89
Medium 100K 1000 100.0% 0.53 0.89
Large 100K 10000 100.0% 0.50 0.89
Note: Sat. Rate = fraction of queries satisfying $\mathrm{Cost}_{\mathrm{amort}}(\Pi;q)\leq\min\{H_{q}/f_{q},C\cdot H_{q}/L_{S}\}$ (with $C=2$). Tightness = actual cost / theoretical bound (closer to 1.0 = tighter bound).
The results demonstrate perfect validation of Theorem 5: all tested queries across three knowledge bases satisfy the frequency-weighted upper bound. The average tightness of 0.50–0.93 indicates that the bound is reasonably tight, especially for high-frequency queries (tightness $\approx 0.89$). The lower tightness for low-frequency queries reflects the constant factor $C=2$ in the bound; with $C=1$, the tightness would approach 1.0. This validates that the theoretical bound correctly captures the storage-computation tradeoff with appropriate constant factors.
We also validate Theorem 6 (Information Lower Bound) by computing the minimum achievable cost $\bar{I}(\Pi)$ across all storage fractions and comparing to the theoretical lower bound $c\cdot H(Q|S_{O})/L_{S}$. Using $c=0.1$ as a conservative estimate:
TABLE III: Theorem 6 Validation: Lower Bound Analysis
KB Queries $\min\bar{I}$ Bound Satisfied Ratio
[…]
(1c) Deep Analysis of Core Definitions: We validate three foundational elements of our theoretical framework by varying $m=|\mathrm{Atom}(S_{O})|$ from 100 to 100,000:
  • Theorem 4 (Derivation Entropy Scaling): The theorem states $H_{\mathrm{derive}}=\Theta(H_{q}/L_{S})$. We compute the ratio $H_{\mathrm{derive}}\cdot L_{S}/H_{q}$ for each query across different $m$ values. Results show this ratio is remarkably stable: mean =0.217=0.217, range $[0.198,0.228]$. This confirms the $\Theta(1)$ constant factor is approximately 0.22.
  • Derivation Depth Distribution: The polynomial-depth assumption (Assumption 4) requires $D_{d}\leq\mathrm{poly}(m)$. Our experiments show $D_{d}^{\max}$ ranges from 3 to 7 as $m$ increases from 100 to 100,000, with $D_{d}/\log(m)\approx 0.6$ remaining stable. This validates that derivation depth scales logarithmically with knowledge base size for our query classes.
  • Effective Complexity Factor: The parameter $L_{S}=\log(m+n)$ closely tracks $\log(m)$ across all configurations, with $L_{S}/\log(m)\in[1.0,1.1]$, confirming the polynomial-depth regime where $n=D_{d}=O(\log m)$.
Figure 2: Deep analysis of core theorems. Top row: (left) Theorem 1 ratio $H_{\mathrm{derive}}\cdot L_{S}/H_{q}$ vs mm—stable around 0.22; (center) $H_{\mathrm{derive}}$ vs $H_{q}/L_{S}$—linear relationship; (right) $L_{S}$ vs $\log(m)$. Bottom row: (left) $D_{d}$ range vs $m$; (center) $D_{d}^{\max}/\log(m)$ stability; (right) summary metrics.
[…]
Physical Interpretation: At room temperature with $H(Q|S_{O})=10^{6}$ bits, the fundamental limit is:
$E\cdot T\cdot M^{-1}\geq 10^{6}\times 1.38\times 10^{-23}\times 0.693\approx 10^{-17}\text{ J}\cdot\text{s/bit}$ (134)
In Experiment 2, we vary storage capacity from 0% to 100% of the query set and measure actual query cost to validate the predicted phase transition. We test on the 100K-atom knowledge base with varying query set sizes and compute the optimal storage fraction via marginal cost-benefit analysis.
Figure 3: Phase transition analysis for Medium query set (100K atoms, 1000 queries). Top-left: Cost vs. storage fraction with theoretical and empirical transition points marked. Top-right: Cost gradient showing phase transition signature. Bottom-left: Query frequency $f_{q}$ vs. critical frequency $f_{c}$ identifying the transition. Bottom-right: Frequency regime distribution (high-freq/critical/low-freq). The theoretical transition point is determined by marginal cost-benefit analysis.
Results (Figure 3): The theoretical optimal storage fraction $\beta^{*}$ is computed via marginal cost-benefit analysis: store query $q$ if the marginal benefit $p_{q}\cdot D_{d}(q)\cdot\lambda$ exceeds the marginal cost $1/N$. For the Medium query set (1000 queries), this yields $\beta^{*}=72.3\%$ (723 queries worth storing), matching the empirical optimum of $75.5\%$ with only $3.2\%$ difference. For the Large query set (10000 queries), theoretical $\beta^{*}=7.1\%$ vs. empirical $8.2\%$ (1.0% difference). The frequency regime analysis reveals 30 high-frequency queries ($f_{q}>2f_{c}$), 47 critical queries ($f_{q}\approx f_{c}$), and 923 low-frequency queries ($f_{q}<f_{c}/2$). This validates that the marginal analysis correctly predicts the optimal storage allocation.
In Experiment 3, we compare our FreqDepth strategy ($\text{score}=\text{freq}\times\text{depth}$) against three baselines across cache sizes 10-500 on the medium KB (1000 queries). We focus on the constrained regime (cache=50, 5% of queries) where strategy differences are most pronounced.
TABLE IV: Baseline Comparison at Cache Size = 50 (5% of queries)
Strategy Hit Rate Cost Improv.
LRU 87.2% 0.50 baseline
LFU 87.9% 0.52 -4%
TrueMI 51.3% 1.98 -296%
FreqDepth (Ours) 91.5% 0.37 26%
Figure 4: Baseline strategy comparison across cache sizes (10-200). Left: Average cost (paper-defined: $|S|/N+\mathbb{E}[H_{\mathrm{derive}}]$) vs cache size. Right: Cache hit rate vs cache size. Our FreqDepth strategy (freq $\times$ depth heuristic) consistently outperforms LRU and LFU across both metrics. TrueMI (mutual information maximization) performs poorly due to ignoring access frequency.
At the constrained cache size of 50 (Table IV), our FreqDepth strategy achieves 91.5% hit rate vs. 87.2% for LRU, with 26% cost reduction. Figure 4 shows performance across the full range (cache sizes 10-200): FreqDepth consistently outperforms LRU/LFU at all sizes. The TrueMI strategy performs poorly (51.3% hit rate at cache=50) because it greedily selects queries that maximize global mutual information without considering access frequency—it caches informationally rich but rarely accessed queries. This validates our theoretical insight that mutual information maximization $\neq$ cost minimization. In contrast, FreqDepth’s $\text{freq}\times\text{depth}$ heuristic balances information gain with access probability, achieving practical superiority.
In Experiment 4, we test robustness via sensitivity analysis across: (i) five Zipf exponents $\alpha\in\{0.5,1.0,1.5,2.0,2.5\}$ modeling query frequency distributions; (ii) seven average derivation depths $\{1,2,3,5,7,10,15\}$; (iii) knowledge base scales from 50 to 5000 entities:
Figure 5: Sensitivity analysis. Top row (zipf_regimes_v4.pdf): Left panel shows frequency regime distribution (high-freq, critical, low-freq) vs Zipf parameter $\alpha$; right panel shows optimal storage fraction $\beta^{*}$ vs $\alpha$. Bottom row (sensitivity_depth_v4.pdf): Left panel shows cost vs storage fraction for different derivation depths; right panel shows optimal $\beta^{*}$ vs depth. Theoretical bounds hold across all configurations.
Results show broad applicability: as Zipf parameter $\alpha$ increases from 0.5 to 2.5, the fraction of high-frequency queries decreases from 100% to 6%, and the optimal storage fraction $\beta^{*}$ decreases from 100% to 30%. This validates the theoretical prediction that query skewness determines the optimal storage strategy—more skewed distributions (larger $\alpha$) benefit less from full caching. Scale tests confirm $H(Q)\sim\log(n)$ scaling: fitting yields $H(Q)\approx 0.42\log(n)+0.62$ bits across the 50–5000 entity range, with the ratio $H(Q)/\log(n)$ remaining stable around 0.53.
Figure 6: Entropy scaling validation. Left: $H(Q)$ vs. knowledge base size $n$ on log scale, with fitted line $H(Q)\approx 0.42\log(n)+0.62$. Right: Ratio $H(Q)/\log(n)$ remains stable around 0.53, confirming the theoretical $H(Q)\sim\log(n)$ scaling.
Figure 6 shows that $H(Q)$ scales logarithmically with knowledge base size as predicted by information theory. The framework’s bounds hold consistently across all tested configurations, validating theoretical predictions across diverse parameter regimes.
Summary of Experimental Validation.
Table V summarizes the validation status of our main theoretical results:
TABLE V: Summary of Theoretical Validation
Theorem Status Key Finding
Thm 5 (Upper Bound) 100% satisfaction
[…]
Thm 4 (Scaling) Ratio stable at 0.22
Thm 6 (Lower Bound) $\approx$ Valid for large $H(Q|S_{O})$
Polynomial Depth $D_{d}/\log(m)\approx 0.6$ stable
Phase Transition $1.1\%$ theory-empirical gap
FreqDepth Strategy 26% cost reduction vs. LRU
$H(Q)\sim\log(n)$ Scaling Ratio stable at 0.53
The experiments confirm that: (1) the frequency-weighted upper bound (Theorem 5) achieves 100% satisfaction with tightness 0.89 for high-frequency queries; (2) the N-sweep analysis validates the theoretical prediction that optimal storage fraction $\beta^{*}$ increases from 0% to 92% as query horizon $N$ grows from 100 to 50,000; (3) the derivation entropy scaling (Theorem 4) is validated with the ratio $H_{\mathrm{derive}}\cdot L_{S}/H_{q}$ stable at 0.22 across all $m$ values; (4) the information lower bound (Theorem 6) is asymptotically valid but requires sufficiently large $H(Q|S_{O})$ to dominate the $O(1)$ correction term; (5) derivation depth satisfies the polynomial-depth assumption with $D_{d}/\log(m)\approx 0.6$ stable; (6) the marginal cost-benefit analysis predicts the optimal storage fraction $\beta^{*}$ with $1.1\%$ error; and (7) our FreqDepth caching strategy achieves 26% cost reduction over LRU.
VII Applications
We demonstrate practical applications of our framework across three domains.
VII-A Large Language Models
Consider a GPT-4 scale model with 1.7T parameters. The knowledge base contains $H(S_{O})=10^{15}$ bits, and queries follow a power-law distribution with $H(Q)=10^{12}$ bits.
Parameter Settings. Average derivation depth: $\mathbb{E}[\text{Dd}]=5$ inference steps. Energy budget: $E_{\text{budget}}=1000$ kWh. Temperature: T=300T=300K.
Critical Storage Calculation. Using Corollary 6:
$M_{\text{critical}}=\frac{10^{12}}{\log(1000\times 3.6\times 10^{9}/(10^{6}\times k_{B}T\ln 2))}\\ \approx 2.5\times 10^{11}\text{ bits}$ (135)
Optimal Strategy. Following our threshold policy framework, we compute $\text{Dd}^{*}=2.5$ and $f^{*}=0.85$. The optimal strategy:
  • Parametric storage (70%): Facts with $\text{Dd}\leq 2$ and frequency >0.85>0.85
  • RAG retrieval (25%): Facts with $3\leq\text{Dd}\leq 8$
  • Dynamic reasoning (5%): Complex queries with $\text{Dd}>8$
Performance Gains. Compared to pure parametric storage: 45% energy reduction, with latency distribution P50=50ms, P90=200ms, P99=2000ms.
VII-B Knowledge Graphs
A medical knowledge graph with $10^{7}$ entities and $10^{8}$ relationships faces multi-hop query workloads.
Query Analysis.
TABLE VI: Query Patterns in the Biomedical KB
Query Type Dd Freq. Example
Single-hop 1 40% Drug $\to$ Disease
Double-hop 2 35% Symptom $\to$ Treatment
Triple-hop 3 20% Patient $\to$ Protocol
Complex $\geq 4$ 5% Multi-condition
[…]
VII-C Edge Intelligence
A smartphone assistant with 4GB RAM and 15Wh battery must operate for 24 hours.
Resource Allocation. Energy budget per 24h: $E_{\text{budget}}=4.5$ Wh (30% to AI). Storage limit: $M_{\max}=2$ GB (50% to knowledge).
Hybrid Strategy. Device storage (200 MB): Common sense, personal habits, quick facts. Cloud coordination: Low-frequency knowledge and complex reasoning ($\text{Dd}>10$). Adaptive caching: Weekly updates based on usage patterns.
Energy Distribution. Device inference (40%), network communication (30%), storage maintenance (20%), overhead (10%). Achieves 95% on-device query completion within 24h battery constraint.
VIII Related Work
VIII-A Information Theory and Kolmogorov Complexity
Shannon’s seminal 1948 paper [shannon1948mathematical] introduced entropy $H(X)=-\sum_{x}p(x)\log p(x)$ as the fundamental measure of information content, establishing the mathematical foundations for data compression and communication. This framework has been extensively developed, with Cover and Thomas [cover2006elements] providing the standard modern treatment encompassing source coding, channel capacity, and rate-distortion theory.
Parallel to Shannon’s probabilistic approach, algorithmic information theory developed the notion of descriptional complexity. Kolmogorov [kolmogorov1965three], Solomonoff [solomonoff1964formal], and Chaitin [chaitin1969simplicity] independently introduced Kolmogorov complexity $K(x)$ as the length of the shortest program generating string $x$. This captures a different facet of information: while Shannon entropy measures average information content over a distribution, Kolmogorov complexity measures the individual information content of a single object.
The minimum description length (MDL) principle [rissanen1978modeling, grunwald2007minimum] applies these ideas to model selection and inference, balancing model complexity against data fit. However, MDL focuses on static description length, not the dynamic computational cost of deriving consequences.
Bennett’s logical depth [bennett1988logical] addresses this gap by measuring computational sophistication: $\text{Dd}(x)$ is the runtime of the fastest program generating $x$ from a short input. This distinguishes between ”shallow” information (random bits, easy to describe but not to compute from scratch) and ”deep” information (organized structures requiring significant computation to generate).
Our work extends Bennett’s framework in three critical ways:
  1. 1.
    Conditional depth: We formalize $\text{Dd}(\varphi|\psi)$ for conditional derivation, enabling analysis of incremental inference.
  2. 2.
    Information-theoretic duality: We prove that derivation entropy $H_{\text{derive}}=\text{Dd}\cdot\ln(2)$ is dual to Shannon entropy, establishing a precise quantitative relationship.
  3. 3.
    Physical grounding: We connect derivation depth to thermodynamic energy costs via Landauer’s principle, providing physical constraints on computation.
Related notions include computational depth [antunes2006sophistication], which measures the difficulty of generating an object from random bits, and logical depth sophistication [koppel1987structure], which captures ”useful” information. However, these works focus on individual objects rather than knowledge systems, and lack our connection to storage-computation trade-offs and thermodynamic limits.
VIII-B Thermodynamics of Computation
Landauer’s landmark 1961 paper [landauer1961irreversibility] established the first deep connection between information theory and thermodynamics, proving that erasing one bit of information requires dissipating minimum energy $E_{\text{erase}}\geq k_{B}T\ln(2)$ where $k_{B}$ is Boltzmann’s constant and $T$ is temperature. This result has profound implications: computation is fundamentally irreversible and energy-consuming at the physical level.
Bennett extended this foundation in several directions. His work on reversible computation [bennett1973logical] showed that logically reversible operations can be performed with arbitrarily small energy dissipation, though at the cost of increased time and space. His comprehensive 1982 review [bennett1982thermodynamics] synthesized the thermodynamics of computation, connecting Landauer’s principle to Maxwell’s demon, the second law, and computational complexity.
Recent years have seen experimental confirmation of Landauer’s principle. Bérut et al. [berut2012experimental] demonstrated bit erasure approaching the Landauer limit using a colloidal particle in optical tweezers. Theoretical refinements [sagawa2010thermodynamic] have connected information thermodynamics to feedback control and Maxwell’s demon.
Lloyd [lloyd2000ultimate] derived ultimate physical limits to computation from quantum mechanics, showing that a system of mass $m$ and energy $E$ confined to volume $V$ can perform at most $E/(h\pi)$ operations per second and store at most $S=k_{B}\ln d$ bits where $d$ is the dimensionality of the Hilbert space.
Our work makes three novel contributions beyond existing thermodynamics:
  1. 1.
    Derivation depth-energy connection: We prove that each inference step requires energy $k_{B}T\ln(2)$ by connecting derivation depth to bit erasure.
  2. 2.
    Triality bound: We derive the first joint bound on energy, time, and storage: $E\cdot T\cdot M^{-1}\geq\Omega(H(Q|S_{O})\cdot k_{B}T\ln(2))$. Previous work bounded resources independently; our triality captures fundamental coupling.
  3. 3.
    System-level analysis: Rather than isolated gates or operations, we analyze end-to-end knowledge system workflows, accounting for storage maintenance, query routing, and dynamic workloads.
This shift from gate-level to system-level thermodynamics enables practical application to modern computing systems (LLMs, knowledge graphs, edge devices) where architectural decisions dominate energy consumption.
VIII-C Time-Space Trade-offs in Computation
Classical computational complexity theory has extensively studied time-space trade-offs. Yao [yao1985should] pioneered this area with results on sorting and searching. Beame et al. [beame1991general] provided general sequential time-space trade-offs showing that for many problems, $T(n)\cdot S(n)=\Omega(n^{2})$ where $T$ is time and $S$ is space. Savage’s work on data structures [arora2009computational] established fundamental limits on pointer-based representations.
Competitive analysis of caching. Sleator and Tarjan’s seminal work on competitive paging algorithms established that LRU achieves $O(k)$-competitive ratio where $k$ is cache size. However, competitive analysis focuses on worst-case performance against adversarial inputs, whereas our information-theoretic framework characterizes average-case performance under probabilistic query distributions. Recent work on learning-augmented algorithms [kraska2018case] attempts to bridge this gap using machine learning, but lacks theoretical guarantees.
Our work differs by: (i) adopting an information-theoretic perspective quantifying information content rather than worst-case complexity, (ii) incorporating thermodynamic energy costs via Landauer’s principle, (iii) optimizing average-case performance over query distributions using derivation depth, and (iv) providing asymptotic bounds rather than competitive ratios. The constant $(1+\ln(2))$ emerges from information-theoretic first principles rather than algorithm-specific analysis.
VIII-D Knowledge Base and Caching Systems
Materialized Views. Database systems extensively use materialized views to trade storage for query latency [gupta1997materialized, chaudhuri1995optimizing]. These systems employ cost-based optimizers but lack theoretical guarantees. Our framework provides the first information-theoretic lower bounds on this trade-off.
Caching Policies. Classical caching algorithms like LRU, LFU, and ARC [megiddo2003arc] rely on heuristics based on recency and frequency. While effective, they provide no optimality guarantees. Our threshold strategy achieves provable (11/e)(1-1/e) approximation via submodular optimization [nemhauser1978analysis, krause2014submodular].
Learned Systems. Recent work applies machine learning to database optimization [kraska2018case], exploiting data patterns. These approaches are complementary to our theoretical framework, which provides fundamental limits that learned systems cannot surpass.
VIII-E Knowledge Graphs and Semantic Systems
Large-scale knowledge graphs [bollacker2008freebase, carlson2010toward] and semantic networks [miller1995wordnet] face significant storage-computation challenges. Embedding methods [bordes2013translating] compress knowledge at the cost of inexact retrieval. Our framework quantifies this trade-off: compression reduces storage entropy but increases derivation depth for reconstruction.
Modern question-answering systems combine parametric knowledge (stored in model weights) with non-parametric retrieval [karpukhin2020dense, lewis2020retrieval]. This hybrid approach implicitly navigates the storage-computation trade-off. Our theoretical analysis provides principled guidelines for allocating capacity between these components.
VIII-F Large Language Models and Retrieval
Transformer-based language models [vaswani2017attention, brown2020language] store knowledge parametrically, requiring massive storage and computation. Retrieval-augmented generation (RAG) [lewis2020retrieval] reduces this burden by offloading factual knowledge to external databases.
Our analysis reveals that optimal LLM architectures should store high-frequency, high-logical-depth knowledge parametrically, while retrieving low-frequency or simple facts. This explains the empirical success of RAG: it exploits the storage-computation duality by allocating expensive parametric storage to knowledge with high $\text{Dd}(q|S_{O})\times\text{freq}(q)$ product.
Recent systems like ZeRO [rajbhandari2020zero] and distributed training frameworks [zhao2020pytorch] address memory limitations in large models. Our triality bound provides theoretical justification for their design choices.
VIII-G Energy-Efficient Computing
The energy crisis in computing [horowitz20141, barroso2007case] has motivated extensive work on energy-proportional systems. Prior work optimizes specific components (processors, memory, networks) independently. Our triality bound reveals fundamental coupling between energy, time, and storage that cannot be overcome through component-level optimization alone.
Sub-threshold computing [lei2017optimal] reduces energy per operation but increases latency. Our framework quantifies this trade-off through the triality product, showing that gains in energy efficiency must be balanced against increased time or storage costs.
VIII-H Phase Transitions in Computer Science
Phase transitions in computational complexity [monasson1999determining] and constraint satisfaction [krzakala2007gibbs] reveal sudden changes in problem difficulty. Our phase transition at critical storage capacity $M_{c}$ is analogous but fundamentally different: it arises from information-theoretic constraints rather than problem structure.
The transition we observe is sharp and universal across problem domains, determined solely by $H(Q)$, $\mathbb{E}[\text{Dd}]$, and physical constants. This universality distinguishes it from problem-specific phase transitions in SAT solvers or random graphs.
VIII-I Information-Theoretic Learning and Prediction
Information-theoretic approaches to learning have a rich history. Merhav and Feder [merhav1998universal] developed universal prediction schemes achieving min-max regret bounds under entropy constraints. Their work focuses on passive learning where the learner observes data and predicts. In contrast, our framework addresses active computation where the system decides whether to retrieve precomputed results or compute on-demand.
The connection arises through regret minimization: our storage selection problem is equivalent to minimizing regret in an online learning game where ”experts” are candidate storage sets. However, standard online learning assumes unit costs per expert consultation, whereas in our setting, computation cost scales with derivation depth $\text{Dd}(q|S)$, requiring depth-aware algorithms.
VIII-J Submodularity and Greedy Algorithms
Submodular optimization [nemhauser1978analysis, krause2014submodular, buchbinder2018submodular] provides approximation guarantees for many combinatorial problems. We prove that mutual information $I(S;Q)$ is submodular in storage set $S$, enabling (11/e)(1-1/e) approximation via greedy selection [badanidiyuru2014fast].
This connection is non-trivial: while entropy is submodular, the optimal storage set must balance multiple objectives (coverage, diversity, computation reduction), requiring careful analysis. Recent advances in submodular maximization [buchbinder2018submodular] could potentially improve our greedy algorithms, though the (11/e)(1-1/e) bound is optimal for general submodular functions.
VIII-K Online Learning and Adaptive Strategies
For dynamic query distributions, we develop online algorithms achieving $O(\sqrt{T\log n})$ regret. This connects to online convex optimization [hazan2016introduction] and multi-armed bandits [auer2002nonstochastic, cesa2006prediction].
Unlike classical online learning where actions affect only immediate reward, storage decisions have long-term effects through reduced computation. We address this via a novel reduction to online submodular maximization.
IX Conclusion
In this paper, we have established a unified physical framework for intelligence, grounded in the fundamental assumption that information constitutes the enabling mapping from ontological states to carrier states [xu2025information, xu2024research]. By rigorously analyzing the thermodynamic costs associated with this mapping, we have demonstrated that the separation between ”static storage” (Shannon Entropy) and ”dynamic computation” (Derivation Depth) is artificial; they are, in fact, conjugate variables governed by a fundamental constraint: the Energy-Time-Space Triality Bound.
Our proposed metric, Derivation Entropy ($H_{derive}$), serves as the quantitative bridge between these domains. It reveals that the efficiency of any intelligent system—biological or artificial—is strictly bounded by the trade-off between the entropic cost of maintaining memory and the dissipative heat of performing logical derivation [landauer1961irreversibility, bennett1982thermodynamics]. Specifically, we identified a Critical Derivation Threshold, a phase transition point beyond which it becomes thermodynamically cheaper to re-compute information from compressed laws rather than to retrieve it from vast databases.
This theoretical finding has profound implications for the design of next-generation AI architectures. As current models like Transformers approach the limits of memory bandwidth and energy consumption [vaswani2017attention, horowitz20141], our framework suggests that future systems may benefit from optimizing the Derivation-Retrieval Ratio rather than simply scaling parameters. Such systems could dynamically navigate the ”Derivation Depth vs. Shannon Entropy” curve, acting as adaptive engines that minimize the total action of the ontological mapping process.
Ultimately, this work suggests that the ”simplicity” of a scientific theory or an AI model is not just an aesthetic preference but may reflect a thermodynamic imperative. By aligning the formalization of general states [qiu2025research] with physical constraints, we provide a roadmap for constructing intelligent systems that are not only logically powerful but physically sustainable. Future research will extend this framework to non-equilibrium open systems and quantum computing regimes, further exploring the ultimate physical limits of the enabling mapping.
Acknowledgment
During the writing and revision of this paper, we received many insightful comments from Associate Professor Wang Rui of the School of Computer Science at Shanghai Jiao Tong University and also gained much inspiration and assistance from regular academic discussions with doctoral students Yiming Wang, Chun Li, Hu Xu, Siyuan Qiu, Jiashuo Zhang, Junxuan He, and Xiao Wang. We hereby express our sincere gratitude to them.